### v1

# The Basis of Foot-Sortable Sock Orderings

Abstract.
A sock ordering is a finite string of elements, each of which has a color. Foot-sorting is a non-deterministic sorting algorithm with the aim of grouping all socks of the same color together, under the constraints of a stack structure; a sock ordering is foot-sortable if there is some way to apply the foot-sorting algorithm such that this aim is realizable. In introducing sock orderings and the foot-sorting algorithm, Defant and Kravitz ask whether there is some list of patterns which all foot-sortable sock orderings must avoid, thus characterizing which sock orderings are foot-sortable. We enumerate this list, the basis of foot-sortable sock orderings, and show that this basis is infinite. We also describe a deterministic polynomial-time algorithm for sorting socks.
1. Introduction
The idea of sorting permutations, introduced by Donald Knuth in [2], considers sorting a given permutation of $\{1,2,\dots,n\}$ to a desired output under the constraints of a stack structure. In [2], Knuth showed that a permutation is sortable if and only if the permutation avoids the pattern 231. In other words, the permutation $a_{1}a_{2}\cdots a_{n}$ is sortable using the stack structure if and only if there is no set of indices $1\leq i<j<k\leq n$ such that $a_{k}<a_{i}<a_{j}$.
Extending Knuth’s findings, Defant and Kravitz introduced the concept of foot sorting for particular patterns in [1]. This alteration allows for single orderings, called sock orderings, to contain multiple socks of the same color, whereas the original work dealt with orderings containing the elements $\{1,2,\dots n\}$ exactly once. A significant difference in sorting a sock ordering when compared to a permutation is that in sock orderings, there is no pre-specified sorted word which the foot-sorting algorithm must produce. Because of the freedom in output, sorting a sock ordering that is given to be foot-sortable may be easier than sorting a permutation pattern.
Defant and Kravitz prove a similar result to Knuth in [1]: that a sock ordering $\rho$ is foot-sortable if and only if there is some word $\nu$ representing $\rho$ such that $\nu$ avoids the pattern 231. While for sock orderings with only a few colors, one could search for such a word which avoids 231, the process of generating all possible words and checking whether they avoid 231 is very computationally intensive. Allowing for multiple iterations of the stack-sorting procedure, Xia investigated a deterministic foot-sorting algorithm in [3] which hinges on pattern avoidance within the stack, though our work focuses on the sock orderings which are foot-sortable in a single iteration of the sorting procedure. Recently, [4] provided a polynomial-time algorithm for deciding the foot-sortability of an arbitrary sock ordering; we later offer a deterministic algorithm which operates in polynomial time and classifies sock orderings as sortable or unsortable, successfully sorting the orderings for which it is possible.
Defant and Kravitz introduce the idea of a basis of foot-sortable sock orderings in [1]. The basis of foot-sortable sock orderings is the smallest set of sock orderings such that a sock ordering $X$ is foot-sortable if and only if no member of the set is a subpattern of $X$; we introduce several key ideas in relation to this basis and subpatterns in the following section. Defant and Kravitz pose two interesting questions about the basis of foot-sortable sock orderings in [1]:
  1. (1)
    Is it possible to enumerate the basis of foot-sortable sock orderings?
  2. (2)
    Is the basis of foot-sortable sock orderings finite?
In Section 2, we will define a infinite set, $\Gamma$, of minimally unsortable sock orderings. The primary result of our work is the following theorem which proves that $\Gamma$ is the basis for foot-sortable sock orderings.
Theorem 1.
If $X$ is an unsortable sock ordering, then a subpattern of $X$ is in $\Gamma$.
The proof of Theorem 1 leads naturally to a deterministic algorithm which we present in Section 3.
2. Notation and Preliminaries
Before enumerating the basis of foot-sortable sock orderings, we lay out some important definitions regarding sock orderings and the patterns they contain as well as provide some examples of pattern containment.
Definition 2.
A sock ordering $X$ is a finite sequence of socks, each of which has a specified color.
We will use $|X|$ to denote the length of a sock ordering $X$, the number of socks in $X$. Throughout our work, we will use capital letters to denote sock orderings, while we use lowercase letters to denote specific socks or colors. At times, we refer to sock orderings as strings in relation to our algorithms.
Definition 3.
Let $X=x_{1}x_{2}\cdots x_{n}$ and $Z=z_{1}z_{2}\cdots z_{n}$ be sock orderings (so $|X|=|Z|$), and let $\mathcal{A}_{X}$ and $\mathcal{A}_{Z}$ be the sets of distinct sock colors appearing in $X$ and $Z$ respectively. $X$ is equivalent to $Z$ (denoted $X\cong Z$) if there exists a bijective mapping $\varphi:\mathcal{A}_{X}\rightarrow\mathcal{A}_{Z}$ such that $\varphi(x_{i})=z_{i}$ for all $1\leq i\leq n$.
Definition 4.
Let $X$ and $Y$ be sock orderings. We say that $X$ contains $Y$ and write $Y\subseteq X$ if deleting zero or more socks from $X$ yields exactly $Y$, and we say $a\in X$ if $a\subseteq X$. If $Y\subseteq X$, we say that $Y$ is a subordering of $X$. In contrast, we say that $Y$ is a subpattern of $X$ and write $Y\preceq X$ if there exists a subordering $Z\subseteq X$ of $X$ such that $Z\cong Y$. If $Y$ is not a subpattern of $X$, then we write $Y\npreceq X$ and say that $X$ avoids $Y$. For a color $a\in X$, we let $X-a$ be the subordering of $X$ formed by removing every sock of color $a$ from $X$.
Because of the many equivalent ways to represent one sock ordering, it may not always be immediately clear whether a sock ordering is a subpatterns of another. The following example will illustrate:
Example 5.
$babc\preceq abacbab$.
We have $babc\nsubseteq abacbab$, but $abac\subseteq abacbab$ and $abac\cong babc$. Therefore, $babc\preceq abacbab$.
To sort, or attempt to sort, sock orderings, we utilize the foot-sorting algorithm proposed by [1], and we outline the rules here.
Definition 6.
foot-sorting is a non-deterministic algorithm which operates on a sock ordering $X$. At any point in the foot-sorting process, one may choose to:
  • Stack a sock: Move the leftmost sock of the remaining input ordering to the top of the stack, or
  • Unstack a sock: Remove the outermost sock on the foot (the topmost element of the stack) and place it on the right of the output.
Though we utilize the idea of foot-sorting and socks, throughout, we will refer to the sorting process using primarily stack-centered language.
At some points, this algorithm has set steps. For example, the first step in foot-sorting always must be to place the first sock, if it exists, into the stack. Similarly, once nothing remains of the original input, all socks in the stack must be removed beginning with the topmost element of the stack and moving downward.
We denote a particular stage in the foot-sorting process as $(S,R)$. $S$ represents the contents of the stack, with left-to-right in $S$ corresponding with bottom-to-top in the stack (or inside-to-outside on the foot, in the language of [1]), while $R$ represents what remains of the input ordering. Before any sorting steps are applied, we may say that an input ordering $X$ appears as $(\emptyset,X)$. To illustrate this process and notation, we provide an example of foot-sorting here.
Example 7.
Using the foot-sorting algorithm, we may obtain the output $aaccbb$ from the input sock ordering $abcabc$.
  1. (1)
    We equivalently write $abcabc$ as $(\emptyset,abcabc)$. Move the leftmost sock of the remaining input ordering, $a$, to the top of the stack. This yields $(a,bcabc)$.
  2. (2)
    Remove the topmost sock of the stack, $a$, and place it on the right of the output, which is currently empty. We now have $(\emptyset,bcabc)$ with output $\kappa=a$.
  3. (3)
    Move the leftmost sock of the remaining input ordering, $b$, to the top of the stack, leaving $(b,cabc)$ and output $\kappa=a$.
  4. (4)
    Stack $c$ to leave $(bc,abc)$ and output $\kappa=a$.
  5. (5)
    Stack $a$ to leave $(bca,bc)$ and output $\kappa=a$.
  6. (6)
    Unstack $a$ to leave $(bc,bc)$ and output $\kappa=aa$.
  7. (7)
    Unstack $c$ to leave $(b,bc)$ and output $\kappa=aac$.
  8. (8)
    Stack $b,c$, leaving $(bbc,\emptyset)$ and output $\kappa=aac$.
  9. (9)
    Unstack, in order, all remaining socks from the stack to leave (,)($\emptyset$,$\emptyset$)( ∅ , ∅ ) and output $\kappa=aaccbb$.
Definition 8.
A stage of the sorting algorithm $(S,R)$ is foot-sortable (or, when clear, sortable) if there is some way to apply the foot-sorting algorithm on $(S,R)$ such that, in the output of the foot-sorting process, all socks of the same color occur consecutively. That is, for some $(S,R)$ containing only the colors $\{a_{1},a_{2},\dots a_{n}\}$, $(S,R)$ may be sorted to $\kappa=a_{\pi(1)}\cdots a_{\pi(1)}a_{\pi(2)}\cdots a_{\pi(2)}\cdots a_{\pi(n)}\cdots a_{\pi(n)}$ where $\pi$ is a permutation of $\{1,\dotsc,n\}$. If there is no such way to apply the foot-sorting algorithm, $X$ is unsortable. We say that a sock ordering $X$ is foot-sortable (resp. unsortable) if $(\emptyset,X)$ is foot-sortable (resp. unsortable).
We see that $abcabc$ in Example 7 is a sortable sock ordering.
The goal of foot-sorting is to begin with an input sock ordering $X$ and apply the steps of the algorithm so that nothing remains in the stack or on the right side. While it can be difficult to decide whether a particular sock ordering is sortable or unsortable, there is an important indicator which can reveal if applying the foot-sorting algorithm has already failed. Note that a sortable sock ordering $X$ may be sorted to a word $\kappa$ such that $aba\not\preceq\kappa$. Due to the nature of the stack structure, once a sock, $a$, appears in the stack, $a$ must appear in the output $\kappa$ before any socks appearing below $a$ in the stack and after any socks appearing above $a$ in the stack. For $(S,R)$ appearing at some point while applying the foot-sorting algorithm, let $S=s_{1}s_{2}\cdots s_{n}$ and set $S^{-1}=s_{n}\cdots s_{2}s_{1}$, the reverse of $S$. It must be that $S^{-1}\subseteq\kappa$. Since in a sorted word $\kappa$, we have $aba\not\preceq\kappa$, and $(aba)^{-1}=aba$, then if $aba\preceq S$, it is not possible to successfully sort the sock ordering. Because of this, if we ever have $aba\preceq S$ while attempting to sort a particular sock ordering, we will immediately stop the sorting attempt as it is not possible to finish sorting after this point. We refer to this $aba$ pattern of distinct colors $a,b$ as a sandwich.
Sandwiches, the colors they contain, and where they appear in sock orderings arise frequently in our work. We may say that the sandwich $aba$ blocks, precedes, or appears before the color $c$, for distinct colors $a,b,c$, at some stage $(S,R)$ if $abac\subseteq SR$. Additionally, we are sometimes concerned about the first sandwich $aba$ which blocks a color $c$. If two distinct sandwiches, $aba$ and $xyx$, precede $c$, then if $a$ and $x$ are distinct colors, $aba$ appears earlier than $xyx$ if the second denoted $a$ appears before the second denoted $x$; if $x$ is $a$, then $aba$ appears earlier than $aya$ if the denoted $b$ appears before the denoted $y$. For example, in $abcabd$, all of the sandwiches $aba,aca,bcb,bab$ block $d$, but $aba$ is the first sandwich which blocks $d$. Finally, we may say $x$-sandwich to describe a sandwich which contains a sock of color $x$.
Interestingly, though it is necessary to avoid $aba$ appearing in the stack while applying the foot-sorting algorithm, a deterministic algorithm which prevents the appearance of sandwiches in the stack does not succeed in sorting all sortable sock orderings; Xia investigated such an algorithm, $\phi_{aba}$, in [3]. We later introduce a deterministic algorithm $\phi_{G}$ which, as a consequence of its operation, prevents $aba$ from appearing in the stack, but has further restrictions that differentiate it from $\phi_{aba}$. It was also shown that $\phi_{aba}$ in [3] may sometimes sort sock orderings which have subpatterns that are unsortable under $\phi_{aba}$. In foot-sorting, however, this is not the case; if there exists no set of steps which may sort some subpattern $Y\preceq X$, there can be no set of steps which may sort $X$. For every foot-sortable sock ordering $S$, $Y\preceq S$ implies that $Y$ is foot-sortable.
Definition 9.
At a particular sorting stage $(S,R)$, some color $a\in SR$ is sortable if both of the following conditions hold:
  1. (1)
    For any $b\in S$ which is distinct from $a$, $ab\not\subseteq S$. (That is, if $a$ appears on the stack, it must be the top element on the stack.)
  2. (2)
    If $a$ appears in $R$, then $bcba\not\subseteq SR$ for any colors $b,c\in SR$ such that $a$, $b$, and $c$ are distinct. (In other words, the color $a$ is not blocked by a sandwich.)
Otherwise, $a$ is unsortable in $(S,R)$. We say that a color $a\in X$ is sortable in $X$ if $a$ is sortable in $(\emptyset,X)$. If no color is sortable in $(S,R)$, then we say that $(S,R)$ is terminal.
The following observation follows from the definitions.
Observation 10.
Let $S$ and $R$ be sock orderings. Suppose $x\in SR$ and that $xy\not\subseteq S$ for any $y\in S$ distinct from $x$. Then the color $x$ is sortable in $(S,R)$ if and only if $x$ is sortable in $SR$.
Definition 11.
A color $a$ is a good sortable color at a stage $(S,R)$ if $a$ is sortable in $(S,R)$ and what remains after sorting $a$ in $(S,R)$ is foot-sortable.
At times, we may reach a stage in which multiple colors are sortable. For some colors $\{a_{1},\dots,a_{n}\}$ in a string $X$, we say $a_{1}$ is the first terminating color from the set if the final sock of color $a_{1}$ appears before the final sock of color $a_{i}$ for all $a_{i}$ in the set distinct from $a_{1}$.
Definition 12.
A trivially unsortable sock ordering is a sock ordering in which no color is sortable. That is, a sock ordering $X$ is trivially unsortable if for every color $a\in X$, there exist distinct colors $b,c\in X$, different from $a$, such that $bcba\subseteq X$.
We often discuss particular sorting choices in an attempt to sort a sock ordering $X$. A move in some sorting procedure consists of sorting exactly one color. That is, when $(S,R)$ appears, a color $x$, identified as a sortable in $(S,R)$, is chosen to be sorted. Any $x$ are removed from the top of $S$, and if $x\in R$, then for every sock in $R$, we apply the following rule:
  • If the sock is the final sock of color $x$, place the sock into the stack, then remove it from the stack and place the sock on the sorted left side. Immediately terminate the procedure at this point.
  • If the sock has color $x$ but is not the final sock of color $x$, stack the sock and immediately unstack it, bringing it to the sorted left side. Look to the next sock on the right side.
  • If the sock has a color which is not $x$, stack the sock, and do not unstack it. Look to the next sock on the right side.
In contrast, a step in the sorting procedure refers to the movement of exactly one sock: either stacking the first sock appearing on the right, or removing the top element of the stack and bringing it to the sorted left side. In general, it takes multiple steps to complete one sorting move. When we say that we sort the color $a$ we mean that we apply a move as describe above of the sortable color $a$.
Definition 13.
Let $X$ be a sock ordering, and let $\mathcal{A}_{X}$ be the set of distinct colors appearing in $X$. A sorting sequence $\theta$ for $X$ is a nonempty sequence $(a_{1},a_{2},\dots,a_{n})$ of elements from $\mathcal{A}_{X}$, each appearing at most once, such that $a_{1}$ is sortable in $(\emptyset,X)$, and for $i>1$, $a_{i}$ is sortable after $a_{i-1}$ has been sorted. We write $\theta(X)=(S,R)$ if, beginning with $(\emptyset,X)$, $(S,R)$ appears after sorting each color from $\theta$ in sequence.
Observation 14.
At some stage $(S,R)$, if a color $x$ only appears as the top color in $S$, then $x$ is sortable with respect to $(S,R)$. Furthermore, if we sort $x$ to form $(S^{\prime},R)$, then $(S,R)$ is sortable if and only if $(S^{\prime},R)$ is sortable.
Definition 15.
We say that a sorting sequence $\theta$ fully sorts $X$ if $\theta(X)=(\emptyset,\emptyset)$. In light of Observation 14, we also say that $\theta$ fully sorts $X$ if we can fully sort $X$ using the prescribed sorting order $\theta$ with the additional stipulation that we immediately sort a color if it only appears as the top color on the stack. Note that, with this definition, a sorting sequence $\theta$ can fully sort a sequence $X$ even if $\theta$ does not contain all of the colors of $X$.
While a sorting sequence $\theta$ must obey the rules of foot-sorting (that is, $\theta$ may never make a move that causes a sandwich to appear in the stack), there is no requirement for $\theta$ to fully sort $X$.
Example 16.
Taking $\theta_{1}=(a,c,b)$, we have that $\theta_{1}(abcabc)=(\emptyset,\emptyset)$, and thus $\theta_{1}$ fully sorts $abcabc$. Choosing instead $\theta_{2}=(a)$, we have $\theta_{2}(abcabc)=(bc,bc)$.
2.1. The basis of foot-sortable sock orderings
Definition 17.
A minimally unsortable, or minimal, sock ordering $X$ is an unsortable sock ordering in which for all $Y\preceq X$, $Y$ is unsortable $\implies$ $|Y|=|X|$.
Observation 18.
Let $\Gamma$ be the set of all minimally unsortable sock orderings, and let $M$ be the basis of foot-sortable sock orderings, the smallest set of sock orderings which foot-sortable sock orderings must avoid. Then $M=\Gamma$.
As mentioned in the introduction, the basis of foot-sortable sock orderings is infinite. The following definition is used to describe our infinite classes in the basis $\Gamma$.
Definition 19.
Let $T:=(t_{0},\dotsc,t_{n})$ be a sequence of $n+1$ distinct colors. Define $I(t_{0}):=I(T)$ to be the empty string when $n=0$, and when $n\geq 1$, let
$I(T):=I(t_{0},\dotsc,t_{n}):=I(t_{0},\dotsc,t_{n-1})t_{n}t_{n-1}=t_{1}t_{0}t_{2}t_{1}\dotsc t_{n}t_{n-1}.$
Observation 20.
Let $n\geq 0$ and $a,b,t_{0},\dotsc,t_{n}$ be distinct colors. For every $0\leq i\leq n$, if $abt_{i}\subseteq S$ and $I(t_{i},\dotsc,t_{n})at_{n}\subseteq R$, then the state $(S,R)$ is unsortable. Furthermore, if $S^{\prime}\subseteq abt_{i}\subseteq S$ and $R^{\prime}\subseteq I(t_{i},\dotsc,t_{n})at_{n}\subseteq R$ are such that $S^{\prime}R^{\prime}\neq abt_{i}I(t_{i},\dotsc,t_{n})at_{n}$, then $(S^{\prime},R^{\prime})$ is sortable.
Proof.
The proof is by induction on $n-i$. For the base case, we have $i=n$, so
$I(t_{i},\dotsc,t_{n})at_{n}=at_{n}$
and it is not hard to see that both statements are true.
For the induction step, first note that sorting $t_{i}$ in
$(abt_{i},I(t_{i},\dotsc,t_{n})at_{n})=(abt_{i},t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n})$
leads to $(abt_{i+1},I(t_{i+1},\dotsc,t_{n})at_{n})$. Note that, when $n-i=1$ the only sortable color is $t_{i}$, so the first statement follows by induction. When $n-i\geq 2$, the only sortable colors are $t_{i}$ and $t_{i+1}$ and sorting $t_{i+1}$ in
$(abt_{i},I(t_{i},\dotsc,t_{n})at_{n})=(abt_{i},t_{i+1}t_{i}t_{i+2}t_{i+1}I(t_{i+2},\dotsc,t_{n})at_{n}),$
leads to $(abt_{i}t_{i}t_{i+2},I(t_{i+2},\dotsc,t_{n})at_{n})$ and the conclusion again follows by the induction hypothesis.
We will now prove the second statement when $n-i\geq 1$. Recall that we are assuming $S^{\prime}\subseteq abt_{i}$ and
$R^{\prime}\subseteq I(t_{i},\dotsc,t_{n})at_{n}=t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$
where at this one of the two is a strict subordering. By possibly adding socks, we also assume that either $S^{\prime}=abt_{i}$ or $R^{\prime}=t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$. First assume that $R^{\prime}=t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$, so $S^{\prime}\neq abt_{i}$. If $b\in S^{\prime}$ and $t_{i}\notin S^{\prime}$, then we sort $b$ and then $t_{i}$. Otherwise, we only sort $t_{i}$. In either case, we reach $(S^{\prime\prime},R^{\prime\prime})$ where $S^{\prime\prime}$ is a strict subordering of $abt_{t+1}$ and $R^{\prime\prime}=I(t_{i+1},\dotsc,t_{n})at_{n}$. The conclusion then follows from the induction hypothesis.
Now, assume $S^{\prime}=abt_{i}$. So, $R^{\prime}\neq t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$. If $R^{\prime}$ does not begin with $t_{i+1}t_{i}$, then we can sort $t_{i}$ and then $b$ to reach $(S^{\prime\prime},R^{\prime\prime})$ where $S^{\prime\prime}\subseteq at_{i+1}$ and $R^{\prime\prime}\subseteq I(t_{i+1},\dotsc,t_{n})at_{n}$. In this case, since $b\notin S^{\prime\prime}$, the conclusion follows from the induction hypothesis. Otherwise, we have $S^{\prime}=abt_{i}$ and $R^{\prime}=t_{i+1}t_{i}R^{\prime\prime}$ where $R^{\prime\prime}$ is a strict subordering of $I(t_{i+1},\dotsc,t_{n})at_{n}$. Then, after sorting the color $t_{i}$, we have $(abt_{i+1},R^{\prime\prime})$ and the conclusion follows by the induction hypothesis. ∎
We conclude this section by enumerating $\Gamma$, the basis of foot-sortable sock orderings. The following list consists of every sock ordering which is minimally unsortable. We describe five classes of minimally unsortable sock orderings which have infinitely many members, and we also outline other classes of sock orderings which are finite in number. We have divided the basis into categories that correspond with particular subpatterns we will later generate using a greedy sorting algorithm. Sock orderings in classes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ have color labels which denote the behavior of particular socks during the greedy sorting algorithm’s procedure.
We have the following orderings in the class $\mathcal{T}$, separated by the number of distinct colors in the sock ordering:
$\mathcal{T}_{1}\cong abacaba$ $\mathcal{T}_{2}\cong abacbab$ $\mathcal{T}_{3}\cong abcabca$ $\mathcal{T}_{4}\cong abcacba$
$\mathcal{T}_{5}\cong abcabac$ $\mathcal{T}_{6}\cong abcbacb$ $\mathcal{T}_{7}\cong abcbabc$
$\mathcal{T}_{8}\cong abcadcab$ $\mathcal{T}_{9}\cong abacdcba$ $\mathcal{T}_{10}\cong abacdcab$
Note that all of the sock orderings in the following class, $\mathcal{C}_{1}$, contain either $atdarat$, $atdrart$, or $atdarbrt$. Furthermore, sorting the color $d$ first for each of these three subpatterns yields $(at,arat)$, $(at,rart)$, and $(at,arbrt)$, respectively, and none of these three states are sortable. To then see that the orderings in $\mathcal{C}_{1}$ are unsortable, it is only necessary to verify that in each of these orderings, we must sort $d$ before the colors $a$, $b$, $r$, and $t$.
$\mathcal{C}_{1,1}\cong datdarat$ $\mathcal{C}_{1,6}\cong atdtrart$ $\mathcal{C}_{1,10}\cong atdtarbrt$
$\mathcal{C}_{1,2}\cong atdtarat$ $\mathcal{C}_{1,7}\cong datdrart$ $\mathcal{C}_{1,11}\cong adtdarbrt$
$\mathcal{C}_{1,3}\cong atdatrat$ $\mathcal{C}_{1,8}\cong adtdrart$ $\mathcal{C}_{1,12}\cong datdarbrt$
$\mathcal{C}_{1,4}\cong adtdarat$ $\mathcal{C}_{1,9}\cong atdedrart$ $\mathcal{C}_{1,13}\cong atdedarbrt$
$\mathcal{C}_{1,5}\cong atdedarat$
The other finite class is $\mathcal{C}_{2}$, consisting of those sock orderings which contain either the subordering $abtdabt$ or $abtdatb$. Since $(abt,at)$ is unsortable and $d$ must be sorted before the colors $a$, $b$ and $t$ in the following orderings, it is not difficult to verify that the the following orderings are minimally unsortable.
$\mathcal{C}_{2,1}\cong abtdatab$ $\mathcal{C}_{2,4}\cong adbtdatb$ $\mathcal{C}_{2,6}\cong abtdedabt$
$\mathcal{C}_{2,2}\cong abtdtatb$ $\mathcal{C}_{2,5}\cong abdtdatb$ $\mathcal{C}_{2,7}\cong abtdedatb$
$\mathcal{C}_{2,3}\cong adbtdabt$
All that remains are the five infinite classes of the basis. For every $n\geq 0$, let $a$, $b$, $d$, $e$, and $t_{0},\dotsc,t_{n}$ be distinct colors. Let $T:=(t_{0},\dotsc,t_{n})$ and define $t:=t_{0}$. Let $E:=I(T)at_{n}$ (so $E=at_{0}=at$ when $n=0$) and define
$I_{1,n}:=dabtdE$ $I_{2,n}:=adbtdbE$ $I_{3,n}:=abdtdbE$ $I_{4,n}:=abtdtbE$ $I_{5,n}:=abtdedbE$
With Observation 20 it is not hard to verify that these orderings are also minimally unsortable.
3. Algorithm
We now present a deterministic recursive algorithm (Algorithm 3) for sock sorting. On input $X$, the algorithm returns a triple $(\theta,(S,R),X^{\prime})$ where $X^{\prime}\subseteq X$ and $(S,R)=\theta(X^{\prime})$. If $X$ is sortable, then $X^{\prime}=X$ and $S=R=\emptyset$, so $\theta$ is a sorting sequence that fully sorts $X$. If $X$ is unsortable, then $X^{\prime}$ is an unsortable subordering of $X$ and $R\neq\emptyset$. Algorithm 3 closely follows the proof Theorem 1, so the proof of Theorem 1 also serves as a proof that Algorithm 3 is correct.
Algorithm 2 (what we call the greedy algorithm) is mentioned extensively in our proof. We often write $(S,R)=\phi_{G}(X)$ to mean that $(S,R)$ is the state when the greedy algorithm terminates. The greedy algorithm simply sorts the sortable color that terminates first in $SR$ until no color is sortable. In particular, if a color only appears on the top of the stack, then that color is sorted next.
At a high-level, the main idea of our sorting algorithm is as follows. Let $(S,R)=\phi_{G}(X)$ be the state after the greedy algorithm terminates. If we assume $X$ is sortable and the greedy algorithm fails, then we can always find a color $d$ such that any sorting sequence that sorts $X-d$ also sorts $X$. Therefore, we can recursively call our sorting routine on $X-d$ and use the resulting sequence to sort $X$. In most cases, this color $d$ is the last color sorted by the greedy algorithm. However, in some cases, we must backtrack in order to find such a color $d$.
For brevity, we do explicitly describe the subroutine Move which is referenced below. For a given state $(S,R)$ with sortable color $a$, the call $\textrm{Move}((S,R),a)$ returns the state $(S^{\prime},R^{\prime})$ after sorting the color $a$.
Algorithm 1 Sort using a prescribed sorting order, except we always immediately sort colors that only appear on the top of the stack. The routine returns the actual sequence of colors sorted.
1:function SortColors($X,\theta$)
2:     $(S,R)\leftarrow(\emptyset,X)$
3:     $\theta^{\prime}\leftarrow\emptyset$triangleright $\theta^{\prime}$ records the actual sorting sequence.
4:     while $R\neq\emptyset$ and $S\neq\emptyset$ do
5:         if $S\neq\emptyset$ and the top color $t$ on $S$ is not in $R$ then
6:              Append $t$ to $\theta^{\prime}$
7:              $(S,R)\leftarrow\textrm{Move}((S,R),t)$
8:         else if $\theta=\emptyset$ then return $\theta^{\prime},(S,R)$
9:         else
10:              $d\leftarrow$ first color in $\theta$
11:              Remove $d$ from $\theta$
12:              if $d$ is in $SR$ then
13:                  if $d$ is not sortable in $(S,R)$ then return $\theta^{\prime},(S,R)$
14:                  end if
15:                  Append $d$ to $\theta^{\prime}$
16:                  $(S,R)\leftarrow\textrm{Move}((S,R),d)$
17:              end if
18:         end if
19:     end while
20:     return $\theta^{\prime},(S,R)$
21:end function
Algorithm 2 When $C=\emptyset$ this is our greedy sorting algorithm. Otherwise, we assume $C\subseteq X$ and follow the greedy sorting algorithm, except we stop when a color in $C$ would be sorted next. We use the second parameter $C$ to backtrack.
1:function Greedy($X$, C)
2:     $(S,R)\leftarrow(\emptyset,X)$
3:     while $\exists d$ a sortable color in $(S,R)$ that terminates first in $SR$ do
4:         if $d\in C$ then return $\theta,(S,R)$
5:         end if
6:         $(S,R)\leftarrow$ $\textrm{Move}((S,R),d)$
7:         Append $d$ to $\theta$
8:     end while
9:     return $\theta,(S,R)$
10:end function
Algorithm 3 The sorting algorithm
1:function FullSort($X$)
2:     $\theta,(S,R)\leftarrow\textrm{Greedy}(X,\emptyset)$
3:     if $R=\emptyset$ or $S=\emptyset$ or $\exists x\in S$ s.t. $(x,R)$ is terminal then return $\theta,(S,R),X$
4:     end if
5:     Let $b$ and $t$ be the top two colors on $S$ and let $d$ be the last color in $\theta$
6:     if $(bt,S)$ is terminal then return $\textrm{SortHelper}(X,d,b)$
7:     end if
8:     $a\leftarrow$ the lowest color in $S$ s.t. $(abt,R)$ is terminal
9:     if $\exists z\in R$ s.t. either $z=b$ or $z$ is between $a$ and $b$ on $S$ then
10:         if $R$ contains $azt$ or $atz$ then return $\textrm{SortHelper}(X,d,a)$
11:         else return $\textrm{SortHelper}(X,d,z)$
12:         end if
13:     end if
14:     Let $A$ and $C$ be the strings that precede and follow the last $b$ in $X$ (so $X=AbC$ and $b\notin C$)
15:     $\theta,(S,R)\leftarrow\textrm{Greedy}(X,C)$triangleright Backtrack to before the final $b$ is pushed onto the stack
16:     $c\leftarrow$ first terminating sortable color in $(S,R)$triangleright $c$ would be sorted next greedily
17:     $d\ \leftarrow$ last color in $\theta$triangleright $d$ is the last color sorted
18:     if $ca\subseteq A$ then
19:         return $\textrm{Greedy}(X,\emptyset),X$triangleright $X$ is unsortable - return greedy output
20:     end if
21:     Let $y$ be the color immediately beneath the top color on the stack $S$
22:     return SortHelper(X, d, y)
23:end function
Algorithm 4 This is the helper routine for the sorting algorithm. When it is called, $X$, $a$, $x$, and $d$ must satisfy the hypothesis of Lemma 25. Furthermore, we also require that if Lemma 25(I) holds, then $X$ satisfies the hypothesis of either Lemma 27 or Lemma 28, so $X$ is unsortable. Note that when both Lemma 25(I) and (II) do not hold, if a sequence that fully sorts $X-d$ does not fully sort $X$, Lemma 25(IV) must hold and Lemma 25(IV) implies that $X$ is unsortable.
1:function SortHelper($X,d,a$)
2:     if $a$ is blocked by a $d$-sandwich in $X$ then return $\textrm{Greedy}(X,\emptyset),X$triangleright $X$ is unsortable
3:     end if
4:     $\theta,(S,R),X^{\prime}\leftarrow\textrm{FullSort}(X-d)$
5:     if  $R\neq\emptyset$  then return $\theta,(S,R),X^{\prime}$triangleright $X^{\prime}\subseteq X-d$ is unsortable
6:     else
7:         return $\textrm{SortColors}(X,\theta),X$triangleright $\theta$ fully sorts $X-d$, so $\theta$ fully sorts $X\iff X$ is sortable
8:     end if
9:end function
4. Prelimanary Observations and Lemmas
In the enumeration of our basis $\Gamma$, the five infinite classes have a common structure. In some sense, the following lemma explains the origin of this commonality.
Lemma 21.
Let $X$ be a sock ordering that, for some $n\geq 0$, contains distinct colors $t_{0},\dotsc,t_{n}$, $b$, and $c$ where the final $b$ precedes the final $c$ in $X$. Define $A$ and $C$ to be the strings such that $X=AbC$ where $b\notin C$ and $c\in C$. Let $\theta$ be a sorting sequence for $X$ that sorts a color in $C$ and let $c\in C$ be the color in $C$ that is sorted first by $\theta$. Taking $(S,R):=\theta(X)$, suppose $b\in S$ and that once $c$ is sorted, $b$ never appears as the top element on the stack. Let $t_{0},\dots,t_{n}$ (in this order) be the $n+1$ distinct colors which appear immediately above the color $b$ in the stack after $c$ is sorted. Then
$I(c,t_{0},\dotsc,t_{n})R\subseteq C.$
Proof.
We will prove the claim by induction on $n$.
For the base case, assume $n=0$, so $t_{0}=t_{n}$ is the unique color that appears directly above $b$ after $c$ is sorted. Since $c$ follows the last $b$ and $b$ can never appear only as the top color on the stack after $c$ is sorted, the color $t_{0}=t_{n}$ must appear above $b$ after $c$ is sorted. Therefore, we have
$C\supseteq t_{0}cR=I(c,t_{0},\dotsc,t_{n})R.$
Now, suppose $n\geq 1$. Let $d$ be the color that is sorted that placed the color $t_{n}$ directly above the color $b$ on the stack. And let $\theta_{d}$ be the initial sequence of $\theta$ up to but not including $d$. Let $(S^{\prime},R^{\prime}):=\theta_{d}(X)$. Note, during the execution of $\theta_{d}$, the color $b$ is never the top element of the stack and that $t_{0},\dotsc,t_{n-1}$ is the sequence of colors that appear immediately above the color $b$ after $c$ is sorted. Therefore, by induction, we have
$I(c,t_{0},\dotsc,t_{n-1})R^{\prime}\subseteq C.$
Recall that $t_{n-1}$ appears immediately above $b$ on the stack $S^{\prime}$, but after $d$ is sorted $t_{n}$ is the color above $b$ on the stack. Since $t_{n-1}\neq t_{n}$, this implies that $d=t_{n-1}$. Furthermore, we also have $t_{n}t_{n-1}R\subseteq R^{\prime}$. This completes the proof because
$C\supseteq I(c,t_{0},\dotsc,t_{n-1})R^{\prime}\supseteq I(c,t_{0},\dotsc,t_{n-1})t_{n}t_{n-1}R=I(c,t_{0},\dotsc,t_{n-1},t_{n})R.\qed$
We now present a few simple observations and lemmas that will be useful throughout.
Lemma 22.
Let $X$ be a sock ordering. Suppose that $x\in X$ is unsortable and that $x$ is one of the first two colors to appears in $X$. Then any color that appears after the final $x$ in $X$ is blocked by an $x$-sandwich.
Proof.
Let $yzy$ be a sandwich that blocks $x$ in $X$. Since $x$ is one of the first two colors to appear, $z$ cannot precede the first occurrence of $x$. Therefore, both $y$ and $z$ appear between the first and last occurrence of $x$. Therefore, any color that follows the final $x$ in $X$ is blocked by one of the sandwiches $xyx$ or $xzx$. ∎
Observation 23.
Let $X$ be a sock ordering that contains a sandwich. If $aba$ is the first sandwich to appear in $X$, then any color that is not $a$ or $b$ is either sortable in $X$ or is blocked by the sandwich $aba$.
Proof.
Let $x$ be an unsortable color in $X$ that is neither $a$ nor $b$. Because $x$ is unsortable, it must follow a sandwich. Since $aba$ is the first sandwich to appear, $x$ must follow $aba$. Therefore, $x$ is blocked by $aba$. ∎
Lemma 24.
Let $S$ and $R$ be strings where $S$ contains at least two distinct colors and does not contain a sandwich. Let $a\in S$ where $a$ is not the top color in $S$ and let $S^{\prime}$ be formed by removing the color $a$ from $S$. If $x$ is unsortable in $(S,R)$ and $x$ is sortable in $(S^{\prime},R)$, then
  • $x=a$, or
  • $ax\subseteq R$, or
  • there exists a color $y$ distinct from the colors $a$ and $x$ such that $yx\subseteq R$ and $y$, $a$, and $x$ are the top three colors on the stack $S$, respectively
Proof.
Assume $x\neq a$. Because $x$ is sortable in $(S^{\prime},R)$, we either have $x\in R$, or that $x$ is the top color on $S^{\prime}$. Suppose $x$ is the top color on $S^{\prime}$. Because $a$ is not the top color on $S$, this implies that $x$ is the top color on $S$. Since $x$ is not sortable in $(S,R)$, we have $x\in R$. So, in all cases, we have $x\in R$.
Let $yzy$ be the first sandwich in $SR$ that blocks $x$. Since $S$ does not contain $yzy$, we have $yx\subseteq R$. Because $x$ is sortable in $S^{\prime}R$, the color $a$ is either $y$ or $z$. If $a$ is $y$, then $ax\subseteq R$, so assume $a$ is $z$. Since $yzy=yay$ does not block $x$ in $S^{\prime}R$, we must have $ya\subseteq S$. Recall that, by assumption, $a$ is not the top color in $S$, so let $w$ be the color that appears directly above $a$ in $S$. If $w\neq x$, then $ywy$ blocks $x$ in $S^{\prime}R$, so we have $w=x$. Since $x$ is sortable in $(S^{\prime},R)$ and $x$ is on $S^{\prime}$, the color $x$ must be the top color on the stack $S^{\prime}$. Therefore, $x$ is the top color on $S$.
We now have that the colors $a$ and $x$ are the top two colors in $S$, respectively, and $yx\subseteq R$. Therefore, we only need to show that the color $y$ is directly below the color $a$ in $S$. This follows from the observation that if there exists $u$ that is distinct from $y$, $a$, or $x$ that is between $y$ and $a$ in $S$, then the sandwich $yuy$ blocks $x$ in $S^{\prime}R$, a contradiction. ∎
The following lemma establishes some general conditions that imply an unsortable sock ordering must either contain a proper subordering that is unsortable or contain $I_{1,n}$ for some $n\geq 0$. This lemma is a crucial piece of our recursive algorithm, because if (I) and (IV) do not hold, then we can determine if a sock ordering is sortable and how to sort it by considering a proper subordering.
Lemma 25.
Let $X$ be a sock ordering and let $a$, $x$, and $d$ be distinct colors in $X$. Let $\overline{\theta}$ be a nonempty sorting sequence for $X$ such that $\overline{\theta}(X)=(S,R)$, where $ax\subseteq S$, $a\in R$, and $d$ is the last color sorted in $\overline{\theta}$. Further suppose that there exists $S^{\prime}\subseteq S$ and $R^{\prime}\subseteq R$ such that $ax\subseteq S^{\prime}$ and $a$ is the only good sortable color in $S^{\prime}R^{\prime}$. One of the following is true:
  1. (I)
    there is a $d$-sandwich that blocks $a$ in $X$;
  2. (II)
    $X-d$ is unsortable;
  3. (III)
    $X-d$ is sortable and if a sequence $\theta$ fully sorts $X-d$, then $\theta$ also fully sorts $X$; or
  4. (IV)
    $I_{1,n}\preceq X$ for some $n\geq 0$.
Proof.
Assume that (I), (II), and (III) do not hold. We will show that (IV) must hold. Let $X=YDW$, where the first element of $D$ is the first $d$ in $X$ and the final element of $D$ is the final $d$ in $X$. Note that $R\subseteq W$, so $a\in W$. Because there are no $d$-sandwiches that block $a$ in $X$, the only colors that could appear in $D$ are $a$ and $d$. Since $d$ is sorted by $\overline{\theta}$ but $a$ and $x$ are not, there is no sandwich that consists only of $a$ and $x$ that blocks $d$ in $X$. Therefore, $D$ cannot contain $a$ as otherwise a sandwich of the form $axa$ would block $d$ in $X$. As such, $d$ is the only color that appears in $D$. Since $d$ is the last color in $\overline{\theta}$, this implies that $Y\supseteq S\supseteq S^{\prime}\supseteq ax$.
Since, by assumption, the statements (II) and (III) are both false, there exists a sorting sequence $\theta$ that fully sorts $X-d=YW$ such that $\theta$ does not fully sort $X=YDW$. Let $\theta_{a}$ be the list of initial colors in $\theta$ up to, but not including, $a$ (so $\theta_{a}$ is empty if and only if $\theta$ begins with $a$). Note that no color in $S^{\prime}R^{\prime}\subseteq X$ could appear before $a$ in $\theta$ because $a$ is the only good sortable color in $S^{\prime}R^{\prime}$. This means that no color that appears after $ax$ in $X$ is in $\theta_{a}$. To see this, note that sorting such a color would put $ax$ onto the stack, but that would make it impossible to sort $a$ before sorting $x$. In particular, this means that no color that appears in $W$ is in $\theta_{a}$. Therefore, $a$ is the color in $W$ that is sorted first by $\theta$ and $\theta_{a}$ is a sorting sequence for $X=YDW$. Furthermore, if we let $(S_{1},R^{\prime}_{1})=\theta_{a}(YW)$, then $(S_{1},R_{1})=\theta_{a}(YDW)$ where $R^{\prime}_{1}$ is formed by removing $d$ from $R_{1}$. Therefore, because $a$ is sortable in $(S_{1},R^{\prime}_{1})$ and no $d$-sandwich blocks $a$ in $X$, the color $a$ is also sortable in $(S_{1},R_{1})$. Recall that no color that appears after $ax$ is in $\theta_{a}$. This, with the fact that $axDa\subseteq SDR\subseteq X$, implies that $xDa\subseteq R_{1}$. Therefore, sorting $a$ in $(S_{1},R_{1})$ pushes $xD$ onto the stack.
Let $z$ be the first color in $\theta$ that causes a sandwich to appear in the stack when $\theta$ is applied to $X$. Such a color $z$ must exist because $\theta$ does not fully sort $X$. Let $\theta_{z}$ be the list of initial colors in $\theta$ up to but not including $z$. Note that by our previous arguments, $a$ must appear in $\theta_{z}$ and $a$ is the color in $W$ that is sorted first by $\theta_{z}$. Furthermore, since $xD$ is pushed onto the stack when $a$ is sorted and $d\notin\theta_{z}$, if we let $(S^{\prime}_{2},R_{2})=\theta_{z}(YW)$ we have $(S_{2},R_{2})=\theta_{z}(YDW)$ where $xD\subseteq S_{2}$ and $S^{\prime}_{2}=S_{2}-d$. Note that while applying $\theta_{z}$ to $X$, after $a$ is sorted the color $d$ can never appear as the top element of the stack since $\theta$ fully sorts $X-d$ but does not not fully sort $X$. Let $t_{0},\dotsc,t_{n}$ be the $n+1$ colors that appear immediately above $d$ on the stack after $a$ is sorted by $\theta_{z}$. Since $X=YDW$ and $D$ consists solely of one or more copies of $d$, Lemma 21 implies that $dI(a,t_{0},\dotsc,t_{n})R_{2}\subseteq dW$. Recall that $xD\subseteq S_{2}$. Therefore, there exists a color $y$ that appears directly below $D$ in $S_{2}$. (We could have $y=x$.) Since $axDa\subseteq X$, we have $ayDa\subseteq X$, so $aydI(a,t_{0},\dotsc,t_{n})R_{2}\subseteq YdW\subseteq X$.
Note that, by the selection of $z$, the color $z$ is unsortable in $(S_{2},R_{2})$, but $z$ is sortable in $(S^{\prime}_{2},R_{2})$. Since $d\notin S^{\prime}_{2}R_{2}\subseteq X-d$, we have $z\neq d$ and $dz\nsubseteq R_{2}$. Therefore, Lemma 24 implies that $y$, $d$, and $z$ are the top three colors in the stack $S_{2}$ and that $yz\subseteq R_{2}$. Note that this means that $z=t_{n}$ since $t_{n}$ is defined to be the last color that appears directly above $d$ on the stack when sorting $X$ with $\theta_{z}$. Therefore,
$X\supseteq aydI(a,t_{0},\dotsc,t_{n})R_{2}\supseteq aydt_{0}aI(t_{0},\dotsc,t_{n})yz=aydt_{0}aI(t_{0},\dotsc,t_{n})yt_{n}\cong I_{1,n}.\qed$
The following three lemmas are needed in the proof of our main theorem to show that a given unsortable sock ordering has a subpattern that is in $\Gamma$. Unsurprisingly, these proofs have a number of cases and are somewhat technical, so we defer their proofs until Section 6. In Lemma 27 and 28, we only need the first conclusion “$X$ is unsortable”, to prove that the algorithm is correct. The second conclusion “a subpattern of $X$ is in $\Gamma$” is needed in our inductive proof of Theorem 1.
Lemma 26.
If $X$ is trivially unsortable, then a subpattern of $X$ is in $\Gamma$.
Lemma 27.
Let $X$ be a sock ordering where, for some $n\geq 0$, the distinct colors $a,b,d,t_{0},\dotsc,t_{n}$ appear in $X$ and define $t:=t_{0}$. Suppose that a nonempty sorting sequence produces $(S,R)$ when it is applied to $X$ and let $d$ be the last color sorted. If $abt\subseteq S$ and either
  • $atb$ or $abt$ is contained in $R$ and $a$ is blocked by a $d$-sandwich in $X$; or
  • $bI(t_{0},\dotsc,t_{n})at_{n}\subseteq R$ and $b$ is blocked by a $d$-sandwich in $X$,
then $X$ is unsortable. If, in addition, every strict subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
Lemma 28.
Let $X$ be a sock ordering and suppose that a nonempty sorting sequence produces $(S,R)$ when it is applied to $X$, where $d$ is the last color sorted by the sequence. Further suppose that the following two conditions hold:
  • For every $x\in S$, the state $(x,R)$ is not terminal.
  • There exist distinct colors $a$ and $t$ in $S$ such that $(at,R)$ is terminal.
Then there exists $P\subseteq R$ such that $a$ is the only good sortable color in $atP$. Furthermore, if $a$ is blocked by a $d$-sandwich in $X$, then $X$ is unsortable. If, in addition, every strict subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
The following lemma shows that if we reach the state $(S,R)$ that is terminal and $S$ has at least three colors, then, in a certain sense, we only need to focus on at most three colors in $S$. Furthermore, two of the three colors can be the top two colors on $S$ (cf. Algorithm 3, Line 8)
Lemma 29.
Let $S$ and $R$ be sock orderings where $S$ does not contain a sandwich and $S$ contains at least three colors. Suppose $(S,R)$ is terminal. If $b$ and $t$ are the top two colors on $S$, then there exists a color $a\in S$ distinct from $b$ and $t$ such that $(abt,S)$ is terminal.
Proof.
Suppose $(S,R)$ is a counterexample. We can assume without loss of generality that no colors are repeated in $S$. So we can write $S=x_{1}x_{2}\dotsm x_{m}bt$ where $x_{1},\dotsc,x_{m}$, $b$, and $t$ are distinct colors. Let $p$ be as small as possible such that there exists $y_{1},\dotsc,y_{p}$, a subsequence of $x_{1},\dotsc,x_{m}$, such that $(y_{1}y_{2}\dotsm y_{p}bt,R)$ is terminal. Since $(S,R)$ is a counterexample, we have $p\geq 2$, and if we let $W$ be the nonempty string $y_{3}\dotsm y_{p}b$, then for $i\in\{1,2\}$, there exists some $z_{i}$ that is sortable in $(y_{i}Wt,R)$.
For $i\in\{1,2\}$, since $z_{i}$ is sortable in $(y_{i}Wt,R)$, we have $z_{i}\notin y_{i}W$ and $y_{i}z_{i}\nsubseteq R$. Since $z_{1}\neq y_{1}$, $y_{1}z_{1}\nsubseteq R$, and $z_{1}$ is not sortable in $(y_{1}y_{2}Wt,R)$, Lemma 24 implies that $z_{1}$ is unsortable in $(y_{2}Wt,R)$. Therefore, because $y_{1}z_{1}\nsubseteq R$ and $z_{1}\notin W$, either $z_{1}=y_{2}$ or there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ that does not contain $y_{1}$. If $z_{1}=y_{2}$, then, because $z_{1}=y_{2}$ is sortable in $(y_{1}Wt,R)$, we have $y_{2}\in R$, so $y_{1}y_{2}=y_{1}z_{1}\nsubseteq R$ implies that $y_{1}$ is not sortable in $(y_{2}Wt,R)$. If there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ that does not contain $y_{1}$, then $y_{1}z_{1}\nsubseteq R$ implies that $y_{1}$ is not sortable in $(y_{2}Wt,R)$. So, in both cases, $y_{1}$ is not sortable in $(y_{2}Wt,R)$, so $y_{1}\neq z_{2}$, and Lemma 24 then implies that $y_{1}z_{2}\subseteq R$. Since $y_{1}z_{2}\subseteq R$ and $y_{2}z_{2}\nsubseteq R$, $y_{2}$ is not sortable in $(y_{1}Wt,R)$. Therefore, $y_{2}\neq z_{1}$. This together with Lemma 24 implies that $y_{2}z_{1}\subseteq R$. Since $y_{1}z_{2}\subseteq R$, $y_{2}z_{2}\nsubseteq R$, and $y_{2}z_{1}\subseteq R$, we have $y_{1}z_{1}\subseteq R$. But we have already established that $y_{1}z_{1}\nsubseteq R$. ∎
Our last lemma is simple despite the somewhat technical statement. It is used in the case when a color, say $x$, is not sorted by the greedy algorithm despite terminating before the color which is sorted by the greedy algorithm, say $y$. Note that, in this case, if a sandwich blocks $x$, then it must be a $y$-sandwich because $y$ is sortable. The lemma deals with the other possibility. Namely, $x$ already being beneath the top color on the stack.
Lemma 30.
Let $S$ and $R$ be sock orderings where $S$ does not contain a sandwich and let $x$ and $y$ be distinct colors in $R$ such that the final $x$ in $R$ appears before the final $y$ in $R$. Suppose that, with respect to $(S,R)$, the color $y$ is sortable and the color $x$ is unsortable. If no $y$-sandwich blocks $x$ in $SR$, then $x$ and $y$ are the top two colors on $S$. Furthermore, no color except possibly $y$ precedes $x$ on $R$.
Proof.
Because $y$ is sortable in $SR$, every sandwich in $SR$ that appears before the final $y$ in $R$ must be a $y$-sandwich. Therefore, because the final $x$ in $R$ precedes the final $y$ in $R$ and $x$ is not blocked by a $y$-sandwich in $SR$, the color $x$ is not blocked by a sandwich in $SR$. The fact that $x$ is unsortable then implies that $x$ is below the top element on the stack $S$. Since $y$ is sortable in $SR$, there is no $x$-sandwich that blocks $y$ in $SR$. Therefore, the colors $x$ and $y$ are the top two colors on $S$. Furthermore, no color except possibly $y$ can precede $x$ on $R$. ∎
5. Proof of Theorem 1
The proof is by induction on $|X|$. The base case is vacuously true. So, by the induction hypothesis, we can assume that there does not exist a proper subordering of $X$ that is unsortable. Note that if we have colors $a$, $x$ and $d$ and substrings $S^{\prime}\subseteq S\subseteq X$ and $R^{\prime}\subseteq R\subseteq X$ that satisfy the hypothesis of Lemma 25, then we can assume (I) holds. That is, we can assume that $a$ is blocked by a $d$-sandwich. This is because (II) and (III) in Lemma 25 cannot hold, and if (IV) holds, then we satisfy the conclusion. This will make it easy to apply either Lemma 27 or 28 to complete the proof.
Let $(S,R):=\phi_{G}(X)$ be the terminal state of the greedy algorithm and let $d$ be the last color sorted by $\phi_{G}$. Since $X$ is unsortable, we have $R\neq\emptyset$.
Case 1: $S=\emptyset$ or there exists $x\in S$ such that $(x,S)$ is unsortable. If $S=\emptyset$, then Lemma 26 implies that a subpattern of $R\subseteq X$ is in $\Gamma$. Similarly, if there exists $x\in S$ such that $(x,R)$ is terminal, then $xR$ is trivially unsortable, and Lemma 26 implies that a subpattern of $xR\subseteq X$ is in $\Gamma$.
From now on, we will assume that we are not in Case 1. That is, for every $x\in S$, some color in $xR$ is sortable. Since $(S,R)$ is terminal, this implies that there are at least two colors on $S$. Define $t$ to be the top color in $S$, and let $b$ be the color in $S$ appearing immediately under $t$.
Case 2: $(bt,R)$ is terminal. By Lemma 28, there exists $P\subseteq R$ such that $b$ is the only good sortable color in $btP$ and that if $b$ is blocked by a $d$-sandwich in $X$, then a subpattern of $X$ is in $\Gamma$. If $b$ is not blocked by a $d$-sandwich in $X$, then we must satisfy Lemma 25(IV). That is, $I_{1,n}\preceq X$, for some $n\geq 0$.
Case 3: $(bt,R)$ is not terminal. With Lemma 29, we can define $a$ to be the lowest color in $S$ such that $(abt,R)$ is terminal.
Claim 31.
$at\subseteq R$.
Proof.
Suppose, by way of contradiction, that $at\nsubseteq R$. Then, because $t$ is unsortable in $(abt,R)$, Lemma 24 implies that $t$ is unsortable in $(bt,R)$. Since $(bt,R)$ is not terminal and both $b$ and $t$ are not sortable in $(bt,R)$, there must exist a color $y\in R$ that is not $b$ or $t$ which is sortable in $(bt,R)$. If $y=a$, then, because $at\nsubseteq R$, the color $y=a$ must follow the last occurrence of $t$ in $btR$. If $y\neq a$, then, since $y$ is sortable in $(bt,R)$ and unsortable in $(abt,R)$, Lemma 24 implies that $ay\subseteq R$. Because $at\nsubseteq R$, this implies that $y$ follows the last occurrence of $t$ in $btR$. So, in both cases, $y$ follows the last occurrence of $t$ in $btR$. This with Lemma 22 and the fact that $t$ is the second color in $btR$ and $t$ is unsortable in $btR$ give us that $y$ is unsortable in $btR$, a contradiction. ∎
Case 3.1: There exists a color $z\in S$ distinct from $a$ and $t$ such that $azt\subseteq S$ and $z\in R$. By Claim 31, we can assume that either $atz$, $azt$, or $zat=zaI(t)t$ is in $R$. But then Lemma 25 and Lemma 27 lead to a contradiction, because $a$ is the only good sortable color in both $aztatz$ and $aztazt$, and $z$ is the only good sortable color in $aztzat$.
Case 3.2: For every $z\in S$ distinct from $a$ and $t$ such that $azt\subseteq S$, we have $z\notin R$. Recall that $abt\subseteq S$. This and the case, implies that $b\notin R$. Let $A$ and $C$ be such that $X=AbC$ and $b\notin C$. Because $b\notin R$ there exists a color in $C$ that is sorted during $\phi_{G}(X)$. Let $c\in C$ be the color in $C$ that is sorted first by the greedy algorithm and let $(S^{\prime},A^{\prime}bC)$ be the state immediately before $c$ is sorted by $\phi_{G}(X)$. Note that we have
$S^{\prime}A^{\prime}\subseteq A.$ (1)
Recall that the color $b$ is never sorted by the greedy algorithm and note that, after $c$ is sorted, the color $b$ only appears on the stack. Therefore, because $ab\subseteq S$, we have
$a\in S^{\prime}A^{\prime}$ (2)
Furthermore, by the definition of the greedy algorithm, the color $b$ can never appear as the top element of the stack after $c$ is sorted by the greedy algorithm. For $n\geq 0$, let $t_{0},\dotsc,t_{n}$ be the $n+1$ distinct colors that appear immediately above $b$ after $c$ is sorted by the greedy algorithm. Note that then $t=t_{n}$, so Claim 31 implies that $at_{n}=at\subseteq R$. With Lemma 21, we have
$C\supseteq I(c,t_{0},\dotsc,t_{n})R\supseteq I(c,t_{0},\dotsc,t_{n})at_{n}.$ (3)
Note that if $ca\subseteq A$, then, with (3), we have
$X=AbC\supseteq cabI(c,t_{0},\dotsc,t_{n})at_{n}=cabt_{0}cI(t_{0},\dotsc,t_{n})at_{n}\cong I_{1,n}.$
So, for the remainder of the proof, we will assume
$ca\nsubseteq A$ (4)
Claim 32.
There are at least three colors in $S^{\prime}$. More specifically, if $y$ is the color immediately beneath the top color in $S^{\prime}$, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC$.
Proof.
First assume that $b$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$. Then, because the greedy algorithm sorted $c$ and not $b$ at stage $(S^{\prime},A^{\prime}bC)$ and the final $b$ occurs before the final $c$ in $S^{\prime}A^{\prime}bC$, with respect to $(S^{\prime},A^{\prime}bC)$, the color $b$ is unsortable and the color $c$ is sortable. Then, Lemma 30 implies that $b$ and $c$ are the top two colors on the stack $S^{\prime}$, so $y=b$. Since (1), (2), and (4), all hold and $b$ and $c$ are the top two colors on the stack $S^{\prime}$, we must have $ayc=abc\subseteq S^{\prime}$. By (3), we also have
$yI(c,t_{0},\dotsc,t_{n})at_{n}=bI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC.$
Now assume that $b$ is blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$. Let $\hat{y}$ be the other color in the first $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$; we will show that $\hat{y}=y$ and our conclusion holds. To see this, first note that we must have $\hat{y}c\subseteq S^{\prime}A^{\prime}\subseteq A$ and $c\hat{y}\subseteq S^{\prime}A^{\prime}\subseteq A$. Furthermore, (2) and (4) and $c\hat{y}\subseteq S^{\prime}A^{\prime}$ imply that $\hat{y}\neq a$ and
$a\hat{y}\subseteq S^{\prime}A^{\prime}.$ (5)
If the final $\hat{y}$ follows the final $c$ in $S^{\prime}A^{\prime}bC$, then, after $c$ is sorted, the color $\hat{y}$ must appear off the stack and, with (5), we will also have $a\hat{y}b$ on the stack. Then, the fact that $b$ is never sorted by the greedy algorithm implies that $\hat{y}$ is never sorted by the greedy algorithm. Furthermore, since we cannot have $\hat{y}$-sandwich on the stack, we must have $a\hat{y}bt_{n}\subseteq S$ and $\hat{y}\in R$. But this contradicts the fact that we are in Case 3.2.
Therefore, the final $\hat{y}$ precedes the final $c$ in $S^{\prime}A^{\prime}bC$. First note that since $c$ is sortable with respect to $(S^{\prime},A^{\prime}bC)$, the color $c$ is not below the top color on the stack $S^{\prime}$. Since there is a sandwich consisting of $\hat{y}$ and $c$ in $S^{\prime}AbC$, we must have $\hat{y}$ in $A^{\prime}bC$. Also, by the definition of $c$, we have $c\in C\subseteq A^{\prime}bC$. Since the greedy algorithm sorts $c$ and not $\hat{y}$ when $(S^{\prime},A^{\prime}bC)$ appears, $c$ is sortable and $\hat{y}$ is unsortable with respect to $(S^{\prime},A^{\prime}bC)$. Furthermore, since the selection of $\hat{y}$ implies that the color $\hat{y}$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$, Lemma 30 implies that $\hat{y}$ and $c$ are the top two colors in $S^{\prime}$. Note that we have shown that $\hat{y}=y$. Furthermore, Lemma 30 implies that and no color except possibly $c$ precedes $\hat{y}=y$ in $A^{\prime}bC$. In particular, $b$ does not precede $\hat{y}=y$ in $A^{\prime}bC$, so $\hat{y}=y\in A^{\prime}$. Then, with (3),
$A^{\prime}bC\supseteq ybI(c,t_{0},\dotsc,t_{n})at_{n}\supseteq yI(c,t_{0},\dotsc,t_{n})at_{n}.$
Since (1), (2), and (4), all hold and $y$ and $c$ are the top two colors in $S^{\prime}$, we have $ayc\subseteq S^{\prime}$. ∎
Defined $d^{\prime}$ to be the color sorted by the greedy algorithm immediately before reaching the state $(S^{\prime},A^{\prime}bC)$. By Claim 32, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,)at_{n}\subseteq A^{\prime}bC$. Note that, with Observation 20, it is not hard to see that $y$ is the only good sortable color in $aycyI(c,t_{0},\dotsc,)at_{n}$. If $y$ is not blocked by a $d^{\prime}$-sandwich in $X$, then Lemma 25 implies that a subpattern of $X$ is in $\Gamma$. Otherwise, Lemma 27 implies that a subpattern of $X$ is in $\Gamma$.
6. Proofs of Lemmas 26, 27, and 28
See 26
Proof.
Assume, for a contradiction, that $X$ is a minimal counterexample. Let $aba$ be the first sandwich in $X$. Recall that since $X$ is trivially unsortable, every color is blocked by a sandwich. That is, for all $z\in X$, there exist distinct colors $x,y$ in $X$ which are distinct from $z$ such that $xyxz\subseteq X$. This implies that every color must follow the first sandwich $aba$. Let $z$ be the first of the colors $a$ and $b$ that terminates in $X$. Let $xyx$ be the first sandwich in $X$ that blocks $z$.
Case 1: $x=a$.
Then $xyx=aya$ blocks $z$, so $z=b$. Suppose the $y$ in the sandwich $aya$ follows the first sandwich $aba$, then $X$ contains $abayaba\cong abacaba\cong\mathcal{T}_{1}$. Otherwise, because $aba$ (and not $aya$) is the first sandwich, $X$ must contain $abya$ which is then followed by $ba$ and the color $y$. Therefore, $X$ contains either $abyabya\cong abcabca\cong\mathcal{T}_{3}$, $abyayba\cong abcacba\cong\mathcal{T}_{4}$, or $abyabay\cong abcabac\cong\mathcal{T}_{5}$.
Case 2: $x=b$.
Then $xyx=byb$ blocks $z$, so $z=a$. Suppose the $y$ in the sandwich $byb$ follows the first sandwich $aba$. Then $X$ contains $abaybab\cong abacbab\cong\mathcal{T}_{2}$. Otherwise, $X$ contains $abya$. Since $byb$ blocks $z=a$ and the final $b$ follows the final $z=a$, the pattern $bab$ must follow this copy of $abya$. Furthermore, since $y$ must follow the first sandwich, the color $y$ must follow this $abya$. So, as in the previous case, $abya$ is followed by $ba$ and the color $y$. Therefore, as before, either $\mathcal{T}_{3}$, $\mathcal{T}_{4}$, or $\mathcal{T}_{5}$ is a subpattern of $X$.
Case 3: $x\neq a$ and $x\neq b$.
Case 3.1: The first $x$ appears before the first sandwich $aba$.
Then, because $aba$ is the first sandwich to appear, the color $x$ must also follow $aba$. Therefore, $X$ contains either $xabaxab\cong abcbabc\cong\mathcal{T}_{6}$ or $xabaxba\cong abcbacb\cong\mathcal{T}_{7}$.
Case 3.2: The first $x$ appears between the first and last $a$ in the first sandwich $aba$.
Note that $X$ must contain either $abxayx$ or $abxyax$ because $aba$ (and not $axa$ or $xyx$) is the first sandwich in $X$. Furthermore, we can assume that both $aba$ and $xyx$ in this subordering are both the first such subordering to appear in $X$. If $z=b$, then $X$ contains $abxaxba\cong abcacba\cong\mathcal{T}_{4}$. So assume $z=a$. If $y=b$, then $X$ must contain $abxabxab$ or $abxbaxab$, but then $bxb$ precedes $xbx$, so $xbx=xyx$ is not the first sandwich which blocks $z=a$. Therefore, we can assume $y\neq b$. As such, $X$ contains either $abxayxab\cong abcadcab\cong\mathcal{T}_{8}$ or $abxyax$, so we assume $X$ contains $abxyax$. Recall that every color must follow $abxya$, so every color that is not $a$ is blocked by either $axa$ or $aya$. Furthermore, $a$ is blocked by $xyx$. Therefore every color is blocked by one of the three sandwiches $axa$, $aya$, or $xyx$. Since none of these three sandwiches contains $b$, if we remove the color $b$ from $X$ we have a trivially unsortable sock ordering which contradicts the minimality of $X$.
Case 3.2: The first $x$ appears after the first sandwich $aba$.
So, we have $abaxyx$ in $X$ were $aba$ and $xyx$ in this subordering are the first such subordering to appear in $X$. If $y=a$, then we must have $z=b$, but then $axy=axa$ appears before $xyx$, a contradiction to the fact that $xyx$ is the first sandwich that blocks $z=b$. Similarly, if $y=b$, then we must have $z=a$, but then $bxy=bxb$ blocks $z=a$ and appears before $xyx$. Therefore, the colors $x$ and $y$ are distinct from the colors $a$ and $b$. Hence, $X$ either contains $abaxyxba\cong abacdcba\cong\mathcal{T}_{9}$ or $abaxyxab\cong abacdcab\cong\mathcal{T}_{10}$. ∎
See 27
Proof.
First note that if a strict subordering of $X$ is unsortable, then $X$ is unsortable. Therefore, to prove the lemma, we can assume that every strict subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$, since this will imply that $X$ is unsortable. Assume for a contradiction that no subpattern of $X$ is in $\Gamma$. Let $T:=(t_{0},\dotsc,t_{n})$ and $P\in\{atb,abt,bI(T)at_{n}\}$ such that $P\subseteq R$. We may write $X=S^{\prime}dR^{\prime}$ where $S\subseteq S^{\prime}$ and $d\notin R^{\prime}$ and $R\subseteq R^{\prime}$.
Note that $a$ is the only sortable color in $abtatb$ and $abtabt$. Furthermore, $b$ is the only good sortable color in $abtbI(T)at_{n}$ (when $n=0$ the color $b$ is the only sortable color in $abtbI(T)at_{n}=abtbat$ and when $n\geq 1$, $b$ and $t$ are the only sortable colors in $abtbI(T)at_{n}$, but sorting $t=t_{0}$ leaves $(abbt_{1},I(t_{1},\dotsc,t_{n})at_{n})$ which is unsortable by Observation 20).
  • If $P\in\{atb,abt\}$, then let $z=a$ and $y=b$.
  • If $P=bI(T)at_{n}$, then let $z=b$ and $y=a$.
We have the following.
Claim 33.
The color $z$ is blocked by a $d$-sandwich in $X$ and the color $z$ is not blocked by a sandwich consisting only of colors in $P$.
Proof.
The first conjunct is given by the hypothesis of the lemma. For the second conjunct, recall that, by the above discussion, $z$ is the only good sortable color in $abtP$. Therefore, because $X-d$ is sortable by assumption, $z$ is not blocked by a sandwich consisting only of colors in $P$. ∎
Claim 34.
One of the following holds.
  1. (i)
    $S^{\prime}$ contains one of $dabt$, $adbt$, $abdt$, or for some color $e\notin P$, $S^{\prime}$ contains $abtde$; or
  2. (ii)
    $R^{\prime}$ contains either $atab$ or $tP$, or
  3. (iii)
    for some color $e\notin P$, $e\in S^{\prime}$ and $R^{\prime}$ contains $eP$.
Proof.
Recall that by Claim 33, $z$ is blocked by a $d$-sandwich in $X$. That is, for some color $e$, we either have a sandwich of the form $ded$ which blocks $z$ or a sandwich of the form $ede$ which blocks $z$.
Case 1: A sandwich of the form $ded$ blocks $z$. In this case, we will show that (i) holds, so suppose $dabt$, $adbt$, and $abdt$ are not contained in $S^{\prime}$. Then, because $d\notin R^{\prime}$ implies $de\subseteq S^{\prime}$, we have $abtde\subseteq S^{\prime}$. To show that (i) holds, we then only need to show that $e\notin P$. Since $ded$ blocks $z$, we have $e\neq z$. We also have $e\neq y$, because, by Claim 33 a sandwich of the form $yty$ cannot block $z$ in $X=S^{\prime}dR^{\prime}$. Furthermore, $e\neq t$, since we are assuming $abdt\subseteq abtdt$ is not in $S^{\prime}$. So, we are left to consider the case when $n\geq 1$, $bI(T)at_{n}$ is in $R$, and $e=t_{i}$ for some $1\leq i\leq n$. Then
$X\supseteq abtdt_{i}dbI(T)at_{n}\supseteq btdt_{i}btt_{i}\cong dabtdat\cong I_{1,0},$
a contradiction.
Case 2: A sandwich of the form $ede$ blocks $z$. Since $d\notin R^{\prime}$, we must have $e\in S^{\prime}$. If $ede\subseteq S^{\prime}$, then a sandwich of the form $ded$ blocks $z$, so we are in the first case. Therefore, we can assume that $ez\subseteq R^{\prime}$.
By Claim 33, the color $z$ is not blocked by a $yty$ sandwich. This, with the fact that $abt\subseteq S\subseteq S^{\prime}$ and $ez\subseteq R^{\prime}$, yields $e\neq y$. Since $ede$ blocks $z$, we have $e\neq z$. Therefore, $e\notin\{a,b\}$.
Case 2.1: $P\in\{atb,abt,bat=bI(t_{0})at_{0}\}$. Recall that Claim 33 implies that a $yty$ sandwich does not block $z$ in $X$. Since $yt\subseteq abt\subseteq S$, $z$ cannot follow the $y$ in $P$. This, with the fact that $ez\subseteq R^{\prime}$, implies that we have $eP\subseteq R^{\prime}$ when $P\in\{abt,bI(t)at=bat\}$. Furthermore, if $P=atb$ and $eP\nsubseteq R^{\prime}$, then, regardless of $e$, we must have $atab\subseteq R^{\prime}$.
First note that if $e=t$, then condition (ii) is satisfied, so assume $e\neq t$. Recall that we have already shown that $e\notin\{a,b\}$, so, since we are in Case 2.1, we have $e\notin P$. Recall that $eP\subseteq R^{\prime}$ when $P\in\{abt,bI(t)at=bat\}$. Therefore, to show that (iii) is satisfied, we only need to show that $eP\subseteq R^{\prime}$ when $P=atb$. For a contradiction, assume $P=atb$ and $eP\nsubseteq R^{\prime}$. So $z=a$ and $y=b$. Recall that we have previously shown that $ea=ez\subseteq R^{\prime}$ and that $z=a$ cannot follow the $y=b$ in $P$. Furthermore, because $eP\nsubseteq R^{\prime}$, we have $atab\subseteq R^{\prime}$. Therefore, we either have $aetab\subseteq R^{\prime}$ or $ateab\subseteq R^{\prime}$. Since $e$ and $t$ are both in $S^{\prime}$, a sandwich consisting of the colors $e$ and $t$ blocks both $a$ and $b$ in $X=S^{\prime}dR^{\prime}$. Since $abt\subseteq S^{\prime}$, a sandwich of the form $ata$ blocks $e$ in $X=S^{\prime}dR^{\prime}$ and a sandwich of the form $aba$ blocks $t$ in $X=S^{\prime}dR^{\prime}$. Therefore, if we consider the pattern $X^{\prime}$ formed by removing all colors that are not $a$, $b$, $t$, or $e$ from $X$, we have that $X^{\prime}$ is a subordering of $X-d$ that is trivially unsortable. This contradicts the fact that every strict subordering of $X$ is sortable.
Case 2.2: $n\geq 1$ and $P=bI(T)at_{n}=bt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$.
In this case, we have $z=b$ and $y=a$. The color $b$ cannot follow $t_{1}t$ in $R^{\prime}$ because Claim 33 implies that $b$ is not blocked by a $tt_{1}t$ sandwich. Therefore, because $P=bt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$ and $eb\subseteq R^{\prime}$, if we do not have $eP\subseteq R^{\prime}$, then, regardless of $e$, we must have $bt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\subseteq R^{\prime}$. So, we have the following two cases.
Case 2.2.1: $R^{\prime}$ contains $eP=ebt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$. Note that if $e=t$, then (ii) holds and if $e\notin P$, then (iii) holds. Since we have already shown that $e\notin\{a,b\}$, we can assume $e=t_{i}$ where $1\leq i\leq n$. Recall that $t_{i}=e\in S^{\prime}$, so since Claim 33 implies that no $t_{i}tt_{i}$ sandwich blocks $b$ in $X$, we must have $abtt_{i}\subseteq S^{\prime}$. Therefore,
$X\supseteq abtt_{i}dt_{i}bt_{1}tI(t_{1},\dotsc t_{n})at_{n}\supseteq btdt_{i}btt_{i}\cong dabtdat\cong I_{1,0},$
a contradiction.
Case 2.2.2: $R^{\prime}$ contains $bt_{1}btI(t_{1},\dotsc,t_{n})at_{n}$. In this case, both when $n=1$ and when $n\geq 2$, we have
$X\supseteq abtdR\subseteq abtdbt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\supseteq btdt_{1}btt_{1}\cong dabtdat\cong I_{1,0},$
a contradiction. ∎
Suppose Claim 34(i) holds. If $dabt\subseteq S^{\prime}$, then we have $at\subseteq P$ or we have $n\geq 1$ and $I(T)at_{n}\subseteq P$. This means that $X$ contains $dabtdat\cong I_{1,0}$ or $dabtdI(T)at_{n}\cong I_{1,n}$. Otherwise, $X$ contains one of the patterns in the following table.
$adbtdatb\cong\mathcal{C}_{2,4}$ $adbtdabt\cong\mathcal{C}_{2,3}$ $adbtdbI(T)at_{n}\cong I_{2,n}$
$abdtdatb\cong\mathcal{C}_{2,5}$ $abdtdabt\supseteq abdtabt\cong I_{1,0}$ $abdtdbI(T)at_{n}\cong I_{3,n}$
$abtdedatb\cong\mathcal{C}_{2,7}$ $abtdedabt\cong\mathcal{C}_{2,6}$ $abtdedbI(T)at_{n}\cong I_{5,n}$
If Claim 34(iii) holds, then by swapping the labels of $d$ and $e$, we can see that $X$ contains either $I_{1,n}$ or one of the patterns from the table above.
Finally, if Claim 34(ii) holds, then $X$ must contain one of the four patterns in the following table.
$abtdatab\cong\mathcal{C}_{2,1}$ $abtdtatb\cong\mathcal{C}_{2,2}$ $abtdtabt\supseteq abdtabt\cong I_{1,0}$ $abtdtbI(T)at_{n}\cong I_{4,n}$
See 28
Proof.
As in the proof of Lemma 27, to prove the lemma, we can assume that every strict subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$. For a contradiction, assume that no subpattern of $X$ is in $\Gamma$. We may write $X=S^{\prime}dR^{\prime}$ where $S\subseteq S^{\prime}$ and $d\notin R^{\prime}$ and $R\subseteq R^{\prime}$. Since $(at,R)$ is terminal, there exists a sandwich that blocks $t$ in $atR$. Therefore, there exists a color that is neither $a$ or $t$ in $R$. Let $r$ be the first color that is neither $a$ nor $t$ in $R$. Furthermore, since a color that is neither $a$ nor $t$ must precede $t$ in $R$, we have $rt\subseteq R$.
Claim 35.
$ar\subseteq R$ and $at\subseteq R$
Proof.
Let $x\in\{r,t\}$ and suppose, for a contradiction, that $ax\nsubseteq R$. Note that then $x$ is one of the first two colors to appear in $tR$. Lemma 24 implies that $x$ is unsortable in $(t,R)$, because $x$ is unsortable in $(at,R)$ and $ax\nsubseteq R$. Recall that by the hypothesis of the lemma, there exists $y$ that is sortable in $(t,R)$. Since $y$ is unsortable in $(at,R)$, Lemma 24 implies that $ay\subseteq R$. Then $y$ must follow the last occurrence of $x$ in $R$, because $ax\nsubseteq R$. However, with Lemma 22, the fact that $y$ is sortable in $(t,R)$ while $x$ is unsortable in $(t,R)$, and the fact that $x$ is one of the first two colors to appear in $tR$, implies that $y$ cannot follow the last occurrence of $x$ in $R$, a contradiction. ∎
Claim 36.
There are distinct colors $c$ and $b$ that are distinct from $a$, $r$, and $t$, such that there exists $P\in\{arat,rart,arbrt,arcbct\}$ such that $P\subseteq R$. In particular, we have $art\subseteq R$.
Proof.
First, suppose that $rat\subseteq R$. Since, by Claim 35, we have $ar\subseteq R$, one of $arat$, $rart$, or $ratr$ appears in $R$. Since $atdratr=atdraI(r)tr\cong I_{1,0}$, we can assume the pattern $ratr$ does not appear $R$, so we are done in this case.
Now, suppose $rat\nsubseteq R$. Since $r$ is the first color in $R$ that is not $a$ or $t$, $rat\nsubseteq R$ implies that the sandwich that blocks $t$ in $atR$ cannot contain $a$. So there exist distinct colors $c,b$ that are not $a$ or $t$, such that a sandwich of the form $cbc$ blocks $t$ in $atR$. Since, by Claim 35, we have $at\subseteq R$, and $r$ is the first color that is not $a$ or $t$ to appear in $R$, and $rat\nsubseteq R$, we have either $r=c$ and $arbrt\subseteq R$ or $r\notin\{b,c\}$ and $arcbct\subseteq R$. ∎
Claim 37.
For distinct colors $a$, $r$, $t$, $b$, and $c$ and every $P\in\{arat,rart,arbrt,arcbct\}$ the color $a$ is the only sortable color in $atP$. Therefore, no sandwich consisting only of the colors that appear in $atP$ blocks $a$ in $X$. In particular, this implies that $rt\nsubseteq S$ and $atPa\nsubseteq X$.
Proof.
It is not hard to see that $a$ is the only sortable color in $atP\subseteq S^{\prime}R$. Furthermore, it is not hard to see that $a$ is a good sortable color in $atP\subseteq S^{\prime}R$. Recall that, by assumption, $X-d$ is sortable. This means that for any sorting sequence that sorts $X-d$, the color $a$ must be sorted first among the colors in $atP$. Therefore, there is no sandwich consisting only of the colors that appear in $atP$ that blocks $a$ in $X$. ∎
Note that Claims 36 and 37 establish the first statement of the conclusion.
Claim 38.
Let $P\in\{arat,rart,arbrt,arcbct\}$ such that $P\subseteq R$ where $a$, $r$, $t$, $b$, and $c$ are distinct colors. One of the following holds.
  1. (i)
    $S^{\prime}$ contains one of $dat$, $adt$, or, for some color $e\notin P$, $S^{\prime}$ contains $atde$; or
  2. (ii)
    $R^{\prime}$ contains one of $taxat$, $atxat$, $txaxt$, or $taxyxt$ where $a$, $t$, $x$, and $y$ are distinct colors; or
  3. (iii)
    for some color $e\notin P$, $R^{\prime}$ contains $eP$ and $S^{\prime}$ contains one of $eat$, $aet$, or $ate$
Proof.
By assumption, there exists a $d$-sandwich that blocks $a$ in $X=S^{\prime}dR^{\prime}$. That is, for some color $e$, we either have a sandwich of the form $ded$ which blocks $a$ or a sandwich of the form $ede$ which blocks $a$. Therefore, we have the following two cases.
Case 1: A sandwich of the form $ded$ blocks $a$ in $X$. Note that since $d\notin R^{\prime}$, we have $d\in S^{\prime}$. So, if $S^{\prime}$ does not contain $dat$ or $adt$, then $S^{\prime}$ must contain $atde$ and $e\neq t$. So to show that $X$ satisfies (i), we only need to show that $e\notin P$. First note that $e\neq a$, since $ded$ blocks $a$. If $e=r$, then, since we always have $art\subseteq P$, the sequence $S^{\prime}dP\subseteq X$ contains $atdrdart\cong abdtdatb\cong C_{2,5}$, a contradiction. We get a similar contradiction if $P=arbrt$ or $P=arcbct$ and $e=b$, since in either case $P$ contains $abt$, so $S^{\prime}dP$ contains $atdbdabt\cong abdtdatb\cong C_{2,5}$. If $P=arcbct$ and $e=c$, then $P$ contains $act$, so $S^{\prime}dP$ contains $atdcdact\cong abdtdatb\cong C_{2,5}$. This completes the proof of this case.
Case 2: A sandwich of the form $ede$ blocks $a$ in $X$. Since $d\notin R$, we have $e\in S^{\prime}$. Furthermore, if $de\subseteq S^{\prime}$, then a sandwich of the form $ded$ blocks $a$ in $X$, and this was handled in Case 1. Therefore, we can assume that $ea\subseteq R^{\prime}$.
Case 2.1: $e=t$. In this case, $ta=ea\subseteq R^{\prime}$ and, by Claim 37 a sandwich of the form $trt$ does not block $a$ in $X=S^{\prime}dR^{\prime}$. The fact that $at\subseteq S^{\prime}$ then implies that $rta\nsubseteq R^{\prime}$. Therefore, there exists a $t$ in $R^{\prime}$ that is before both an $a$ in $R^{\prime}$ and the first $r$ in $R^{\prime}$. So, if $P=arat$, then we have either $tarat=eP$ or $atrat$ in $R^{\prime}$, so (ii) holds. Furthermore, if $P=rart$, we have $trart=eP$ in $R^{\prime}$, so (ii) holds. If $P=arbrt$, then one option is $tarbrt=eP\subseteq R^{\prime}$, and (ii) holds in this case. By Claim 37, there is no $rbr$ sandwich that blocks $a$, so the only other possibilities when $P=arbrt$ are for $R^{\prime}$ to contain $atrabrt$ or $atrbart$, and these both contain $trart$, so (ii) holds. Finally, assume $P=arcbct$. If $tarcbct=eP\subseteq R^{\prime}$, then $R^{\prime}$ contains $tacbct$, so (ii) holds. Otherwise, because there is no $cbc$ sandwich that blocks $a$, the pattern $R^{\prime}$ must contain one of $atracbct$, $atrcabct$, or $atrcbact$. These contain $tacbct$, $tcact$, and $tcact$, respectively, so (ii) holds.
Case 2.2: $e=r$. In this case, $r\in S^{\prime}$ and $ra=ea\subseteq R^{\prime}$. Note that, by Claim 37, there is no sandwich of the form $rtr$ that blocks $a$ in $X$. Therefore, $r$ must follow $at$ in $S^{\prime}$, so $atr\subseteq S^{\prime}$. Since $ar\subseteq P$ and $t\in P$, either $art$, $atr$, or $tar=tI(r)ar$ is in $P\subseteq R$. Therefore, Lemma 27 implies that a subpattern of $X$ is in $\Gamma$, a contradiction.
Case 2.3: $e\neq r$ and $e\neq t$. Recall that we have $e\in S^{\prime}$ and $ea\subseteq R^{\prime}$. Since $e\in S^{\prime}$ and $at\subseteq S^{\prime}$, one of $eat$, $aet$, or $ate$ is in $S^{\prime}$. We will first prove that $e\notin P$ and $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
Assume first that $e\in R$ and recall that $r$ was defined to be the first color in $R$ that is neither $t$ nor $a$. Therefore, $rea\subseteq R^{\prime}$, which implies that a sandwich of the form $ere$ blocks $a$ in $X\supseteq S^{\prime}dR^{\prime}$. Note that this, with Claim 37, implies that $e\notin P$. Recall that Claim 37 implies that every color in $atP$ is unsortable except $a$. If $e$ is blocked by a sandwich consisting of colors only in $P$, then the pattern formed by removing all colors except those in $P$ and $e$ from $X-d$ is trivially unsortable. This contradicts the assumption that all strict subpatterns of $X$ are sortable. Therefore, we can assume that $e$ is not blocked by a sandwich consisting of colors only in $P$ in $X$.
Now assume that $e\notin R$. In this case, $P\subseteq R$ implies that $e\notin P$. By the definition of $R^{\prime}$, the fact that $e\in R^{\prime}$ implies that the last $e$ follows the last $d$ in $X$. Since $e\notin R$ and the state after $d$ is sorted is $(S,R)$, this means that at some sorting step before the color $d$ was sorted, $de$ appeared on the stack. Therefore, $e$ was sorted before $d$. Since $d$ was sorted before any color in $P$, the color $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
So, in all cases, $e\notin P$ and $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$. In particular, the pattern $ata$ does not block $e$ in $X$. This means that $e$ does not follow $a$ in $R^{\prime}$. Note first that if $eP\subseteq R^{\prime}$, then (iii) holds. So we can assume $eP\nsubseteq R^{\prime}$. This, with the fact that $e$ does does not follow $a$ in $R^{\prime}$ and $e\in R^{\prime}$, implies that $P=rart$ and $reart\subseteq R^{\prime}$. Therefore, $X=S^{\prime}dR^{\prime}$ contains either $eatdreart$, $aetdreart$, or $atedreart$. These contain $eatrear\cong dabtdat\cong I_{1,0}$, $aetreart\cong adbtdatb\cong\mathcal{C}_{2,4}$, and $atereart\cong abdtdatb\cong\mathcal{C}_{2,5}$, respectively, which is a contradiction. ∎
When Claim 38(i) holds, then $X$ must contain one of the patterns in the following table.
$datdarat\cong\mathcal{C}_{1,1}$ $datdrart\cong\mathcal{C}_{1,7}$ $datdarbrt\cong\mathcal{C}_{1,12}$ $datdarcbct\succeq\mathcal{C}_{1,12}$
$adtdarat\cong\mathcal{C}_{1,4}$ $adtdrart\cong\mathcal{C}_{1,8}$ $adtdarbrt\cong\mathcal{C}_{1,11}$ $adtdarcbct\succeq\mathcal{C}_{1,11}$
$atdedarat\cong\mathcal{C}_{1,5}$ $atdedrart\cong\mathcal{C}_{1,9}$ $atdedarbrt\cong\mathcal{C}_{1,13}$ $atdedarcbct\succeq\mathcal{C}_{1,13}$
If Claim 38(iii) holds, then by swapping $d$ and $e$, we can see that $X$ must contain one of these same patterns. Finally, if Claim 38(ii) holds, then $X$ must contain one of the patterns in the following table.
$atdtarat\cong\mathcal{C}_{1,2}$ $atdatrat\cong\mathcal{C}_{1,3}$ $atdtrart\cong\mathcal{C}_{1,6}$ $atdtarbrt\cong\mathcal{C}_{1,10}$

### v2

@@
 Abstract.
-A sock ordering is a finite string of elements, each of which has a color.
-Foot-sorting is a non-deterministic sorting algorithm with the aim of grouping all socks of the same color together, under the constraints of a stack structure; a sock ordering is foot-sortable if there is some way to apply the foot-sorting algorithm such that this aim is realizable.
-In introducing sock orderings and the foot-sorting algorithm, Defant and Kravitz ask whether there is some list of patterns which all foot-sortable sock orderings must avoid, thus characterizing which sock orderings are foot-sortable.
-We enumerate this list, the basis of foot-sortable sock orderings, and show that this basis is infinite.
-We also describe a deterministic polynomial-time algorithm for sorting socks.
+Defant and Kravitz considered the following problem: Suppose that, to the right of a foot, there is a line of colored socks that needs to be sorted.
+However, at any point in time, one can only either place the leftmost sock to the right of the foot onto the foot (stack) or remove the outermost sock on the foot and make it the rightmost sock to the left of the foot (unstack).
+In this paper, we explicitly describe all minimal initial sock orderings that are unsortable.
+TM: Department of Mathematics and Statistics, University of South Florida, Tampa, FL. Email: molla@usf.edu. Research supported in part by NSF grants DMS-1800761 and DMS-2154313.
+CN: Department of Mathematics, University of Tennessee, Knoxville, TN.
+Email: cnelso74@vols.utk.edu.
+The majority of this research effort was performed while this author was an undergraduate student at the University of South Florida.
 1. Introduction
-The idea of sorting permutations, introduced by Donald Knuth in [2], considers sorting a given permutation of $\{1,2,\dots,n\}$ to a desired output under the constraints of a stack structure.
-In [2], Knuth showed that a permutation is sortable if and only if the permutation avoids the pattern 231.
-In other words, the permutation $a_{1}a_{2}\cdots a_{n}$ is sortable using the stack structure if and only if there is no set of indices $1\leq i<j<k\leq n$ such that $a_{k}<a_{i}<a_{j}$.
-Extending Knuth’s findings, Defant and Kravitz introduced the concept of foot sorting for particular patterns in [1].
-This alteration allows for single orderings, called sock orderings, to contain multiple socks of the same color, whereas the original work dealt with orderings containing the elements $\{1,2,\dots n\}$ exactly once.
-A significant difference in sorting a sock ordering when compared to a permutation is that in sock orderings, there is no pre-specified sorted word which the foot-sorting algorithm must produce.
-Because of the freedom in output, sorting a sock ordering that is given to be foot-sortable may be easier than sorting a permutation pattern.
-Defant and Kravitz prove a similar result to Knuth in [1]: that a sock ordering $\rho$ is foot-sortable if and only if there is some word $\nu$ representing $\rho$ such that $\nu$ avoids the pattern 231.
-While for sock orderings with only a few colors, one could search for such a word which avoids 231, the process of generating all possible words and checking whether they avoid 231 is very computationally intensive.
-Allowing for multiple iterations of the stack-sorting procedure, Xia investigated a deterministic foot-sorting algorithm in [3] which hinges on pattern avoidance within the stack, though our work focuses on the sock orderings which are foot-sortable in a single iteration of the sorting procedure.
-Recently, [4] provided a polynomial-time algorithm for deciding the foot-sortability of an arbitrary sock ordering; we later offer a deterministic algorithm which operates in polynomial time and classifies sock orderings as sortable or unsortable, successfully sorting the orderings for which it is possible.
-Defant and Kravitz introduce the idea of a basis of foot-sortable sock orderings in [1].
-The basis of foot-sortable sock orderings is the smallest set of sock orderings such that a sock ordering $X$ is foot-sortable if and only if no member of the set is a subpattern of $X$; we introduce several key ideas in relation to this basis and subpatterns in the following section.
-Defant and Kravitz pose two interesting questions about the basis of foot-sortable sock orderings in [1]:
+Consider the following problem, which was introduced by Defant and Kravitz [1].
+Starting with an arbitrary line of colored socks to the right of a foot, one aims to line the socks up by color to the left of the foot, but, at any point in time, only one of the following two actions can be performed.
+  1. (stack)
+    Take the leftmost sock on the right and place it on the foot (potentially over any socks that are already on the foot).
+  2. (unstack)
+    Remove the outermost sock from the foot and place it so that it becomes the rightmost sock on the left.
+The initial ordering of socks on the right is a sock ordering and a sock ordering is foot-sortable (or simply sortable) if, using only the two specified operations, one can produce a sorted ordering of the socks on the left.
+Otherwise, it is unsortable.
+Throughout, we represent sock orderings as strings in the obvious way; left-to-right in the string corresponds to left-to-right in the sock ordering.
+For example, if $r$, $g$, and $b$ refer to socks of color red, green, and blue, respectively, the sock ordering $rgbrgb$ is sortable (to $rrbbgg$ on the left) while the sock ordering $rgbgrgb$ is unsortable.
+Let $X$ and $Y$ be sock orderings.
+We say that $Y$ is a subordering of $X$ and write $Y\subseteq X$ (or $X\supseteq Y$) if deleting zero or more socks from $X$ yields $Y$ exactly.
+The subordering $Y\subseteq X$ is proper if $X\neq Y$.
+An unsortable sock ordering is minimally unsortable if every proper subordering is sortable.
+We write $a\in X$ if the color $a$ appears in $X$ and we let $X-a$ be the subordering of $X$ obtained by deleting all socks of color $a$ from $X$.
+We use $\emptyset$ to denote the empty ordering and $a$ to denote the ordering consisting of a single sock of color $a$.
+We let $XY$ be the concatenation of the orderings $X$ and $Y$.
+We say $X$ and $Y$ are equivalent and write $X\cong Y$ if $Y$ can be obtained from $X$ via a bijective mapping of the colors of $X$ to the colors of $Y$.
+We call $Y$ a subpattern of $X$ and write $Y\preceq X$ (or $X\succeq Y$) if there exists a sock ordering $Z$ such that $Z\subseteq X$ and $Z\cong Y$.
+Otherwise, we say that $X$ avoids $Y$.
+Furthermore, $X$ avoids a collection of sock orderings if it avoids every ordering in the collection.
+Modulo the relation $\cong$ and with respect to the partial ordering $\preceq$, Defant and Kravitz called the set of minimal elements among the collection of unsortable sock orderings the basis of foot-sortable sock orderings [1].
+One may equivalently view the basis of foot-sortable sock orderings as a collection of pairwise inequivalent minimally unsortable sock orderings that no unsortable sock ordering avoids.
+In this paper, we give an explicit description of the basis of foot-sortable sock orderings (Section 2.2) and a proof that no unsortable sock ordering avoids this basis (Theorem 3).
+1.1. Prior Work
+As observed by Defant and Kravitz [1], this problem fits into the rich line of research dealing with sorting procedures in which the employed data structures are restricted.
+For example, sorting permutations with a stack was considered by Knuth [3].
+Defant and Kravitz also considered several variants of this problem, including employing multiple stacks (in particular two stacks or two feet) and restricting the number of socks of a given color.
+Of particular interest with respect to socks, is the case when there are at most two socks of every color.
+Independent of our work, Yu explicitly described the basis for foot-sortable orderings in which each color appears at most twice [6].
+In doing so, Yu determined that this basis is infinite, and this implies that the basis of foot-sortable sock orderings is also infinite.
+This answered a question of Defant and Kravitz which our work also addresses.
+Also independent of our work, Yu devised a fast deterministic algorithm that decides if a sock ordering of length $N$ is sortable in time $O(N\log N)$.
+Provided that the ordering is sortable, this algorithm can also procedure a valid sorting.
+This problem was also investigated by Xia, who considered a deterministic algorithm that forces the stack to remain sorted [5].
+2. Results
+2.1. Preliminaries
+Throughout, we will refer to the sorting process using primarily stack-centered language.
+We denote a particular state of the foot-sorting process as $(S,R)$.
+Here $S$ represents the contents of the stack, with left-to-right in $S$ corresponding with bottom-to-top in the stack (or inside-to-outside on the foot), while $R$ represents what remains on the right.
+We say that the state $(S,R)$ is foot-sortable (or just sortable) if we can apply the operations stack and unstack so that all of the socks in $SR$ are arranged by color to the left of the foot.
+Note that if the stack $S$ is not sorted, then $(S,R)$ is unsortable.
+As such, we implicitly assume throughout that the stack $S$ is sorted.
+Instead of dealing with the two operations stack and unstack, we work almost exclusively at a slightly higher level.
+At each point, we are mainly concerned with the next sock color that will be moved completely to the left and we say that a color is sortable if it is possible to do this while keeping the stack is sorted.
+More explicitly, given a state $(S,R)$, a color $a\in SR$ is sortable if the following two conditions hold.
   1. (1)
-    Is it possible to enumerate the basis of foot-sortable sock orderings?
+    For any $b\in S$ which is distinct from $a$, $ab\not\subseteq S$. (That is, if $a$ appears on the stack, it must be the top color on the stack.)
   2. (2)
-    Is the basis of foot-sortable sock orderings finite?
-In Section 2, we will define a infinite set, $\Gamma$, of minimally unsortable sock orderings.
-The primary result of our work is the following theorem which proves that $\Gamma$ is the basis for foot-sortable sock orderings.
-Theorem 1.
-If $X$ is an unsortable sock ordering, then a subpattern of $X$ is in $\Gamma$.
-The proof of Theorem 1 leads naturally to a deterministic algorithm which we present in Section 3.
-2. Notation and Preliminaries
-Before enumerating the basis of foot-sortable sock orderings, we lay out some important definitions regarding sock orderings and the patterns they contain as well as provide some examples of pattern containment.
+    If $a$ appears in $R$, then $bcba\not\subseteq SR$ for any colors $b,c\in SR$ such that $a$, $b$, and $c$ are distinct.
+If a color $a\in R$ does not meet the second condition, we say that the color $a$ is blocked by a sandwich.
+We refer to a sandwich of the form $bcb$ as either a $b$-sandwich or a $c$-sandwich.
+Given a state $(S,R)$, a color $a\in SR$ is a good sortable color if $a$ is sortable in $(S,R)$ and what remains after sorting $a$ in $(S,R)$ is foot-sortable.
+If no color is sortable in $(S,R)$, then we say that $(S,R)$ is terminal.
+Given a sock ordering $X$, a color $a\in X$ is sortable (resp.
+unsortable) if it is sortable (resp.
+unsortable) in $(\emptyset,X)$, and a sock ordering $X$ is trivially unsortable if no color in $X$ is sortable.
+In other words, $X$ is trivially unsortable if the state $(\emptyset,X)$ is terminal.
+Note that sorting the color $a$ transitions the state $(S,R)$ to the state $(S^{\prime}R^{\prime},R^{\prime\prime})$ where $S^{\prime}=S-a$ (so $S^{\prime}=S$ if $a$ is not on the stack $S$); and, when $a\in R$, $R^{\prime}$ is the string which precedes the final $a$ in $R$ minus the color $a$ and $R^{\prime\prime}$ is the string which follows the final $a$ in $R$.
+When $a$ does not appear in $R$, we have $R^{\prime}=\emptyset$ and $R^{\prime\prime}=R$.
+A sequence of distinct colors $\theta=a_{1},\dotsc,a_{\ell}$ is a sorting sequence for $(S,R)$ if there are orderings $S_{1},\dotsc,S_{\ell}$ and $R_{1},\dotsc,R_{\ell}$ such that when $S_{0}:=S$ and $R_{0}:=R$, for every $i\in[\ell]$, the color $a_{i}$ is sortable with respect to the state $(S_{i-1},R_{i-1})$ and sorting $a_{i}$ in this state leads to $(S_{i},R_{i})$.
+We define $\theta(S,R):=(S_{\ell},R_{\ell})$ and we say that $\theta$ fully sorts $X$ if $S_{\ell}=R_{\ell}=\emptyset$.
+For a sock ordering $X$, we let $\theta(X):=\theta(\emptyset,X)$.
+2.2. The basis of foot-sortable sock orderings
+We use $\Gamma$ to denote the basis of foot-sortable sock orderings. As mentioned previously, the set $\Gamma$ is infinite and we use the following definition to describe the infinite classes.
 Definition 1.
-A sock ordering $X$ is a finite sequence of socks, each of which has a specified color.
-We will use $|X|$ to denote the length of a sock ordering $X$, the number of socks in $X$.
-Throughout our work, we will use capital letters to denote sock orderings, while we use lowercase letters to denote specific socks or colors.
-At times, we refer to sock orderings as strings in relation to our algorithms.
-Definition 3.
-Let $X=x_{1}x_{2}\cdots x_{n}$ and $Z=z_{1}z_{2}\cdots z_{n}$ be sock orderings (so $|X|=|Z|$), and let $\mathcal{A}_{X}$ and $\mathcal{A}_{Z}$ be the sets of distinct sock colors appearing in $X$ and $Z$ respectively.
-$X$ is equivalent to $Z$ (denoted $X\cong Z$) if there exists a bijective mapping $\varphi:\mathcal{A}_{X}\rightarrow\mathcal{A}_{Z}$ such that $\varphi(x_{i})=z_{i}$ for all $1\leq i\leq n$.
-Definition 4.
-Let $X$ and $Y$ be sock orderings.
-We say that $X$ contains $Y$ and write $Y\subseteq X$ if deleting zero or more socks from $X$ yields exactly $Y$, and we say $a\in X$ if $a\subseteq X$.
-If $Y\subseteq X$, we say that $Y$ is a subordering of $X$.
-In contrast, we say that $Y$ is a subpattern of $X$ and write $Y\preceq X$ if there exists a subordering $Z\subseteq X$ of $X$ such that $Z\cong Y$.
-If $Y$ is not a subpattern of $X$, then we write $Y\npreceq X$ and say that $X$ avoids $Y$.
-For a color $a\in X$, we let $X-a$ be the subordering of $X$ formed by removing every sock of color $a$ from $X$.
-Because of the many equivalent ways to represent one sock ordering, it may not always be immediately clear whether a sock ordering is a subpatterns of another. The following example will illustrate:
-Example 5.
-$babc\preceq abacbab$.
-We have $babc\nsubseteq abacbab$, but $abac\subseteq abacbab$ and $abac\cong babc$. Therefore, $babc\preceq abacbab$.
-To sort, or attempt to sort, sock orderings, we utilize the foot-sorting algorithm proposed by [1], and we outline the rules here.
-Definition 6.
-foot-sorting is a non-deterministic algorithm which operates on a sock ordering $X$. At any point in the foot-sorting process, one may choose to:
-  • Stack a sock: Move the leftmost sock of the remaining input ordering to the top of the stack, or
-  • Unstack a sock: Remove the outermost sock on the foot (the topmost element of the stack) and place it on the right of the output.
-Though we utilize the idea of foot-sorting and socks, throughout, we will refer to the sorting process using primarily stack-centered language.
-At some points, this algorithm has set steps.
-For example, the first step in foot-sorting always must be to place the first sock, if it exists, into the stack.
-Similarly, once nothing remains of the original input, all socks in the stack must be removed beginning with the topmost element of the stack and moving downward.
-We denote a particular stage in the foot-sorting process as $(S,R)$.
-$S$ represents the contents of the stack, with left-to-right in $S$ corresponding with bottom-to-top in the stack (or inside-to-outside on the foot, in the language of [1]), while $R$ represents what remains of the input ordering.
-Before any sorting steps are applied, we may say that an input ordering $X$ appears as $(\emptyset,X)$.
-To illustrate this process and notation, we provide an example of foot-sorting here.
-Example 7.
-Using the foot-sorting algorithm, we may obtain the output $aaccbb$ from the input sock ordering $abcabc$.
-  1. (1)
-    We equivalently write $abcabc$ as $(\emptyset,abcabc)$. Move the leftmost sock of the remaining input ordering, $a$, to the top of the stack. This yields $(a,bcabc)$.
-  2. (2)
-    Remove the topmost sock of the stack, $a$, and place it on the right of the output, which is currently empty. We now have $(\emptyset,bcabc)$ with output $\kappa=a$.
-  3. (3)
-    Move the leftmost sock of the remaining input ordering, $b$, to the top of the stack, leaving $(b,cabc)$ and output $\kappa=a$.
-  4. (4)
-    Stack $c$ to leave $(bc,abc)$ and output $\kappa=a$.
-  5. (5)
-    Stack $a$ to leave $(bca,bc)$ and output $\kappa=a$.
-  6. (6)
-    Unstack $a$ to leave $(bc,bc)$ and output $\kappa=aa$.
-  7. (7)
-    Unstack $c$ to leave $(b,bc)$ and output $\kappa=aac$.
-  8. (8)
-    Stack $b,c$, leaving $(bbc,\emptyset)$ and output $\kappa=aac$.
-  9. (9)
-    Unstack, in order, all remaining socks from the stack to leave (,)($\emptyset$,$\emptyset$)( ∅ , ∅ ) and output $\kappa=aaccbb$.
-Definition 8.
-A stage of the sorting algorithm $(S,R)$ is foot-sortable (or, when clear, sortable) if there is some way to apply the foot-sorting algorithm on $(S,R)$ such that, in the output of the foot-sorting process, all socks of the same color occur consecutively.
-That is, for some $(S,R)$ containing only the colors $\{a_{1},a_{2},\dots a_{n}\}$, $(S,R)$ may be sorted to $\kappa=a_{\pi(1)}\cdots a_{\pi(1)}a_{\pi(2)}\cdots a_{\pi(2)}\cdots a_{\pi(n)}\cdots a_{\pi(n)}$ where $\pi$ is a permutation of $\{1,\dotsc,n\}$.
-If there is no such way to apply the foot-sorting algorithm, $X$ is unsortable.
-We say that a sock ordering $X$ is foot-sortable (resp.
-unsortable) if $(\emptyset,X)$ is foot-sortable (resp.
-unsortable).
-We see that $abcabc$ in Example 7 is a sortable sock ordering.
-The goal of foot-sorting is to begin with an input sock ordering $X$ and apply the steps of the algorithm so that nothing remains in the stack or on the right side.
-While it can be difficult to decide whether a particular sock ordering is sortable or unsortable, there is an important indicator which can reveal if applying the foot-sorting algorithm has already failed.
-Note that a sortable sock ordering $X$ may be sorted to a word $\kappa$ such that $aba\not\preceq\kappa$.
-Due to the nature of the stack structure, once a sock, $a$, appears in the stack, $a$ must appear in the output $\kappa$ before any socks appearing below $a$ in the stack and after any socks appearing above $a$ in the stack.
-For $(S,R)$ appearing at some point while applying the foot-sorting algorithm, let $S=s_{1}s_{2}\cdots s_{n}$ and set $S^{-1}=s_{n}\cdots s_{2}s_{1}$, the reverse of $S$.
-It must be that $S^{-1}\subseteq\kappa$.
-Since in a sorted word $\kappa$, we have $aba\not\preceq\kappa$, and $(aba)^{-1}=aba$, then if $aba\preceq S$, it is not possible to successfully sort the sock ordering.
-Because of this, if we ever have $aba\preceq S$ while attempting to sort a particular sock ordering, we will immediately stop the sorting attempt as it is not possible to finish sorting after this point.
-We refer to this $aba$ pattern of distinct colors $a,b$ as a sandwich.
-Sandwiches, the colors they contain, and where they appear in sock orderings arise frequently in our work.
-We may say that the sandwich $aba$ blocks, precedes, or appears before the color $c$, for distinct colors $a,b,c$, at some stage $(S,R)$ if $abac\subseteq SR$.
-Additionally, we are sometimes concerned about the first sandwich $aba$ which blocks a color $c$.
-If two distinct sandwiches, $aba$ and $xyx$, precede $c$, then if $a$ and $x$ are distinct colors, $aba$ appears earlier than $xyx$ if the second denoted $a$ appears before the second denoted $x$; if $x$ is $a$, then $aba$ appears earlier than $aya$ if the denoted $b$ appears before the denoted $y$.
-For example, in $abcabd$, all of the sandwiches $aba,aca,bcb,bab$ block $d$, but $aba$ is the first sandwich which blocks $d$.
-Finally, we may say $x$-sandwich to describe a sandwich which contains a sock of color $x$.
-Interestingly, though it is necessary to avoid $aba$ appearing in the stack while applying the foot-sorting algorithm, a deterministic algorithm which prevents the appearance of sandwiches in the stack does not succeed in sorting all sortable sock orderings; Xia investigated such an algorithm, $\phi_{aba}$, in [3].
-We later introduce a deterministic algorithm $\phi_{G}$ which, as a consequence of its operation, prevents $aba$ from appearing in the stack, but has further restrictions that differentiate it from $\phi_{aba}$.
-It was also shown that $\phi_{aba}$ in [3] may sometimes sort sock orderings which have subpatterns that are unsortable under $\phi_{aba}$.
-In foot-sorting, however, this is not the case; if there exists no set of steps which may sort some subpattern $Y\preceq X$, there can be no set of steps which may sort $X$.
-For every foot-sortable sock ordering $S$, $Y\preceq S$ implies that $Y$ is foot-sortable.
-Definition 9.
-At a particular sorting stage $(S,R)$, some color $a\in SR$ is sortable if both of the following conditions hold:
-  1. (1)
-    For any $b\in S$ which is distinct from $a$, $ab\not\subseteq S$. (That is, if $a$ appears on the stack, it must be the top element on the stack.)
-  2. (2)
-    If $a$ appears in $R$, then $bcba\not\subseteq SR$ for any colors $b,c\in SR$ such that $a$, $b$, and $c$ are distinct. (In other words, the color $a$ is not blocked by a sandwich.)
-Otherwise, $a$ is unsortable in $(S,R)$.
-We say that a color $a\in X$ is sortable in $X$ if $a$ is sortable in $(\emptyset,X)$.
-If no color is sortable in $(S,R)$, then we say that $(S,R)$ is terminal.
-The following observation follows from the definitions.
-Observation 10.
-Let $S$ and $R$ be sock orderings. Suppose $x\in SR$ and that $xy\not\subseteq S$ for any $y\in S$ distinct from $x$. Then the color $x$ is sortable in $(S,R)$ if and only if $x$ is sortable in $SR$.
-Definition 11.
-A color $a$ is a good sortable color at a stage $(S,R)$ if $a$ is sortable in $(S,R)$ and what remains after sorting $a$ in $(S,R)$ is foot-sortable.
-At times, we may reach a stage in which multiple colors are sortable.
-For some colors $\{a_{1},\dots,a_{n}\}$ in a string $X$, we say $a_{1}$ is the first terminating color from the set if the final sock of color $a_{1}$ appears before the final sock of color $a_{i}$ for all $a_{i}$ in the set distinct from $a_{1}$.
-Definition 12.
-A trivially unsortable sock ordering is a sock ordering in which no color is sortable.
-That is, a sock ordering $X$ is trivially unsortable if for every color $a\in X$, there exist distinct colors $b,c\in X$, different from $a$, such that $bcba\subseteq X$.
-We often discuss particular sorting choices in an attempt to sort a sock ordering $X$.
-A move in some sorting procedure consists of sorting exactly one color.
-That is, when $(S,R)$ appears, a color $x$, identified as a sortable in $(S,R)$, is chosen to be sorted.
-Any $x$ are removed from the top of $S$, and if $x\in R$, then for every sock in $R$, we apply the following rule:
-  • If the sock is the final sock of color $x$, place the sock into the stack, then remove it from the stack and place the sock on the sorted left side.
-Immediately terminate the procedure at this point.
-  • If the sock has color $x$ but is not the final sock of color $x$, stack the sock and immediately unstack it, bringing it to the sorted left side. Look to the next sock on the right side.
-  • If the sock has a color which is not $x$, stack the sock, and do not unstack it. Look to the next sock on the right side.
-In contrast, a step in the sorting procedure refers to the movement of exactly one sock: either stacking the first sock appearing on the right, or removing the top element of the stack and bringing it to the sorted left side.
-In general, it takes multiple steps to complete one sorting move.
-When we say that we sort the color $a$ we mean that we apply a move as describe above of the sortable color $a$.
-Definition 13.
-Let $X$ be a sock ordering, and let $\mathcal{A}_{X}$ be the set of distinct colors appearing in $X$.
-A sorting sequence $\theta$ for $X$ is a nonempty sequence $(a_{1},a_{2},\dots,a_{n})$ of elements from $\mathcal{A}_{X}$, each appearing at most once, such that $a_{1}$ is sortable in $(\emptyset,X)$, and for $i>1$, $a_{i}$ is sortable after $a_{i-1}$ has been sorted.
-We write $\theta(X)=(S,R)$ if, beginning with $(\emptyset,X)$, $(S,R)$ appears after sorting each color from $\theta$ in sequence.
-Observation 14.
-At some stage $(S,R)$, if a color $x$ only appears as the top color in $S$, then $x$ is sortable with respect to $(S,R)$.
-Furthermore, if we sort $x$ to form $(S^{\prime},R)$, then $(S,R)$ is sortable if and only if $(S^{\prime},R)$ is sortable.
-Definition 15.
-We say that a sorting sequence $\theta$ fully sorts $X$ if $\theta(X)=(\emptyset,\emptyset)$.
-In light of Observation 14, we also say that $\theta$ fully sorts $X$ if we can fully sort $X$ using the prescribed sorting order $\theta$ with the additional stipulation that we immediately sort a color if it only appears as the top color on the stack.
-Note that, with this definition, a sorting sequence $\theta$ can fully sort a sequence $X$ even if $\theta$ does not contain all of the colors of $X$.
-While a sorting sequence $\theta$ must obey the rules of foot-sorting (that is, $\theta$ may never make a move that causes a sandwich to appear in the stack), there is no requirement for $\theta$ to fully sort $X$.
-Example 16.
-Taking $\theta_{1}=(a,c,b)$, we have that $\theta_{1}(abcabc)=(\emptyset,\emptyset)$, and thus $\theta_{1}$ fully sorts $abcabc$.
-Choosing instead $\theta_{2}=(a)$, we have $\theta_{2}(abcabc)=(bc,bc)$.
-2.1. The basis of foot-sortable sock orderings
-Definition 17.
-A minimally unsortable, or minimal, sock ordering $X$ is an unsortable sock ordering in which for all $Y\preceq X$, $Y$ is unsortable $\implies$ $|Y|=|X|$.
-Observation 18.
-Let $\Gamma$ be the set of all minimally unsortable sock orderings, and let $M$ be the basis of foot-sortable sock orderings, the smallest set of sock orderings which foot-sortable sock orderings must avoid.
-Then $M=\Gamma$.
-As mentioned in the introduction, the basis of foot-sortable sock orderings is infinite. The following definition is used to describe our infinite classes in the basis $\Gamma$.
-Definition 19.
-Let $T:=(t_{0},\dotsc,t_{n})$ be a sequence of $n+1$ distinct colors. Define $I(t_{0}):=I(T)$ to be the empty string when $n=0$, and when $n\geq 1$, let
+Let $T:=(t_{0},\dotsc,t_{n})$ be a sequence of $n+1$ colors. Define $I(t_{0}):=I(T)$ to be the empty ordering when $n=0$, and when $n\geq 1$, let
 $I(T):=I(t_{0},\dotsc,t_{n}):=I(t_{0},\dotsc,t_{n-1})t_{n}t_{n-1}=t_{1}t_{0}t_{2}t_{1}\dotsc t_{n}t_{n-1}.$
-Observation 20.
+Observation 2.
 Let $n\geq 0$ and $a,b,t_{0},\dotsc,t_{n}$ be distinct colors.
@@
 $I(t_{i},\dotsc,t_{n})at_{n}=at_{n}$
-and it is not hard to see that both statements are true.
+and it is not hard to see $(abt_{n},at_{n})$ is terminal and removing any sock from $(abt_{n},at_{n})$ makes the state sortable. Therefore, both of the statements are true when $i=n$.
 For the induction step, first note that sorting $t_{i}$ in
@@
 leads to $(abt_{i+1},I(t_{i+1},\dotsc,t_{n})at_{n})$.
-Note that, when $n-i=1$ the only sortable color is $t_{i}$, so the first statement follows by induction.
+Note that when $n-i=1$ the only sortable color is $t_{i}$, so the first statement follows by induction.
 When $n-i\geq 2$, the only sortable colors are $t_{i}$ and $t_{i+1}$ and sorting $t_{i+1}$ in
 $(abt_{i},I(t_{i},\dotsc,t_{n})at_{n})=(abt_{i},t_{i+1}t_{i}t_{i+2}t_{i+1}I(t_{i+2},\dotsc,t_{n})at_{n}),$
-leads to $(abt_{i}t_{i}t_{i+2},I(t_{i+2},\dotsc,t_{n})at_{n})$ and the conclusion again follows by the induction hypothesis.
+leads to $(abt_{i}t_{i}t_{i+2},I(t_{i+2},\dotsc,t_{n})at_{n})$. Therefore, the first statement again follows by the induction hypothesis.
 We will now prove the second statement when $n-i\geq 1$. Recall that we are assuming $S^{\prime}\subseteq abt_{i}$ and
 $R^{\prime}\subseteq I(t_{i},\dotsc,t_{n})at_{n}=t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$
-where at this one of the two is a strict subordering.
+where at least one of the two, $S^{\prime}$ or $R^{\prime}$, is a proper subordering.
 By possibly adding socks, we also assume that either $S^{\prime}=abt_{i}$ or $R^{\prime}=t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$.
@@
 Otherwise, we only sort $t_{i}$.
-In either case, we reach $(S^{\prime\prime},R^{\prime\prime})$ where $S^{\prime\prime}$ is a strict subordering of $abt_{t+1}$ and $R^{\prime\prime}=I(t_{i+1},\dotsc,t_{n})at_{n}$.
+In either case, we reach $(S^{\prime\prime},R^{\prime\prime})$ where $S^{\prime\prime}$ is a proper subordering of $abt_{t+1}$ and $R^{\prime\prime}=I(t_{i+1},\dotsc,t_{n})at_{n}$.
 The conclusion then follows from the induction hypothesis.
@@
 So, $R^{\prime}\neq t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$.
-If $R^{\prime}$ does not begin with $t_{i+1}t_{i}$, then we can sort $t_{i}$ and then $b$ to reach $(S^{\prime\prime},R^{\prime\prime})$ where $S^{\prime\prime}\subseteq at_{i+1}$ and $R^{\prime\prime}\subseteq I(t_{i+1},\dotsc,t_{n})at_{n}$.
+If $R^{\prime}$ does not begin with $t_{i+1}t_{i}$, then we can sort $t_{i}$ and then $b$.
+After possibly pushing $t_{i+1}$ onto the stack, we arrive at the state $(S^{\prime\prime},R^{\prime\prime})$ where $S^{\prime\prime}\subseteq at_{i+1}$ and $R^{\prime\prime}\subseteq I(t_{i+1},\dotsc,t_{n})at_{n}$.
 In this case, since $b\notin S^{\prime\prime}$, the conclusion follows from the induction hypothesis.
-Otherwise, we have $S^{\prime}=abt_{i}$ and $R^{\prime}=t_{i+1}t_{i}R^{\prime\prime}$ where $R^{\prime\prime}$ is a strict subordering of $I(t_{i+1},\dotsc,t_{n})at_{n}$.
+Otherwise, we have $S^{\prime}=abt_{i}$ and $R^{\prime}=t_{i+1}t_{i}R^{\prime\prime}$ where $R^{\prime\prime}$ is a proper subordering of $I(t_{i+1},\dotsc,t_{n})at_{n}$.
 Then, after sorting the color $t_{i}$, we have $(abt_{i+1},R^{\prime\prime})$ and the conclusion follows by the induction hypothesis.
 ∎
-We conclude this section by enumerating $\Gamma$, the basis of foot-sortable sock orderings.
-The following list consists of every sock ordering which is minimally unsortable.
-We describe five classes of minimally unsortable sock orderings which have infinitely many members, and we also outline other classes of sock orderings which are finite in number.
-We have divided the basis into categories that correspond with particular subpatterns we will later generate using a greedy sorting algorithm.
-Sock orderings in classes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ have color labels which denote the behavior of particular socks during the greedy sorting algorithm’s procedure.
-We have the following orderings in the class $\mathcal{T}$, separated by the number of distinct colors in the sock ordering:
+The following tables describe the basis $\Gamma$. The first table consists of trivially unsortable orderings. It is not difficult to verify that the orderings are minimally unsortable.
 $\mathcal{T}_{1}\cong abacaba$ $\mathcal{T}_{2}\cong abacbab$ $\mathcal{T}_{3}\cong abcabca$ $\mathcal{T}_{4}\cong abcacba$
@@
 $\mathcal{T}_{8}\cong abcadcab$ $\mathcal{T}_{9}\cong abacdcba$ $\mathcal{T}_{10}\cong abacdcab$
-Note that all of the sock orderings in the following class, $\mathcal{C}_{1}$, contain either $atdarat$, $atdrart$, or $atdarbrt$.
-Furthermore, sorting the color $d$ first for each of these three subpatterns yields $(at,arat)$, $(at,rart)$, and $(at,arbrt)$, respectively, and none of these three states are sortable.
-To then see that the orderings in $\mathcal{C}_{1}$ are unsortable, it is only necessary to verify that in each of these orderings, we must sort $d$ before the colors $a$, $b$, $r$, and $t$.
+Note that all of the sock orderings in the next table have either $atdarat$, $atdrart$, or $atdarbrt$ as a subordering.
+Furthermore, sorting the color $d$ (preceded or followed by sorting the color $e$ when it appears) for each of these three possible suborderings yields $(at,arat)$, $(at,rart)$, and $(at,arbrt)$, respectively, and each of these three states are terminal.
+Further note that in the orderings below, $d$ and $e$ are the only sortable colors, which implies that the orderings are unsortable.
+It is then not difficult to verify that the orderings are minimally unsortable.
 $\mathcal{C}_{1,1}\cong datdarat$ $\mathcal{C}_{1,6}\cong atdtrart$ $\mathcal{C}_{1,10}\cong atdtarbrt$
@@
 $\mathcal{C}_{1,5}\cong atdedarat$
-The other finite class is $\mathcal{C}_{2}$, consisting of those sock orderings which contain either the subordering $abtdabt$ or $abtdatb$.
-Since $(abt,at)$ is unsortable and $d$ must be sorted before the colors $a$, $b$ and $t$ in the following orderings, it is not difficult to verify that the the following orderings are minimally unsortable.
+The next table consists of the sock orderings that have either $abtdabt$ or $abtdatb$ as a subordering.
+In a similar fashion to the other classes, sorting $d$ (and possibly $e$) first in these orderings yields some $(S,R)$ where $abt\subseteq S$ and $at\subseteq R)$.
+Since $(abt,at)$ is unsortable and $d$ and $e$ are the only sortable colors in the following orderings, we have that orderings in this class are unsortable, and it is not difficult to verify that these are minimally unsortable.
 $\mathcal{C}_{2,1}\cong abtdatab$ $\mathcal{C}_{2,4}\cong adbtdatb$ $\mathcal{C}_{2,6}\cong abtdedabt$
@@
 $I_{1,n}:=dabtdE$ $I_{2,n}:=adbtdbE$ $I_{3,n}:=abdtdbE$ $I_{4,n}:=abtdtbE$ $I_{5,n}:=abtdedbE$
-With Observation 20 it is not hard to verify that these orderings are also minimally unsortable.
-3. Algorithm
-We now present a deterministic recursive algorithm (Algorithm 3) for sock sorting.
-On input $X$, the algorithm returns a triple $(\theta,(S,R),X^{\prime})$ where $X^{\prime}\subseteq X$ and $(S,R)=\theta(X^{\prime})$.
-If $X$ is sortable, then $X^{\prime}=X$ and $S=R=\emptyset$, so $\theta$ is a sorting sequence that fully sorts $X$.
-If $X$ is unsortable, then $X^{\prime}$ is an unsortable subordering of $X$ and $R\neq\emptyset$.
-Algorithm 3 closely follows the proof Theorem 1, so the proof of Theorem 1 also serves as a proof that Algorithm 3 is correct.
-Algorithm 2 (what we call the greedy algorithm) is mentioned extensively in our proof.
-We often write $(S,R)=\phi_{G}(X)$ to mean that $(S,R)$ is the state when the greedy algorithm terminates.
-The greedy algorithm simply sorts the sortable color that terminates first in $SR$ until no color is sortable.
-In particular, if a color only appears on the top of the stack, then that color is sorted next.
-At a high-level, the main idea of our sorting algorithm is as follows.
-Let $(S,R)=\phi_{G}(X)$ be the state after the greedy algorithm terminates.
-If we assume $X$ is sortable and the greedy algorithm fails, then we can always find a color $d$ such that any sorting sequence that sorts $X-d$ also sorts $X$.
-Therefore, we can recursively call our sorting routine on $X-d$ and use the resulting sequence to sort $X$.
-In most cases, this color $d$ is the last color sorted by the greedy algorithm.
-However, in some cases, we must backtrack in order to find such a color $d$.
-For brevity, we do explicitly describe the subroutine Move which is referenced below.
-For a given state $(S,R)$ with sortable color $a$, the call $\textrm{Move}((S,R),a)$ returns the state $(S^{\prime},R^{\prime})$ after sorting the color $a$.
-Algorithm 1 Sort using a prescribed sorting order, except we always immediately sort colors that only appear on the top of the stack. The routine returns the actual sequence of colors sorted.
-1:function SortColors($X,\theta$)
-2:     $(S,R)\leftarrow(\emptyset,X)$
-3:     $\theta^{\prime}\leftarrow\emptyset$triangleright $\theta^{\prime}$ records the actual sorting sequence.
-4:     while $R\neq\emptyset$ and $S\neq\emptyset$ do
-5:         if $S\neq\emptyset$ and the top color $t$ on $S$ is not in $R$ then
-6:              Append $t$ to $\theta^{\prime}$
-7:              $(S,R)\leftarrow\textrm{Move}((S,R),t)$
-8:         else if $\theta=\emptyset$ then return $\theta^{\prime},(S,R)$
-9:         else
-10:              $d\leftarrow$ first color in $\theta$
-11:              Remove $d$ from $\theta$
-12:              if $d$ is in $SR$ then
-13:                  if $d$ is not sortable in $(S,R)$ then return $\theta^{\prime},(S,R)$
-14:                  end if
-15:                  Append $d$ to $\theta^{\prime}$
-16:                  $(S,R)\leftarrow\textrm{Move}((S,R),d)$
-17:              end if
-18:         end if
-19:     end while
-20:     return $\theta^{\prime},(S,R)$
-21:end function
-Algorithm 2 When $C=\emptyset$ this is our greedy sorting algorithm.
-Otherwise, we assume $C\subseteq X$ and follow the greedy sorting algorithm, except we stop when a color in $C$ would be sorted next.
-We use the second parameter $C$ to backtrack.
-1:function Greedy($X$, C)
-2:     $(S,R)\leftarrow(\emptyset,X)$
-3:     while $\exists d$ a sortable color in $(S,R)$ that terminates first in $SR$ do
-4:         if $d\in C$ then return $\theta,(S,R)$
-5:         end if
-6:         $(S,R)\leftarrow$ $\textrm{Move}((S,R),d)$
-7:         Append $d$ to $\theta$
-8:     end while
-9:     return $\theta,(S,R)$
-10:end function
-Algorithm 3 The sorting algorithm
-1:function FullSort($X$)
-2:     $\theta,(S,R)\leftarrow\textrm{Greedy}(X,\emptyset)$
-3:     if $R=\emptyset$ or $S=\emptyset$ or $\exists x\in S$ s.t. $(x,R)$ is terminal then return $\theta,(S,R),X$
-4:     end if
-5:     Let $b$ and $t$ be the top two colors on $S$ and let $d$ be the last color in $\theta$
-6:     if $(bt,S)$ is terminal then return $\textrm{SortHelper}(X,d,b)$
-7:     end if
-8:     $a\leftarrow$ the lowest color in $S$ s.t. $(abt,R)$ is terminal
-9:     if $\exists z\in R$ s.t. either $z=b$ or $z$ is between $a$ and $b$ on $S$ then
-10:         if $R$ contains $azt$ or $atz$ then return $\textrm{SortHelper}(X,d,a)$
-11:         else return $\textrm{SortHelper}(X,d,z)$
-12:         end if
-13:     end if
-14:     Let $A$ and $C$ be the strings that precede and follow the last $b$ in $X$ (so $X=AbC$ and $b\notin C$)
-15:     $\theta,(S,R)\leftarrow\textrm{Greedy}(X,C)$triangleright Backtrack to before the final $b$ is pushed onto the stack
-16:     $c\leftarrow$ first terminating sortable color in $(S,R)$triangleright $c$ would be sorted next greedily
-17:     $d\ \leftarrow$ last color in $\theta$triangleright $d$ is the last color sorted
-18:     if $ca\subseteq A$ then
-19:         return $\textrm{Greedy}(X,\emptyset),X$triangleright $X$ is unsortable - return greedy output
-20:     end if
-21:     Let $y$ be the color immediately beneath the top color on the stack $S$
-22:     return SortHelper(X, d, y)
-23:end function
-Algorithm 4 This is the helper routine for the sorting algorithm.
-When it is called, $X$, $a$, $x$, and $d$ must satisfy the hypothesis of Lemma 25.
-Furthermore, we also require that if Lemma 25(I) holds, then $X$ satisfies the hypothesis of either Lemma 27 or Lemma 28, so $X$ is unsortable.
-Note that when both Lemma 25(I) and (II) do not hold, if a sequence that fully sorts $X-d$ does not fully sort $X$, Lemma 25(IV) must hold and Lemma 25(IV) implies that $X$ is unsortable.
-1:function SortHelper($X,d,a$)
-2:     if $a$ is blocked by a $d$-sandwich in $X$ then return $\textrm{Greedy}(X,\emptyset),X$triangleright $X$ is unsortable
-3:     end if
-4:     $\theta,(S,R),X^{\prime}\leftarrow\textrm{FullSort}(X-d)$
-5:     if  $R\neq\emptyset$  then return $\theta,(S,R),X^{\prime}$triangleright $X^{\prime}\subseteq X-d$ is unsortable
-6:     else
-7:         return $\textrm{SortColors}(X,\theta),X$triangleright $\theta$ fully sorts $X-d$, so $\theta$ fully sorts $X\iff X$ is sortable
-8:     end if
-9:end function
-4. Prelimanary Observations and Lemmas
+With Observation 2 it is not hard to verify that these orderings are also minimally unsortable.
+As we have now specified a collection $\Gamma$ of minimally unsortable sock orderings, to prove that $\Gamma$ is the basis of foot-sortable sock orderings, we only need the following theorem.
+Theorem 3.
+If $X$ is an unsortable sock ordering, then a subpattern of $X$ is in $\Gamma$.
+2.3. Algorithm
+Our proof of Theorem 3 is inherently algorithmic, and we feel it is instructive to explicitly describe a deterministic algorithm that, given an arbitrary sock order, either sorts the ordering using only the operations stack and unstack, or indicates that the given sock ordering is unsortable.
+Our proof can be used to extend this algorithm so that, when the given sock ordering is unsortable, it also identifies a subpattern which is an element of the basis, thereby justifying the indication that the sock ordering is unsortable.
+For simplicity, we do not explicitly describe this extension.
+We made no attempt to optimize the running time of this algorithm.
+(Note that, as referenced in the introduction, independent of our work, Yu has already described a fast algorithm [6].)
+The algorithm first attempts to sort using the following simple greedy algorithm.
+This greedy algorithm is the only part of this section that is referenced in the other sections of this paper.
+Greedy algorithm - Input: A sock ordering $X$.
+  1. (G.1)
+    While the current state is not terminal, do the following.
+    1. (G.1.1)
+      Sort the sortable color which terminates first.
+That is, if there is a color that only appears on the top of the stack, then sort that color, and otherwise sort the sortable color on the right which terminates first.
+The main algorithm also makes use of the following helper routine.
+The input to the helper routine is a sock ordering $X$ and colors $d$ and $a$ where we assume the input satisfies the hypothesis of Lemma 8 in the following sense: There is a way to sort $X$ to a state $(S,R)$ so that
+  • $d$ was the color sorted immediately before entering the state $(S,R)$;
+  • for some substack $S^{\prime}\subseteq S$ and subordering $R^{\prime}\subseteq R$, the color $a$ is beneath the top color on $S^{\prime}$ and the color $a$ is the only good sortable color in $S^{\prime}R^{\prime}$.
+(In particular, this implies that $(S^{\prime},R^{\prime})$, and hence $(S,R)$, is unsortable.)
+When this routine is called, the colors $a$ and $d$ will also have the property that if $a$ is blocked by a $d$-sandwich in $X$ (i.e.
+Lemma 8(I) holds), then (by either Lemma 10 or Lemma 11) a subpattern of $X$ is in $\Gamma$, so $X$ is unsortable.
+ Helper routine - Input: A sock ordering $X$ and colors $d$ and $a$.
+  1. (H.1)
+    If $a$ is blocked by a $d$-sandwich, then $X$ is unsortable.
+  2. (H.2)
+    Otherwise, we recursively run the main algorithm on the smaller ordering $X-d$.
+  3. (H.3)
+    If it fails, then by our recursive assumption $X-d$ is unsortable, so $X$ is unsortable.
+  4. (H.4)
+    Otherwise, $X-d$ is sortable and we have the following two possibilities.
+    1. (H.4.1)
+      When we attempt to sort $X$ using the same sequence of colors that fully sorted $X-d$, there is a stage in which the color $d$ only appears as the top color of the stack, which means that we can fully sort $X$ by sorting $d$ at this stage and then continuing with the sequence that was used to sort $X-d$.
+    2. (H.4.2)
+      $X$ is unsortable because $I_{1,n}\in\Gamma$ is a subpattern of $X$.
+We now present the main algorithm. It closely follows the proof Theorem 3 in Section 2.5, so the proof of Theorem 3 also serves as a proof that this algorithm is correct.
+ Main Algorithm - Input: A sock ordering $X$.
+  1. (M.1)
+    Let $(S,R)$ be the output of the greedy algorithm performed on $X$.
+  2. (M.2)
+    If the greedy algorithm succeeds (i.e. $S=R=\emptyset$), then we are done.
+  3. (M.3)
+    If $S=\emptyset$ or there is a color $x\in S$ on the stack such that $(x,R)$ is terminal, then $X$ has a trivially unsortable subpattern, so $X$ is unsortable.
+  4. (M.4)
+    Otherwise, there are at least two colors on the stack, so let $t$ be the top color and let $b$ be the color immediately beneath it on the stack.
+We can also let $d$ be the final color sorted by the greedy algorithm.
+  5. (M.5)
+    If $(bt,R)$ is terminal, then we can finish by running the helper routine with input $X$, $d$, and $b$ (c.f. Lemma 11 and Case 2 in Theorem 3).
+  6. (M.6)
+    Otherwise, let $a$ be the lowest color on the stack such that $(abt,R)$ is terminal.
+(Such a color exists by Lemma 12.) We must have $at\subseteq R$.
+(This follows from Claim 14 in the proof of Theorem 3.)
+  7. (M.7)
+    If there exists a color $z$ that is both in $R$ and above $a$ and below or equal to $b$ on the stack $S$, then do the following (c.f. Lemma 10 and Case 3.1 in Theorem 3).
+    1. (M.7.1)
+      If $azt\subseteq R$ or $atz\subseteq R$, then we can finish by running the helper routine with input $X$, $d$, and $a$.
+    2. (M.7.2)
+      Otherwise, $zat\subseteq R$, and we can finish by running the helper routine with input $X$, $d$, and $z$.
+  8. (M.8)
+    Otherwise, we have $b\notin R$, so if we backtrack to the state $(S^{\prime},R^{\prime})$ immediately before the final $b$ was pushed onto the stack the following holds:
+    • There exists a string $C$ with $b\notin C$ and strings $A$ and $A^{\prime}$ such that $X=AbC$ and $R^{\prime}=A^{\prime}bC$ and $S^{\prime}A^{\prime}\subseteq A$.
+    • The next color sorted by the greedy algorithm (i.e. the color that pushes the final $b$ onto the stack) is a color $c\in C$.
+    • For some nonnegative integer $n$, there are $n+1$ distinct colors $t_{0},\dotsc,t_{n}$ that are distinct from $b$, $c$ and $a$, such that $t_{0}cI(t_{0},\dotsc,t_{n})at_{n}\subseteq C$.
+(This follows from Lemma 4.)
+  9. (M.9)
+    Therefore, if $ca\subseteq A$, then $I_{1,n}\in\Gamma$ is a subpattern of $X$, so $X$ is unsortable.
+  10. (M.10)
+    Otherwise, there are at least three distinct colors on the stack $S^{\prime}$. (This follows from Claim 15 in the proof of Theorem 3.)
+  11. (M.11)
+    In particular, this implies that there is a color $d^{\prime}$ that was sorted immediately before the greedy algorithm reached the state $(S^{\prime},R^{\prime})$.
+  12. (M.12)
+    Furthermore, if we let $y$ be the color immediately beneath the top color on the stack $S^{\prime}$, we can finish by running the helper routine with input $X$, $d^{\prime}$, and $y$ (c.f.
+Lemma 10 and Claim 15 in the proof of Theorem 3).
+2.4. Observations and Lemmas
 In the enumeration of our basis $\Gamma$, the five infinite classes have a common structure. In some sense, the following lemma explains the origin of this commonality.
 Lemma 4.
-Let $X$ be a sock ordering that, for some $n\geq 0$, contains distinct colors $t_{0},\dotsc,t_{n}$, $b$, and $c$ where the final $b$ precedes the final $c$ in $X$.
+Let $X$ be a sock ordering that, for some $n\geq 0$, has distinct colors $b$, and $c$ where the final $b$ precedes the final $c$ in $X$.
 Define $A$ and $C$ to be the strings such that $X=AbC$ where $b\notin C$ and $c\in C$.
-Let $\theta$ be a sorting sequence for $X$ that sorts a color in $C$ and let $c\in C$ be the color in $C$ that is sorted first by $\theta$.
-Taking $(S,R):=\theta(X)$, suppose $b\in S$ and that once $c$ is sorted, $b$ never appears as the top element on the stack.
-Let $t_{0},\dots,t_{n}$ (in this order) be the $n+1$ distinct colors which appear immediately above the color $b$ in the stack after $c$ is sorted.
+Let $\theta$ be a nonempty sorting sequence for $X$ that sorts a color in $C$ and suppose $c$ is the color in $C$ that is sorted first by $\theta$.
+Taking $(S,R):=\theta(X)$, suppose $b\in S$ and that once $c$ is sorted, $b$ never appears as the top color on the stack.
+Suppose $t_{0},\dots,t_{n}$ (in this order) are the $n+1$ colors which appear immediately above the color $b$ in the stack after $c$ is sorted.
 Then
 $I(c,t_{0},\dotsc,t_{n})R\subseteq C.$
+Furthermore, the colors $t_{0},\dotsc,t_{n}$ are distinct and disjoint from $b$ and $c$.
 Proof.
-We will prove the claim by induction on $n$.
-For the base case, assume $n=0$, so $t_{0}=t_{n}$ is the unique color that appears directly above $b$ after $c$ is sorted.
-Since $c$ follows the last $b$ and $b$ can never appear only as the top color on the stack after $c$ is sorted, the color $t_{0}=t_{n}$ must appear above $b$ after $c$ is sorted.
+We will prove the first statement by induction on $n$.
+For the base case, assume $n=0$, so $t_{0}=t_{n}$ is the only color that appears directly above $b$ after $c$ is sorted.
+Since $c$ follows the last $b$ and $b$ can never appear as the top color on the stack after $c$ is sorted, the color $t_{0}=t_{n}$ must appear above $b$ after $c$ is sorted.
 Therefore, we have
@@
 Let $(S^{\prime},R^{\prime}):=\theta_{d}(X)$.
-Note, during the execution of $\theta_{d}$, the color $b$ is never the top element of the stack and that $t_{0},\dotsc,t_{n-1}$ is the sequence of colors that appear immediately above the color $b$ after $c$ is sorted.
+Note, during the execution of $\theta_{d}$, the color $b$ is never the top color on the stack and that $t_{0},\dotsc,t_{n-1}$ is the sequence of colors that appear immediately above the color $b$ after $c$ is sorted.
 Therefore, by induction, we have
@@
 $C\supseteq I(c,t_{0},\dotsc,t_{n-1})R^{\prime}\supseteq I(c,t_{0},\dotsc,t_{n-1})t_{n}t_{n-1}R=I(c,t_{0},\dotsc,t_{n-1},t_{n})R.\qed$
+For the final statement, note that, for every $i\in\{0,\dotsc,n\}$, the color $t_{i}\neq b$ by definition and $t_{i}\neq c$ since it appears in the stack after $c$ is sorted.
+Furthermore, the colors $t_{0},\dotsc,t_{n}$ must be $n+1$ distinct colors because the only way for a color to removed from the stack is for it to be sorted.
 We now present a few simple observations and lemmas that will be useful throughout.
@@
 ∎
-Observation 23.
-Let $X$ be a sock ordering that contains a sandwich.
-If $aba$ is the first sandwich to appear in $X$, then any color that is not $a$ or $b$ is either sortable in $X$ or is blocked by the sandwich $aba$.
+There is a natural ordering of the sandwiches that appear in a given sock ordering.
+If $aba$ and $xyx$ are suborderings of a sock ordering $X$ we say that $aba$ precedes $xyx$ if the final $a$ in $aba$ precedes the final $x$ in $xyx$; or when $a=x$ and the final socks in $aba$ and $xyx$ are the same, the $b$ in $aba$ precedes the $y$ in $xyx$; or when $a=x$ and $b=y$ and the last two sock in $aba$ and $xyx$ are the same, the first $a$ in $aba$ precedes the first in $xyx$.
+Observation 6.
+If $aba$ is the first sandwich to appear in a sock ordering $X$, then any color that is not $a$ or $b$ is either sortable in $X$ or is blocked by the sandwich $aba$.
 Proof.
@@
 Lemma 7.
-Let $S$ and $R$ be strings where $S$ contains at least two distinct colors and does not contain a sandwich.
+Let $S$ and $R$ be strings where $S$ is sorted and has at least two distinct colors.
 Let $a\in S$ where $a$ is not the top color in $S$ and let $S^{\prime}$ be formed by removing the color $a$ from $S$.
@@
   • $ax\subseteq R$, or
-  • there exists a color $y$ distinct from the colors $a$ and $x$ such that $yx\subseteq R$ and $y$, $a$, and $x$ are the top three colors on the stack $S$, respectively
+  • there exists a color $y$ distinct from the colors $a$ and $x$ such that $yx\subseteq R$ and $y$ is the top color, $a$ is immediately below the color $y$, and $x$ is immediately below the color $a$ on the stack $S$.
 Proof.
@@
 Let $yzy$ be the first sandwich in $SR$ that blocks $x$.
-Since $S$ does not contain $yzy$, we have $yx\subseteq R$.
+Since $S$ is sorted, $yzy$ is not a subordering of $S$, so we have $yx\subseteq R$.
 Because $x$ is sortable in $S^{\prime}R$, the color $a$ is either $y$ or $z$.
@@
 Therefore, $x$ is the top color on $S$.
-We now have that the colors $a$ and $x$ are the top two colors in $S$, respectively, and $yx\subseteq R$.
+We now have that the colors $a$ and $x$ are the top two colors on $S$, respectively, and $yx\subseteq R$.
 Therefore, we only need to show that the color $y$ is directly below the color $a$ in $S$.
-This follows from the observation that if there exists $u$ that is distinct from $y$, $a$, or $x$ that is between $y$ and $a$ in $S$, then the sandwich $yuy$ blocks $x$ in $S^{\prime}R$, a contradiction.
+This follows from the observation that if there exists a color $u$ that is distinct from $y$, $a$, or $x$ that is between $y$ and $a$ in $S$, then the sandwich $yuy$ blocks $x$ in $S^{\prime}R$, a contradiction.
 ∎
-The following lemma establishes some general conditions that imply an unsortable sock ordering must either contain a proper subordering that is unsortable or contain $I_{1,n}$ for some $n\geq 0$.
-This lemma is a crucial piece of our recursive algorithm, because if (I) and (IV) do not hold, then we can determine if a sock ordering is sortable and how to sort it by considering a proper subordering.
+The following lemma is a crucial piece of our inductive proof, because if (I) and (IV) do not hold for a sock ordering $X$ which satisfies the hypothesis, then, because (III) implies that $X$ is sortable, we can determine if a sock ordering is sortable and exactly how to sort $X$ by considering a proper subordering.
 Lemma 8.
-Let $X$ be a sock ordering and let $a$, $x$, and $d$ be distinct colors in $X$.
-Let $\overline{\theta}$ be a nonempty sorting sequence for $X$ such that $\overline{\theta}(X)=(S,R)$, where $ax\subseteq S$, $a\in R$, and $d$ is the last color sorted in $\overline{\theta}$.
-Further suppose that there exists $S^{\prime}\subseteq S$ and $R^{\prime}\subseteq R$ such that $ax\subseteq S^{\prime}$ and $a$ is the only good sortable color in $S^{\prime}R^{\prime}$.
-One of the following is true:
+Let $X$ be a sock ordering and let $a$ and $d$ be distinct colors in $X$.
+Let $\overline{\theta}$ be a nonempty sorting sequence for $X$ and let $(S,R):=\overline{\theta}(\emptyset,X)$.
+Let $d$ be the last color in $\overline{\theta}$.
+Furthermore, assume there exists a substack $S^{\prime}\subseteq S$ and subordering $R^{\prime}\subseteq R$ such that the color $a$ is beneath the top color on $S^{\prime}$, $a$ appears in $R^{\prime}$, and $a$ is the only good sortable color in $S^{\prime}R^{\prime}$.
+Then, one of the following is true:
   1. (I)
@@
   3. (III)
-    $X-d$ is sortable and if a sequence $\theta$ fully sorts $X-d$, then $\theta$ also fully sorts $X$; or
+    $X-d$ is sortable and if a sequence $\theta$ fully sorts $X-d$, then when we attempt to sort $X$ with the sequence $\theta$, there is a stage where the color $d$ only appears as the top color on the stack (this condition implies that $X$ is sortable); or
   4. (IV)
-    $I_{1,n}\preceq X$ for some $n\geq 0$.
+    for some $n\geq 0$, $I_{1,n}$ is a subpattern of $X$.
 Proof.
@@
 We will show that (IV) must hold.
-Let $X=YDW$, where the first element of $D$ is the first $d$ in $X$ and the final element of $D$ is the final $d$ in $X$.
-Note that $R\subseteq W$, so $a\in W$.
-Because there are no $d$-sandwiches that block $a$ in $X$, the only colors that could appear in $D$ are $a$ and $d$.
-Since $d$ is sorted by $\overline{\theta}$ but $a$ and $x$ are not, there is no sandwich that consists only of $a$ and $x$ that blocks $d$ in $X$.
-Therefore, $D$ cannot contain $a$ as otherwise a sandwich of the form $axa$ would block $d$ in $X$.
-As such, $d$ is the only color that appears in $D$.
-Since $d$ is the last color in $\overline{\theta}$, this implies that $Y\supseteq S\supseteq S^{\prime}\supseteq ax$.
-Since, by assumption, the statements (II) and (III) are both false, there exists a sorting sequence $\theta$ that fully sorts $X-d=YW$ such that $\theta$ does not fully sort $X=YDW$.
+Let $x$ be the top color on $S^{\prime}$ and $D\subseteq X$ be such that $X=YDW$ where the first element of $D$ is the first $d$ in $X$ and the final element of $D$ is the final $d$ in $X$.
+Note that the subordering $R$ must follow the last $d$ in $X$, so $R\subseteq W$.
+Similarly, everything in $S$ must precede the last $d$ in $X$, so $S\subseteq YD$.
+Since $R\subseteq W$ and $a\in R^{\prime}\subseteq R$, we have $a\in W$, so the fact there are no $d$-sandwiches that block $a$ in $X$ implies that the only colors that could appear in $D$ are $a$ and $d$.
+In particular, the color $x$ does not appear in $D$, so the fact that $a$ is beneath $x$ on $S^{\prime}$ and $S^{\prime}\subseteq S\subseteq YD$ implies that $ax\subseteq Y$.
+Note that because $d$ is sorted by $\overline{\theta}$ but $a$ and $x$ are not, there is no sandwich that consists only of $a$ and $x$ that blocks $d$ in $X$.
+Therefore, the color $a$ does not appear in $D$ as otherwise the fact that $ax\subseteq Y$ would imply the existence of a sandwich of the form $axa$ before the final $d$ in $X$.
+As such, $d$ is the only color that appears in $D$ and we can assume without loss of generality that the color $d$ only appears once in $X$, so $D=d$.
+By assumption, the statements (II) and (III) are both false, so there exists a sorting sequence $\theta$ that fully sorts $X-d=YW$ such that when we apply $\theta$ to $X=YDW=YdW$ the color $d$ never appears only as the top color on the stack.
 Let $\theta_{a}$ be the list of initial colors in $\theta$ up to, but not including, $a$ (so $\theta_{a}$ is empty if and only if $\theta$ begins with $a$).
-Note that no color in $S^{\prime}R^{\prime}\subseteq X$ could appear before $a$ in $\theta$ because $a$ is the only good sortable color in $S^{\prime}R^{\prime}$.
-This means that no color that appears after $ax$ in $X$ is in $\theta_{a}$.
+Note that no color in $S^{\prime}R^{\prime}$ could appear before $a$ in $\theta$ because $a$ is the only good sortable color in $S^{\prime}R^{\prime}\subseteq YW$ and $\theta$ fully sorts $YW$.
+We claim that this implies that no color that appears after $ax$ in $YW$ is in $\theta_{a}$.
 To see this, note that sorting such a color would put $ax$ onto the stack, but that would make it impossible to sort $a$ before sorting $x$.
-In particular, this means that no color that appears in $W$ is in $\theta_{a}$.
-Therefore, $a$ is the color in $W$ that is sorted first by $\theta$ and $\theta_{a}$ is a sorting sequence for $X=YDW$.
-Furthermore, if we let $(S_{1},R^{\prime}_{1})=\theta_{a}(YW)$, then $(S_{1},R_{1})=\theta_{a}(YDW)$ where $R^{\prime}_{1}$ is formed by removing $d$ from $R_{1}$.
+In particular, because $ax\subseteq Y$, this means that no color that appears in $W$ is in $\theta_{a}$.
+Therefore, $a$ is the color in $W$ that is sorted first by $\theta$, and $\theta_{a}$ is a sorting sequence for $X=YdW$.
+Furthermore, if we let $(S_{1},R^{\prime}_{1})=\theta_{a}(YW)$, then $(S_{1},R_{1})=\theta_{a}(YdW)$ where $R^{\prime}_{1}=R_{1}-d$.
 Therefore, because $a$ is sortable in $(S_{1},R^{\prime}_{1})$ and no $d$-sandwich blocks $a$ in $X$, the color $a$ is also sortable in $(S_{1},R_{1})$.
-Recall that no color that appears after $ax$ is in $\theta_{a}$.
+Recall that no color that appears after $ax$ is in $\theta_{a}$ and $x$ is not in $\theta_{a}$.
 This, with the fact that $axda\subseteq SdR\subseteq X$, implies that $xda\subseteq R_{1}$.
-Therefore, sorting $a$ in $(S_{1},R_{1})$ pushes $xD$ onto the stack.
-Let $z$ be the first color in $\theta$ that causes a sandwich to appear in the stack when $\theta$ is applied to $X$.
-Such a color $z$ must exist because $\theta$ does not fully sort $X$.
-Let $\theta_{z}$ be the list of initial colors in $\theta$ up to but not including $z$.
-Note that by our previous arguments, $a$ must appear in $\theta_{z}$ and $a$ is the color in $W$ that is sorted first by $\theta_{z}$.
-Furthermore, since $xD$ is pushed onto the stack when $a$ is sorted and $d\notin\theta_{z}$, if we let $(S^{\prime}_{2},R_{2})=\theta_{z}(YW)$ we have $(S_{2},R_{2})=\theta_{z}(YDW)$ where $xD\subseteq S_{2}$ and $S^{\prime}_{2}=S_{2}-d$.
-Note that while applying $\theta_{z}$ to $X$, after $a$ is sorted the color $d$ can never appear as the top element of the stack since $\theta$ fully sorts $X-d$ but does not not fully sort $X$.
-Let $t_{0},\dotsc,t_{n}$ be the $n+1$ colors that appear immediately above $d$ on the stack after $a$ is sorted by $\theta_{z}$.
-Since $X=YDW$ and $D$ consists solely of one or more copies of $d$, Lemma 21 implies that $dI(a,t_{0},\dotsc,t_{n})R_{2}\subseteq dW$.
-Recall that $xD\subseteq S_{2}$.
-Therefore, there exists a color $y$ that appears directly below $D$ in $S_{2}$.
-(We could have $y=x$.) Since $axDa\subseteq X$, we have $ayDa\subseteq X$, so $aydI(a,t_{0},\dotsc,t_{n})R_{2}\subseteq YdW\subseteq X$.
-Note that, by the selection of $z$, the color $z$ is unsortable in $(S_{2},R_{2})$, but $z$ is sortable in $(S^{\prime}_{2},R_{2})$.
-Since $d\notin S^{\prime}_{2}R_{2}\subseteq X-d$, we have $z\neq d$ and $dz\nsubseteq R_{2}$.
-Therefore, Lemma 24 implies that $y$, $d$, and $z$ are the top three colors in the stack $S_{2}$ and that $yz\subseteq R_{2}$.
-Note that this means that $z=t_{n}$ since $t_{n}$ is defined to be the last color that appears directly above $d$ on the stack when sorting $X$ with $\theta_{z}$.
-Therefore,
-$X\supseteq aydI(a,t_{0},\dotsc,t_{n})R_{2}\supseteq aydt_{0}aI(t_{0},\dotsc,t_{n})yz=aydt_{0}aI(t_{0},\dotsc,t_{n})yt_{n}\cong I_{1,n}.\qed$
+Therefore, sorting $a$ in the state $(S_{1},R_{1})$ pushes $xd$ onto the stack.
+Let $y$ be the color that immediately precedes $d$ in $R_{1}$ (so we could have $y=x$).
+Note that since $ax\subseteq Y$ and either $x=y$ or $xy\subseteq Y$, in all cases, we have
+$ay\subseteq Y.$ (1)
+Recall that, by the previous arguments, $d$ is pushed on the stack when we attempt to apply $\theta$ to $X$, so $\theta$ cannot be a sorting sequence for $X$ as otherwise we would have $(d,\emptyset)=\theta(X)$, a contradiction to the fact that the color $d$ never appears as the top color on the stack.
+Therefore, there must exist a color in $\theta$ that causes the stack to be unsorted when $\theta$ is applied to $X$.
+Let $z$ be the first such color and let $\theta_{z}$ be the list of initial colors in $\theta$ up to but not including $z$.
+That is, $\theta_{z}$ is the longest initial sequence of $\theta$ that is a sorting sequence for $X$.
+Let $(S_{2},R_{2}):=\theta_{z}(X)$.
+Note that by our previous arguments, $a$ must appear in $\theta_{z}$ and $a$ is the color in $W$ that is sorted first by $\theta_{z}$ and $d\notin W$.
+By Lemma 4, we have
+$W\supseteq I(a,t_{0},\dotsc,t_{n})R_{2}=t_{0}aI(t_{0},\dotsc,t_{n})R_{2},$ (2)
+where $t_{0},\dotsc,t_{n}$ is the sequence of colors that appear above $d$ on the stack after $a$ is sorted.
+Note that since $y$ is immediately below $d$ on the stack after the sequence $\theta_{a}$ is applied to $X$ and $d$ is never removed from the stack while $\theta_{z}$ is applied to $X$, the color $y$ is immediately below $d$ on the stack $S_{2}$.
+Let $S_{2}^{\prime}:=S_{2}-d$ and note that $(S_{2}^{\prime},R_{2})=\theta_{z}(X-d)$.
+By the selection of $z$, the color $z$ is unsortable in $(S_{2},R_{2})$, but $z$ is sortable in $(S^{\prime}_{2},R_{2})$.
+Since $d\notin S^{\prime}_{2}R_{2}\subseteq X-d$, we have $z\neq d$ and $dz\nsubseteq R_{2}$, so Lemma 7 implies that $y$, $d$, and $z$ are the top three colors in the stack $S_{2}$ and that $yz\subseteq R_{2}$.
+Note that this means that $z=t_{n}$, since $t_{n}$ is defined to be the last color that appears immediately above $d$ on the stack when we sort $X$ with $\theta_{z}$.
+So, $yt_{n}\subseteq R_{2}$.
+This, with (1) and (2), implies that
+$X=YdW\supseteq aydt_{0}aI(t_{0},\dotsc,t_{n})R_{2}\supseteq aydIt_{0}a(t_{0},\dotsc,t_{n})yt_{n}\cong I_{1,n}.\qed$
 The following three lemmas are needed in the proof of our main theorem to show that a given unsortable sock ordering has a subpattern that is in $\Gamma$.
-Unsurprisingly, these proofs have a number of cases and are somewhat technical, so we defer their proofs until Section 6.
-In Lemma 27 and 28, we only need the first conclusion “$X$ is unsortable”, to prove that the algorithm is correct.
-The second conclusion “a subpattern of $X$ is in $\Gamma$” is needed in our inductive proof of Theorem 1.
+Unsurprisingly, these proofs have a number of cases and are somewhat technical, so we defer their proofs until Section 2.6.
 Lemma 9.
@@
 If $abt\subseteq S$ and either
-  • $atb$ or $abt$ is contained in $R$ and $a$ is blocked by a $d$-sandwich in $X$; or
+  • $atb\subseteq R$ or $abt\subseteq R$ and $a$ is blocked by a $d$-sandwich in $X$; or
   • $bI(t_{0},\dotsc,t_{n})at_{n}\subseteq R$ and $b$ is blocked by a $d$-sandwich in $X$,
-then $X$ is unsortable. If, in addition, every strict subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
+then $X$ is unsortable. If, in addition, every proper subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
 Lemma 11.
-Let $X$ be a sock ordering and suppose that a nonempty sorting sequence produces $(S,R)$ when it is applied to $X$, where $d$ is the last color sorted by the sequence.
+Let $X$ be a sock ordering and suppose that a nonempty sorting sequence produces $(S,R)$ when it is applied to $X$ where $d$ is the last color sorted by the sequence.
 Further suppose that the following two conditions hold:
@@
   • There exist distinct colors $a$ and $t$ in $S$ such that $(at,R)$ is terminal.
-Then there exists $P\subseteq R$ such that $a$ is the only good sortable color in $atP$.
+Then there exists $P\subseteq R$ such that $a$ appears in $P$ and $a$ is the only good sortable color in $atP$.
 Furthermore, if $a$ is blocked by a $d$-sandwich in $X$, then $X$ is unsortable.
-If, in addition, every strict subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
+If, in addition, every proper subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
 The following lemma shows that if we reach the state $(S,R)$ that is terminal and $S$ has at least three colors, then, in a certain sense, we only need to focus on at most three colors in $S$.
-Furthermore, two of the three colors can be the top two colors on $S$ (cf.
-Algorithm 3, Line 8)
+Furthermore, two of the three colors can be the top two colors on $S$.
 Lemma 12.
-Let $S$ and $R$ be sock orderings where $S$ does not contain a sandwich and $S$ contains at least three colors.
+Let $S$ and $R$ be sock orderings where $S$ is sorted and at least three distinct colors appear in $S$.
 Suppose $(S,R)$ is terminal.
@@
 Let $p$ be as small as possible such that there exists $y_{1},\dotsc,y_{p}$, a subsequence of $x_{1},\dotsc,x_{m}$, such that $(y_{1}y_{2}\dotsm y_{p}bt,R)$ is terminal.
-Since $(S,R)$ is a counterexample, we have $p\geq 2$, and if we let $W$ be the nonempty string $y_{3}\dotsm y_{p}b$, then for $i\in\{1,2\}$, there exists some $z_{i}$ that is sortable in $(y_{i}Wt,R)$.
-For $i\in\{1,2\}$, since $z_{i}$ is sortable in $(y_{i}Wt,R)$, we have $z_{i}\notin y_{i}W$ and $y_{i}z_{i}\nsubseteq R$.
-Since $z_{1}\neq y_{1}$, $y_{1}z_{1}\nsubseteq R$, and $z_{1}$ is not sortable in $(y_{1}y_{2}Wt,R)$, Lemma 24 implies that $z_{1}$ is unsortable in $(y_{2}Wt,R)$.
-Therefore, because $y_{1}z_{1}\nsubseteq R$ and $z_{1}\notin W$, either $z_{1}=y_{2}$ or there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ that does not contain $y_{1}$.
-If $z_{1}=y_{2}$, then, because $z_{1}=y_{2}$ is sortable in $(y_{1}Wt,R)$, we have $y_{2}\in R$, so $y_{1}y_{2}=y_{1}z_{1}\nsubseteq R$ implies that $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
-If there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ that does not contain $y_{1}$, then $y_{1}z_{1}\nsubseteq R$ implies that $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
-So, in both cases, $y_{1}$ is not sortable in $(y_{2}Wt,R)$, so $y_{1}\neq z_{2}$, and Lemma 24 then implies that $y_{1}z_{2}\subseteq R$.
-Since $y_{1}z_{2}\subseteq R$ and $y_{2}z_{2}\nsubseteq R$, $y_{2}$ is not sortable in $(y_{1}Wt,R)$.
-Therefore, $y_{2}\neq z_{1}$.
-This together with Lemma 24 implies that $y_{2}z_{1}\subseteq R$.
+Since $(S,R)$ is a counterexample, we have $p\geq 2$.
+If $p=2$, let $W=b$ and otherwise let $W=y_{3}\dotsm y_{p}b$.
+By the minimality of $p$, for $i\in\{1,2\}$, there exists some $z_{i}$ that is sortable in $(y_{i}Wt,R)$.
+For $i\in\{1,2\}$, since $z_{i}$ is sortable in $(y_{i}Wt,R)$, we have $z_{i}\notin y_{i}W$.
+In particular, this implies that $z_{i}\neq y_{i}$ and $z_{i}\neq b$.
+Furthermore, since $z_{i}$ is sortable in $(y_{i}Wt,R)$, there is no $y_{i}by_{i}$ sandwich blocking $z_{i}$ in $y_{i}WtR$, so $y_{i}z_{i}\nsubseteq R$.
+We will now show that $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
+Recall that $(y_{1}y_{2}Wt,R)=(y_{1}y_{2}\dotsm y_{p}bt,R)$ is terminal, so $z_{1}$ is not sortable in $(y_{1}y_{2}Wt,R)$.
+Therefore, we can apply Lemma 7 to deduce that $z_{1}$ is unsortable in $(y_{2}Wt,R)$, because $z_{1}\neq y_{1}$, $y_{1}z_{1}\nsubseteq R$, and $y_{2}$ is not the second color on the stack $y_{1}y_{2}Wt$.
+Since $z_{1}$ is unsortable in $(y_{2}Wt,R)$, $y_{1}z_{1}\nsubseteq R$, and $z_{1}\notin W$, either $z_{1}=y_{2}$ or there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ in which $y_{1}$ does not appear.
+If $z_{1}=y_{2}$, then, because $z_{1}=y_{2}$ is sortable in $(y_{1}Wt,R)$, we have $y_{2}\in R$, so $y_{1}y_{2}=y_{1}z_{1}\nsubseteq R$ implies that either $y_{1}\notin R$ or that $y_{1}$ is blocked by a sandwich of the form $y_{2}by_{2}$ in $y_{2}WtR$.
+So, $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
+If there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ in which $y_{1}$ does not appear, then $y_{1}z_{1}\nsubseteq R$ implies that $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
+So, in all cases, $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
+Since $y_{1}$ is not sortable in $(y_{2}Wt,R)$, we have $y_{1}\neq z_{2}$.
+This, with the fact that $z_{2}$ is not sortable in $(y_{1}y_{2}Wt,R)$ but is sortable $(y_{2}Wt,R)$ together with Lemma 7, yields $y_{1}z_{2}\subseteq R$.
+Since $y_{1}z_{2}\subseteq R$ and $y_{2}z_{2}\nsubseteq R$, $y_{2}$ is not sortable in $(y_{1}Wt,R)$, so $y_{2}\neq z_{1}$.
+Using this fact with Lemma 7, allows us to conclude that $y_{2}z_{1}\subseteq R$ because $z_{1}$ is not sortable in $(y_{1}y_{2}Wt,R)$ but is sortable $(y_{1}Wt,R)$.
 Since $y_{1}z_{2}\subseteq R$, $y_{2}z_{2}\nsubseteq R$, and $y_{2}z_{1}\subseteq R$, we have $y_{1}z_{1}\subseteq R$.
@@
 Our last lemma is simple despite the somewhat technical statement.
-It is used in the case when a color, say $x$, is not sorted by the greedy algorithm despite terminating before the color which is sorted by the greedy algorithm, say $y$.
+Recall the greedy algorithm described in Section 2.3.
+This lemma is used in the case when a color, say $x$, is not sorted by the greedy algorithm at a particular stage despite terminating before another color, say $y$, which is sorted by the greedy algorithm at that stage.
 Note that, in this case, if a sandwich blocks $x$, then it must be a $y$-sandwich because $y$ is sortable.
@@
 Lemma 13.
-Let $S$ and $R$ be sock orderings where $S$ does not contain a sandwich and let $x$ and $y$ be distinct colors in $R$ such that the final $x$ in $R$ appears before the final $y$ in $R$.
+Let $S$ and $R$ be sock orderings where $S$ is sorted and let $x$ and $y$ be distinct colors in $R$ such that the final $x$ in $R$ appears before the final $y$ in $R$.
 Suppose that, with respect to $(S,R)$, the color $y$ is sortable and the color $x$ is unsortable.
 If no $y$-sandwich blocks $x$ in $SR$, then $x$ and $y$ are the top two colors on $S$.
-Furthermore, no color except possibly $y$ precedes $x$ on $R$.
+Furthermore, no color except possibly $y$ precedes $x$ in $R$.
 Proof.
@@
 Therefore, because the final $x$ in $R$ precedes the final $y$ in $R$ and $x$ is not blocked by a $y$-sandwich in $SR$, the color $x$ is not blocked by a sandwich in $SR$.
-The fact that $x$ is unsortable then implies that $x$ is below the top element on the stack $S$.
+The fact that $x$ is unsortable then implies that $x$ is below the top color on the stack $S$.
 Since $y$ is sortable in $SR$, there is no $x$-sandwich that blocks $y$ in $SR$.
 Therefore, the colors $x$ and $y$ are the top two colors on $S$.
-Furthermore, no color except possibly $y$ can precede $x$ on $R$.
+Furthermore, no color except possibly $y$ can precede $x$ in $R$.
 ∎
-5. Proof of Theorem 1
-The proof is by induction on $|X|$.
+2.5. Proof of Theorem 3
+The proof is by induction on the number of socks in $X$.
 The base case is vacuously true.
-So, by the induction hypothesis, we can assume that there does not exist a proper subordering of $X$ that is unsortable.
-Note that if we have colors $a$, $x$ and $d$ and substrings $S^{\prime}\subseteq S\subseteq X$ and $R^{\prime}\subseteq R\subseteq X$ that satisfy the hypothesis of Lemma 25, then we can assume (I) holds.
+So assume that $X$ is an unsortable sock ordering.
+With the induction hypothesis, we can assume that there does not exist a proper subordering of $X$ that is unsortable.
+Note that if we have colors $a$ and $d$ and substrings $S^{\prime}\subseteq S\subseteq X$ and $R^{\prime}\subseteq R\subseteq X$ that satisfy the hypothesis of Lemma 8, then we can assume Lemma 8(I) holds.
 That is, we can assume that $a$ is blocked by a $d$-sandwich.
-This is because (II) and (III) in Lemma 25 cannot hold, and if (IV) holds, then we satisfy the conclusion.
-This will make it easy to apply either Lemma 27 or 28 to complete the proof.
-Let $(S,R):=\phi_{G}(X)$ be the terminal state of the greedy algorithm and let $d$ be the last color sorted by $\phi_{G}$. Since $X$ is unsortable, we have $R\neq\emptyset$.
-Case 1: $S=\emptyset$ or there exists $x\in S$ such that $(x,S)$ is unsortable.
-If $S=\emptyset$, then Lemma 26 implies that a subpattern of $R\subseteq X$ is in $\Gamma$.
-Similarly, if there exists $x\in S$ such that $(x,R)$ is terminal, then $xR$ is trivially unsortable, and Lemma 26 implies that a subpattern of $xR\subseteq X$ is in $\Gamma$.
+To see this, note first that Lemma 8(II) cannot hold because every proper subordering of $X$ is sortable.
+Additionally, Lemma 8(III) cannot hold because $X$ is unsortable.
+Furthermore, if Lemma 8(IV) holds, then we satisfy the conclusion.
+This will make it possible to apply either Lemma 10 or 11 to complete the proof.
+Let $(S,R)$ be output of the greedy algorithm (as described in Section 2.3) applied to $X$. This implies that $(S,R)$ is terminal. Since $X$ is unsortable, we have $R\neq\emptyset$.
+Case 1: $S=\emptyset$ or there exists $x\in S$ such that $(x,R)$ is terminal.
+If $S=\emptyset$, then $R$ is trivially unsortable, so Lemma 9 implies that a subpattern of $R\subseteq X$ is in $\Gamma$.
+Similarly, if there exists $x\in S$ such that $(x,R)$ is terminal, then $xR$ is trivially unsortable.
+Lemma 9 then implies that a subpattern of $xR\subseteq X$ is in $\Gamma$.
 From now on, we will assume that we are not in Case 1.
@@
 Define $t$ to be the top color in $S$, and let $b$ be the color in $S$ appearing immediately under $t$.
+We can also now let $d$ be the last color sorted by the greedy algorithm.
 Case 2: $(bt,R)$ is terminal.
-By Lemma 28, there exists $P\subseteq R$ such that $b$ is the only good sortable color in $btP$ and that if $b$ is blocked by a $d$-sandwich in $X$, then a subpattern of $X$ is in $\Gamma$.
-If $b$ is not blocked by a $d$-sandwich in $X$, then we must satisfy Lemma 25(IV).
-That is, $I_{1,n}\preceq X$, for some $n\geq 0$.
-Case 3: $(bt,R)$ is not terminal. With Lemma 29, we can define $a$ to be the lowest color in $S$ such that $(abt,R)$ is terminal.
+By Lemma 11, there exists $P\subseteq R$ such that $b$ appears in $P$ and such that $b$ is the only good sortable color in $btP$.
+This means that the hypothesis of Lemma 8 is satisfied with $b$, $d$, $bt$, and $P$ playing the roles of $a$, $d$, $S^{\prime}$, and $R^{\prime}$, respectively.
+Therefore, we can assume that $b$ is blocked by a $d$-sandwich in $X$.
+But then Lemma 11 implies that a subpattern of $X$ is in $\Gamma$.
+Case 3: $(bt,R)$ is not terminal. With Lemma 12, we can define $a$ to be the lowest color on the stack $S$ such that $(abt,R)$ is terminal.
 Claim 14.
@@
 Suppose, by way of contradiction, that $at\nsubseteq R$.
-Then, because $t$ is unsortable in $(abt,R)$, Lemma 24 implies that $t$ is unsortable in $(bt,R)$.
-Since $(bt,R)$ is not terminal and both $b$ and $t$ are not sortable in $(bt,R)$, there must exist a color $y\in R$ that is not $b$ or $t$ which is sortable in $(bt,R)$.
+Then, because $t$ is unsortable in $(abt,R)$ and $t\neq a$ and $a$ is not the second color on the stack $abt$, Lemma 7 implies that $t$ is unsortable in $(bt,R)$.
+By the case, $(bt,R)$ is not terminal, so since both $b$ and $t$ are not sortable in $(bt,R)$, there must exist a color $y\in R$ that is not $b$ or $t$ which is sortable in $(bt,R)$.
 If $y=a$, then, because $at\nsubseteq R$, the color $y=a$ must follow the last occurrence of $t$ in $btR$.
@@
 By Claim 14, we can assume that either $atz$, $azt$, or $zat=zaI(t)t$ is in $R$.
-But then Lemma 25 and Lemma 27 lead to a contradiction, because $a$ is the only good sortable color in both $aztatz$ and $aztazt$, and $z$ is the only good sortable color in $aztzat$.
+But then Lemma 8 and Lemma 10 give us the desired conclusion, because $a$ is the only good sortable color in both $aztatz$ and $aztazt$, and $z$ is the only good sortable color in $aztzat$.
 Case 3.2: For every $z\in S$ distinct from $a$ and $t$ such that $azt\subseteq S$, we have $z\notin R$.
@@
 Let $A$ and $C$ be such that $X=AbC$ and $b\notin C$.
-Because $b\notin R$ there exists a color in $C$ that is sorted during $\phi_{G}(X)$.
-Let $c\in C$ be the color in $C$ that is sorted first by the greedy algorithm and let $(S^{\prime},A^{\prime}bC)$ be the state immediately before $c$ is sorted by $\phi_{G}(X)$.
+Because $b\in S$ and $b\notin R$, there exists a color in $C$ that is sorted by the greedy algorithm.
+Let $c\in C$ be the color in $C$ that is sorted first by the greedy algorithm and let $(S^{\prime},A^{\prime}bC)$ be the state immediately before $c$ is sorted.
 Note that we have
@@
 $a\in S^{\prime}A^{\prime}$ (4)
-Furthermore, by the definition of the greedy algorithm, the color $b$ can never appear as the top element of the stack after $c$ is sorted by the greedy algorithm.
+Furthermore, since $b$ is not sorted by the greedy algorithm, the color $b$ can never appear as the top color on the stack after $c$ is sorted by the greedy algorithm.
 For $n\geq 0$, let $t_{0},\dotsc,t_{n}$ be the $n+1$ distinct colors that appear immediately above $b$ after $c$ is sorted by the greedy algorithm.
@@
 There are at least three colors in $S^{\prime}$.
-More specifically, if $y$ is the color immediately beneath the top color in $S^{\prime}$, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC$.
+Furthermore, if $y$ is the color immediately beneath the top color in $S^{\prime}$, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC$.
 Proof.
 First assume that $b$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$.
-Then, because the greedy algorithm sorted $c$ and not $b$ at stage $(S^{\prime},A^{\prime}bC)$ and the final $b$ occurs before the final $c$ in $S^{\prime}A^{\prime}bC$, with respect to $(S^{\prime},A^{\prime}bC)$, the color $b$ is unsortable and the color $c$ is sortable.
+Then, because the greedy algorithm sorted $c$ and not $b$ at the stage $(S^{\prime},A^{\prime}bC)$ and the final $b$ occurs before the final $c$ in $S^{\prime}A^{\prime}bC$, with respect to $(S^{\prime},A^{\prime}bC)$, the color $b$ is unsortable and the color $c$ is sortable.
 Then, Lemma 13 implies that $b$ and $c$ are the top two colors on the stack $S^{\prime}$, so $y=b$.
@@
 Now assume that $b$ is blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$.
-Let $\hat{y}$ be the other color in the first $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$; we will show that $\hat{y}=y$ and our conclusion holds.
+Let $\hat{y}$ be the other color in the first $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$.
+We will show that $\hat{y}=y$ and our conclusion holds.
 To see this, first note that we must have $\hat{y}c\subseteq S^{\prime}A^{\prime}\subseteq A$ and $c\hat{y}\subseteq S^{\prime}A^{\prime}\subseteq A$.
@@
 Then, the fact that $b$ is never sorted by the greedy algorithm implies that $\hat{y}$ is never sorted by the greedy algorithm.
-Furthermore, since we cannot have $\hat{y}$-sandwich on the stack, we must have $a\hat{y}bt_{n}\subseteq S$ and $\hat{y}\in R$.
+Furthermore, since we cannot have a $\hat{y}b\hat{y}$-sandwich on the stack, we must have $a\hat{y}bt_{n}\subseteq S$ and $\hat{y}\in R$.
 But this contradicts the fact that we are in Case 3.2.
 Therefore, the final $\hat{y}$ precedes the final $c$ in $S^{\prime}A^{\prime}bC$.
-First note that since $c$ is sortable with respect to $(S^{\prime},A^{\prime}bC)$, the color $c$ is not below the top color on the stack $S^{\prime}$.
-Since there is a sandwich consisting of $\hat{y}$ and $c$ in $S^{\prime}AbC$, we must have $\hat{y}$ in $A^{\prime}bC$.
-Also, by the definition of $c$, we have $c\in C\subseteq A^{\prime}bC$.
-Since the greedy algorithm sorts $c$ and not $\hat{y}$ when $(S^{\prime},A^{\prime}bC)$ appears, $c$ is sortable and $\hat{y}$ is unsortable with respect to $(S^{\prime},A^{\prime}bC)$.
-Furthermore, since the selection of $\hat{y}$ implies that the color $\hat{y}$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$, Lemma 30 implies that $\hat{y}$ and $c$ are the top two colors in $S^{\prime}$.
-Note that we have shown that $\hat{y}=y$.
-Furthermore, Lemma 30 implies that and no color except possibly $c$ precedes $\hat{y}=y$ in $A^{\prime}bC$.
-In particular, $b$ does not precede $\hat{y}=y$ in $A^{\prime}bC$, so $\hat{y}=y\in A^{\prime}$.
-Then, with (3),
+Recall that $\hat{y}$ is the color other than $c$ that is in the first $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$.
+Since $b\notin C$, we have $\hat{y}b\subseteq S^{\prime}A^{\prime}b$.
+Because there is no sandwich consisting of $b$ and $\hat{y}$ which blocks $c$ in $S^{\prime}A^{\prime}bC$ and both the final $b$ and the final $\hat{y}$ in $S^{\prime}A^{\prime}bC$ preceed the final $c$ in $S^{\prime}A^{\prime}bC$, this implies that $b\hat{y}\nsubseteq S^{\prime}A^{\prime}bC$.
+In particular, we have $\hat{y}\notin C$.
+We claim that the color $\hat{y}$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$.
+For a contradiction, assume that $\bar{y}$ is the other color in such a sandwich.
+Since $\hat{y}\notin C$, this sandwich is in $S^{\prime}A^{\prime}b$, and $\bar{y}\hat{y}\subseteq S^{\prime}A^{\prime}b$.
+Since $b\hat{y}\nsubseteq S^{\prime}A^{\prime}bC$, this implies that $\bar{y}\neq b$.
+Therefore, there is a sandwich consisting of $\bar{y}\neq b$ and $c$ in $S^{\prime}A^{\prime}b$, so this sandwich blocks $b$ as well as $\hat{y}$.
+Furthermore, since there is no sandwich consisting of $\bar{y}$ and $\hat{y}$ that blocks $c$ in $S^{\prime}AbC$, we have $\hat{y}\bar{y}\nsubseteq S^{\prime}Ab$.
+But this contradicts the fact that $\hat{y}$ is in the first $c$-sandwich in $S^{\prime}A^{\prime}bC$ that blocks $b$.
+So we have established that the color $\hat{y}$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$.
+Recall that the final $\hat{y}$ precedes the final $c$ in $S^{\prime}A^{\prime}bC$, so Lemma 13 implies that $c$ is on top of the stack $S^{\prime}$ and the color $\hat{y}$ is immediately below the color $c$ on $S^{\prime}$.
+Note that we now have $\hat{y}=y$ since $y$ was defined to be the color below the top color on $S^{\prime}$.
+Since $y=\hat{y}$ is in a $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$ this also implies that $y=\hat{y}\in A^{\prime}$.
+Furthermore, with (5),
 $A^{\prime}bC\supseteq ybI(c,t_{0},\dotsc,t_{n})at_{n}\supseteq yI(c,t_{0},\dotsc,t_{n})at_{n}.$
-Since (1), (2), and (4), all hold and $y$ and $c$ are the top two colors in $S^{\prime}$, we have $ayc\subseteq S^{\prime}$. ∎
-Defined $d^{\prime}$ to be the color sorted by the greedy algorithm immediately before reaching the state $(S^{\prime},A^{\prime}bC)$.
-By Claim 32, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,)at_{n}\subseteq A^{\prime}bC$.
-Note that, with Observation 20, it is not hard to see that $y$ is the only good sortable color in $aycyI(c,t_{0},\dotsc,)at_{n}$.
-If $y$ is not blocked by a $d^{\prime}$-sandwich in $X$, then Lemma 25 implies that a subpattern of $X$ is in $\Gamma$.
-Otherwise, Lemma 27 implies that a subpattern of $X$ is in $\Gamma$.
-6. Proofs of Lemmas 26, 27, and 28
-See 26
+Since (3), (4), and (6), all hold and $y$ and $c$ are the top two colors on $S^{\prime}$, we have $ayc\subseteq S^{\prime}$. ∎
+Define $d^{\prime}$ to be the color sorted by the greedy algorithm immediately before reaching the state $(S^{\prime},A^{\prime}bC)$.
+By Claim 15, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC$.
+Note that $y$ and $c$ are the only sortable colors in $aycyI(c,t_{0},\dotsc,t_{n})at_{n}$, because every other color is blocked by a $ycy$ sandwich.
+Furthermore, after sorting $c$ we have $(ayyt_{0},I(t_{0},\dotsc,t_{n})at_{n})$ which Observation 2 implies is unsortable.
+Therefore, with Observation 2, $y$ is the only good sortable color in $aycyI(c,t_{0},\dotsc,t_{n})at_{n}$.
+This means that the hypothesis of Lemma 8 is satisfied with $y$, $d^{\prime}$, $ayc$, and $yI(c,t_{0},\dotsc,t_{n})at_{n}$ playing the roles of $a$, $d$, $S^{\prime}$ and $R^{\prime}$, respectively.
+Therefore, we can assume that $y$ is blocked by a $d^{\prime}$-sandwich in $X$.
+Then Lemma 10 implies that a subpattern of $X$ is in $\Gamma$.
+2.6. Proofs of Lemmas 9, 10, and 11
+See 9
 Proof.
@@
 Recall that since $X$ is trivially unsortable, every color is blocked by a sandwich.
-That is, for all $z\in X$, there exist distinct colors $x,y$ in $X$ which are distinct from $z$ such that $xyxz\subseteq X$.
+That is, for all $z\in X$, there exist distinct colors $x$ and $y$ in $X$ which are distinct from $z$ such that $xyxz\subseteq X$.
 This implies that every color must follow the first sandwich $aba$.
@@
 Then $xyx=aya$ blocks $z$, so $z=b$.
-Suppose the $y$ in the sandwich $aya$ follows the first sandwich $aba$, then $X$ contains $abayaba\cong abacaba\cong\mathcal{T}_{1}$.
-Otherwise, because $aba$ (and not $aya$) is the first sandwich, $X$ must contain $abya$ which is then followed by $ba$ and the color $y$.
-Therefore, $X$ contains either $abyabya\cong abcabca\cong\mathcal{T}_{3}$, $abyayba\cong abcacba\cong\mathcal{T}_{4}$, or $abyabay\cong abcabac\cong\mathcal{T}_{5}$.
+Suppose the $y$ in the sandwich $aya$ follows the first sandwich $aba$, then $X\supseteq abayaba\cong abacaba\cong\mathcal{T}_{1}$.
+Otherwise, because $aba$ (and not $aya$) is the first sandwich, we have $X\supseteq abya$.
+This $abya$ subordering is followed by $ba$ and the color $y$ in $X$.
+Therefore, either $X\supseteq abyabya\cong abcabca\cong\mathcal{T}_{3}$, $abyayba\cong abcacba\cong\mathcal{T}_{4}$, or $X\supseteq abyabay\cong abcabac\cong\mathcal{T}_{5}$.
 Case 2: $x=b$.
@@
 Suppose the $y$ in the sandwich $byb$ follows the first sandwich $aba$.
-Then $X$ contains $abaybab\cong abacbab\cong\mathcal{T}_{2}$.
-Otherwise, $X$ contains $abya$.
-Since $byb$ blocks $z=a$ and the final $b$ follows the final $z=a$, the pattern $bab$ must follow this copy of $abya$.
+Then $X\supseteq abaybab\cong abacbab\cong\mathcal{T}_{2}$.
+Otherwise, because $aba$ is the first sandwich (and not $aya$), we have $X\supseteq abya$.
+Since $byb$ blocks $z=a$ and the final $b$ follows the final $z=a$, the subordering $bab$ must follow the subordering $abya$.
 Furthermore, since $y$ must follow the first sandwich, the color $y$ must follow this $abya$.
 So, as in the previous case, $abya$ is followed by $ba$ and the color $y$.
-Therefore, as before, either $\mathcal{T}_{3}$, $\mathcal{T}_{4}$, or $\mathcal{T}_{5}$ is a subpattern of $X$.
+Therefore, as in the previous case, either $\mathcal{T}_{3}$, $\mathcal{T}_{4}$, or $\mathcal{T}_{5}$ is a subpattern of $X$.
 Case 3: $x\neq a$ and $x\neq b$.
@@
 Then, because $aba$ is the first sandwich to appear, the color $x$ must also follow $aba$.
-Therefore, $X$ contains either $xabaxab\cong abcbabc\cong\mathcal{T}_{6}$ or $xabaxba\cong abcbacb\cong\mathcal{T}_{7}$.
+Therefore, $X\subseteq xabaxab\cong abcbacb\cong\mathcal{T}_{6}$ or $X\subseteq xabaxba\cong abcbabc\cong\mathcal{T}_{7}$.
 Case 3.2: The first $x$ appears between the first and last $a$ in the first sandwich $aba$.
-Note that $X$ must contain either $abxayx$ or $abxyax$ because $aba$ (and not $axa$ or $xyx$) is the first sandwich in $X$.
-Furthermore, we can assume that both $aba$ and $xyx$ in this subordering are both the first such subordering to appear in $X$.
-If $z=b$, then $X$ contains $abxaxba\cong abcacba\cong\mathcal{T}_{4}$.
-So assume $z=a$.
-If $y=b$, then $X$ must contain $abxabxab$ or $abxbaxab$, but then $bxb$ precedes $xbx$, so $xbx=xyx$ is not the first sandwich which blocks $z=a$.
+By the case, an $axa$ sandwich precedes the first $xyx$ sandwich. Since the first sandwich that blocks $z$ is $xyx$, we must have $z=a$. Note that, because $xyx$ blocks $z=a$, we also have $y\neq a$.
+The color $x$ must follow the $ab$ in the first sandwich $aba$, as otherwise an $axa$ sandwich would proceed the first $aba$ sandwich.
+Since $xyx$ is not the first sandwich in $X$ and $y\neq a$, either $X\supseteq abxayx$ or $X\supseteq abxyax$.
+We can assume that both $aba$ and $xyx$ in this subordering are the first such subordering to appear in $X$.
+If $y=b$, then the subordering is $abxabx$ or $abxbax$.
+But then $bxb$ precedes $xbx=xyx$, so $xbx=xyx$ is not the first sandwich which blocks $z=a$.
 Therefore, we can assume $y\neq b$.
-As such, $X$ contains either $abxayxab\cong abcadcab\cong\mathcal{T}_{8}$ or $abxyax$, so we assume $X$ contains $abxyax$.
-Recall that every color must follow $abxya$, so every color that is not $a$ is blocked by either $axa$ or $aya$.
+Therefore, the colors $a$ and $b$ are distinct from the colors $x$ and $y$, so either $X\supseteq abxayxab\cong abcadcab\cong\mathcal{T}_{8}$ or the subordering is $abxyax$.
+So, we can assume the subordering is $abxyax$.
+Recall that every color must follow the first $aba$ sandwich, so every color must follow $abxya$.
+This with the fact that $a$ is neither $x$ nor $y$ implies that every color that is not $a$ is blocked by either $axa$ or $aya$.
 Furthermore, $a$ is blocked by $xyx$.
-Therefore every color is blocked by one of the three sandwiches $axa$, $aya$, or $xyx$.
-Since none of these three sandwiches contains $b$, if we remove the color $b$ from $X$ we have a trivially unsortable sock ordering which contradicts the minimality of $X$.
-Case 3.2: The first $x$ appears after the first sandwich $aba$.
+Therefore, every color is blocked by one of the three sandwiches $axa$, $aya$, or $xyx$.
+Since none of these three sandwiches includes the color $b$, if we remove the color $b$ from $X$ we have a trivially unsortable sock ordering which contradicts the minimality of $X$.
+Case 3.3: The first $x$ appears after the first sandwich $aba$.
 So, we have $abaxyx$ in $X$ were $aba$ and $xyx$ in this subordering are the first such subordering to appear in $X$.
@@
 Therefore, the colors $x$ and $y$ are distinct from the colors $a$ and $b$.
-Hence, $X$ either contains $abaxyxba\cong abacdcba\cong\mathcal{T}_{9}$ or $abaxyxab\cong abacdcab\cong\mathcal{T}_{10}$.
+Hence, $X\supseteq abaxyxba\cong abacdcba\cong\mathcal{T}_{9}$ or $X\supseteq abaxyxab\cong abacdcab\cong\mathcal{T}_{10}$.
 ∎
-See 27
+See 10
 Proof.
-First note that if a strict subordering of $X$ is unsortable, then $X$ is unsortable.
-Therefore, to prove the lemma, we can assume that every strict subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$, since this will imply that $X$ is unsortable.
+First note that if a proper subordering of $X$ is unsortable, then $X$ is unsortable.
+Therefore, to prove the lemma, we can assume that every proper subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$, since this will imply that $X$ is unsortable.
 Assume for a contradiction that no subpattern of $X$ is in $\Gamma$.
@@
 Note that $a$ is the only sortable color in $abtatb$ and $abtabt$.
-Furthermore, $b$ is the only good sortable color in $abtbI(T)at_{n}$ (when $n=0$ the color $b$ is the only sortable color in $abtbI(T)at_{n}=abtbat$ and when $n\geq 1$, $b$ and $t$ are the only sortable colors in $abtbI(T)at_{n}$, but sorting $t=t_{0}$ leaves $(abbt_{1},I(t_{1},\dotsc,t_{n})at_{n})$ which is unsortable by Observation 20).
+We claim that $b$ is the only good sortable color in $abtbI(T)at_{n}$.
+To see this note that when $n=0$ the color $b$ is the only sortable color in $abtbI(T)at_{n}=abtbat$ and when $n\geq 1$, $b$ and $t$ are the only sortable colors in $abtbI(T)at_{n}$, but sorting $t=t_{0}$ leaves $(abbt_{1},I(t_{1},\dotsc,t_{n})at_{n})$ which is unsortable by Observation 2.
   • If $P\in\{atb,abt\}$, then let $z=a$ and $y=b$.
@@
 The first conjunct is given by the hypothesis of the lemma.
-For the second conjunct, recall that, by the above discussion, $z$ is the only good sortable color in $abtP$.
-Therefore, because $X-d$ is sortable by assumption, $z$ is not blocked by a sandwich consisting only of colors in $P$.
+For the second conjunct, recall that, by the above discussion, $z$ is the only good sortable color in $abtP\subseteq X-d$.
+Therefore, because $X-d$ is a proper subordering of $X$ and is therefore sortable, $z$ is not blocked by a sandwich consisting only of colors in $P$.
 ∎
@@
   1. (i)
-    $S^{\prime}$ contains one of $dabt$, $adbt$, $abdt$, or for some color $e\notin P$, $S^{\prime}$ contains $abtde$; or
+    one of $dabt$, $adbt$, $abdt$, is a subordering of $S^{\prime}$ or $abtde\subseteq S^{\prime}$ for some color $e\notin P$; or
   2. (ii)
-    $R^{\prime}$ contains either $atab$ or $tP$, or
+    $atab\subseteq R^{\prime}$ or $tP\subseteq R^{\prime}$; or
   3. (iii)
-    for some color $e\notin P$, $e\in S^{\prime}$ and $R^{\prime}$ contains $eP$.
+    for some color $e\notin P$, $e\in S^{\prime}$ and $eP\subseteq R^{\prime}$.
 Proof.
@@
 Case 1: A sandwich of the form $ded$ blocks $z$.
-In this case, we will show that (i) holds, so suppose $dabt$, $adbt$, and $abdt$ are not contained in $S^{\prime}$.
+In this case, we will show that (i) holds, so suppose that neither $dabt$, $adbt$, nor $abdt$ is a subordering of $S^{\prime}$.
 Then, because $d\notin R^{\prime}$ implies $de\subseteq S^{\prime}$, we have $abtde\subseteq S^{\prime}$.
@@
 Since $ded$ blocks $z$, we have $e\neq z$.
-We also have $e\neq y$, because, by Claim 33 a sandwich of the form $yty$ cannot block $z$ in $X=S^{\prime}dR^{\prime}$.
+We also have $e\neq y$, since otherwise we would have a $yty$ sandwich in $S^{\prime}$ which would block $z$ and this is forbidden by Claim 16.
 Furthermore, $e\neq t$, since we are assuming $abdt\subseteq abtdt$ is not in $S^{\prime}$.
@@
 Since $ede$ blocks $z$, we have $e\neq z$.
-Therefore, $e\notin\{a,b\}$.
+Therefore, $e\notin\{y,z\}=\{a,b\}$.
 Case 2.1: $P\in\{atb,abt,bat=bI(t_{0})at_{0}\}$.
@@
 Since $yt\subseteq abt\subseteq S$, $z$ cannot follow the $y$ in $P$.
-This, with the fact that $ez\subseteq R^{\prime}$, implies that we have $eP\subseteq R^{\prime}$ when $P\in\{abt,bI(t)at=bat\}$.
-Furthermore, if $P=atb$ and $eP\nsubseteq R^{\prime}$, then, regardless of $e$, we must have $atab\subseteq R^{\prime}$.
-First note that if $e=t$, then condition (ii) is satisfied, so assume $e\neq t$.
+This, with the fact that $ez\subseteq R^{\prime}$, implies that we have $eP\subseteq R^{\prime}$ when $P$ begins with $zy$.
+So, we have $eP\subseteq R^{\prime}$ when $P\in\{abt,bI(t)at=bat\}$.
+If $P=atb$ and $eP\nsubseteq R^{\prime}$, then, because $ea=ez\subseteq R^{\prime}$ and $z=a$ cannot follow $y=b$ in $P$, we either have $aetab\subseteq R^{\prime}$ or $ateab\subseteq R^{\prime}$.
+Note that, in both cases, we have $atab\subseteq R^{\prime}$.
+First note that if $e=t$, then the preceding argument implies that condition (ii) is satisfied, so assume $e\neq t$.
 Recall that we have already shown that $e\notin\{a,b\}$, so, since we are in Case 2.1, we have $e\notin P$.
@@
 For a contradiction, assume $P=atb$ and $eP\nsubseteq R^{\prime}$.
-So $z=a$ and $y=b$.
-Recall that we have previously shown that $ea=ez\subseteq R^{\prime}$ and that $z=a$ cannot follow the $y=b$ in $P$.
-Furthermore, because $eP\nsubseteq R^{\prime}$, we have $atab\subseteq R^{\prime}$.
-Therefore, we either have $aetab\subseteq R^{\prime}$ or $ateab\subseteq R^{\prime}$.
+Recall that we have previously shown that $e\in S^{\prime}$ and that, in this case, we either have $aetab\subseteq R^{\prime}$ or $ateab\subseteq R^{\prime}$.
 Since $e$ and $t$ are both in $S^{\prime}$, a sandwich consisting of the colors $e$ and $t$ blocks both $a$ and $b$ in $X=S^{\prime}dR^{\prime}$.
 Since $abt\subseteq S^{\prime}$, a sandwich of the form $ata$ blocks $e$ in $X=S^{\prime}dR^{\prime}$ and a sandwich of the form $aba$ blocks $t$ in $X=S^{\prime}dR^{\prime}$.
-Therefore, if we consider the pattern $X^{\prime}$ formed by removing all colors that are not $a$, $b$, $t$, or $e$ from $X$, we have that $X^{\prime}$ is a subordering of $X-d$ that is trivially unsortable.
-This contradicts the fact that every strict subordering of $X$ is sortable.
+Therefore, if we consider the subordering $X^{\prime}$ formed by removing all colors that are not $a$, $b$, $t$, or $e$ from $X$, we have that $X^{\prime}$ is a subordering of $X-d$ that is trivially unsortable.
+This contradicts the fact that every proper subordering of $X$ is sortable.
 Case 2.2: $n\geq 1$ and $P=bI(T)at_{n}=bt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$.
@@
 The color $b$ cannot follow $t_{1}t$ in $R^{\prime}$ because Claim 16 implies that $b$ is not blocked by a $tt_{1}t$ sandwich.
-Therefore, because $P=bt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$ and $eb\subseteq R^{\prime}$, if we do not have $eP\subseteq R^{\prime}$, then, regardless of $e$, we must have $bt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\subseteq R^{\prime}$.
-So, we have the following two cases.
-Case 2.2.1: $R^{\prime}$ contains $eP=ebt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$.
+Therefore, because $P=bt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$ and $eb\subseteq R^{\prime}$, if we do not have $eP\subseteq R^{\prime}$, then we must have $bt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\subseteq R^{\prime}$.
+But then,
+$X\supseteq abtdR\supseteq abtdbt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\supseteq btdt_{1}btt_{1}\cong dabtdat\cong I_{1,0},$
+a contradiction.
+Therefore, we can assume that $R^{\prime}\supseteq eP=ebt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$.
 Note that if $e=t$, then (ii) holds and if $e\notin P$, then (iii) holds.
@@
 $X\supseteq abtt_{i}dt_{i}bt_{1}tI(t_{1},\dotsc t_{n})at_{n}\supseteq btdt_{i}btt_{i}\cong dabtdat\cong I_{1,0},$
-a contradiction.
-Case 2.2.2: $R^{\prime}$ contains $bt_{1}btI(t_{1},\dotsc,t_{n})at_{n}$. In this case, both when $n=1$ and when $n\geq 2$, we have
-$X\supseteq abtdR\subseteq abtdbt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\supseteq btdt_{1}btt_{1}\cong dabtdat\cong I_{1,0},$
 a contradiction. ∎
 Suppose Claim 17(i) holds.
-If $dabt\subseteq S^{\prime}$, then we have $at\subseteq P$ or we have $n\geq 1$ and $I(T)at_{n}\subseteq P$.
-This means that $X$ contains $dabtdat\cong I_{1,0}$ or $dabtdI(T)at_{n}\cong I_{1,n}$.
-Otherwise, $X$ contains one of the patterns in the following table.
+Note that $at\subseteq P$ or $n\geq 1$ and $I(T)at_{n}\subseteq P$.
+Therefore, if $dabt\subseteq S^{\prime}$, then either $X\supseteq dabtdat\cong I_{1,0}$ or $X\supseteq dabtdI(T)at_{n}\cong I_{1,n}$.
+Otherwise, one of the orderings in the following table is a subordering of $X$.
 $adbtdatb\cong\mathcal{C}_{2,4}$ $adbtdabt\cong\mathcal{C}_{2,3}$ $adbtdbI(T)at_{n}\cong I_{2,n}$
@@
 $abtdedatb\cong\mathcal{C}_{2,7}$ $abtdedabt\cong\mathcal{C}_{2,6}$ $abtdedbI(T)at_{n}\cong I_{5,n}$
-If Claim 34(iii) holds, then by swapping the labels of $d$ and $e$, we can see that $X$ contains either $I_{1,n}$ or one of the patterns from the table above.
-Finally, if Claim 34(ii) holds, then $X$ must contain one of the four patterns in the following table.
+If Claim 17(iii) holds, then by swapping the labels of $d$ and $e$, we can see that $X$ either $I_{1,n}\preceq X$ or one of the orderings in the table above is a subpattern of $X$.
+Finally, if Claim 17(ii) holds, then one of the four orderings in the following table is a subordering of $X$.
 $abtdatab\cong\mathcal{C}_{2,1}$ $abtdtatb\cong\mathcal{C}_{2,2}$ $abtdtabt\supseteq abdtabt\cong I_{1,0}$ $abtdtbI(T)at_{n}\cong I_{4,n}$
-See 28
+See 11
 Proof.
-As in the proof of Lemma 27, to prove the lemma, we can assume that every strict subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$.
+As in the proof of Lemma 10, to prove the lemma, we can assume that every proper subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$.
 For a contradiction, assume that no subpattern of $X$ is in $\Gamma$.
@@
 Since $(at,R)$ is terminal, there exists a sandwich that blocks $t$ in $atR$.
-Therefore, there exists a color that is neither $a$ or $t$ in $R$.
+Therefore, there exists a color that is neither $a$ nor $t$ in $R$.
 Let $r$ be the first color that is neither $a$ nor $t$ in $R$.
@@
 Note that then $x$ is one of the first two colors to appear in $tR$.
-Lemma 24 implies that $x$ is unsortable in $(t,R)$, because $x$ is unsortable in $(at,R)$ and $ax\nsubseteq R$.
+Lemma 7 implies that $x$ is unsortable in $(t,R)$, because $x$ is unsortable in $(at,R)$ and $x\neq a$ and $ax\nsubseteq R$.
 Recall that by the hypothesis of the lemma, there exists $y$ that is sortable in $(t,R)$.
-Since $y$ is unsortable in $(at,R)$, Lemma 24 implies that $ay\subseteq R$.
-Then $y$ must follow the last occurrence of $x$ in $R$, because $ax\nsubseteq R$.
-However, with Lemma 22, the fact that $y$ is sortable in $(t,R)$ while $x$ is unsortable in $(t,R)$, and the fact that $x$ is one of the first two colors to appear in $tR$, implies that $y$ cannot follow the last occurrence of $x$ in $R$, a contradiction.
+Since $y$ is unsortable in $(at,R)$, Lemma 7 implies that either $y=a$ or $ay\subseteq R$.
+In either case, $y$ must follow the last occurrence of $x$ in $R$, because $ax\nsubseteq R$.
+However, with Lemma 5, the fact that $y$ is sortable in $(t,R)$ while $x$ is unsortable in $(t,R)$, and the fact that $x$ is one of the first two colors to appear in $tR$, imply that $y$ cannot follow the last occurrence of $x$ in $R$, a contradiction.
 ∎
@@
 Since, by Claim 18, we have $ar\subseteq R$, one of $arat$, $rart$, or $ratr$ appears in $R$.
-Since $atdratr=atdraI(r)tr\cong I_{1,0}$, we can assume the pattern $ratr$ does not appear $R$, so we are done in this case.
+Since $atdratr=atdraI(r)tr\cong I_{1,0}$, we can assume the subordering $ratr$ does not appear $R$, so we are done in this case.
 Now, suppose $rat\nsubseteq R$.
-Since $r$ is the first color in $R$ that is not $a$ or $t$, $rat\nsubseteq R$ implies that the sandwich that blocks $t$ in $atR$ cannot contain $a$.
-So there exist distinct colors $c,b$ that are not $a$ or $t$, such that a sandwich of the form $cbc$ blocks $t$ in $atR$.
-Since, by Claim 35, we have $at\subseteq R$, and $r$ is the first color that is not $a$ or $t$ to appear in $R$, and $rat\nsubseteq R$, we have either $r=c$ and $arbrt\subseteq R$ or $r\notin\{b,c\}$ and $arcbct\subseteq R$.
+Since $r$ is the first color in $R$ that is not $a$ or $t$, $rat\nsubseteq R$ implies that the first sandwich that blocks $t$ in $atR$ does not include the color $a$.
+So there exist distinct colors $c$ and $b$ that are both neither $a$ nor $t$, such that a sandwich of the form $cbc$ blocks $t$ in $atR$.
+Since $r$ is the first color that is not $a$ or $t$ to appear in $R$, we have either $r=c$ and $rbrt\subseteq R$ or $r\notin\{b,c\}$ and $rcbct\subseteq R$.
+By Claim 18, we have $at\subseteq R$.
+This, with the fact that $rat\nsubseteq R$, implies that either $arbrt\subseteq R$ or $arcbct\subseteq R$, where the distinct colors $b$ and $c$ are distinct from $a$, $r$, and $t$.
 ∎
 Claim 20.
-For distinct colors $a$, $r$, $t$, $b$, and $c$ and every $P\in\{arat,rart,arbrt,arcbct\}$ the color $a$ is the only sortable color in $atP$.
+For distinct colors $a$, $r$, $t$, $b$, and $c$ and every $P\in\{arat,rart,arbrt,arcbct\}$ the color $a$ is the only sortable color in $atP$ and $a$ is a good sortable color in $atP$.
 Therefore, no sandwich consisting only of the colors that appear in $atP$ blocks $a$ in $X$.
-In particular, this implies that $rt\nsubseteq S$ and $atPa\nsubseteq X$.
 Proof.
@@
 ∎
-Note that Claims 36 and 37 establish the first statement of the conclusion.
+Note that Claims 19 and 20 establish the first statement of the conclusion.
 Claim 21.
@@
   1. (i)
-    $S^{\prime}$ contains one of $dat$, $adt$, or, for some color $e\notin P$, $S^{\prime}$ contains $atde$; or
+    $dat$ or $adt$ is a subordering of $S^{\prime}$ or $atde\subseteq S^{\prime}$ for some color $e\notin P$; or
   2. (ii)
-    $R^{\prime}$ contains one of $taxat$, $atxat$, $txaxt$, or $taxyxt$ where $a$, $t$, $x$, and $y$ are distinct colors; or
+    $taxat$, $atxat$, $txaxt$, or $taxyxt$ is a subordering of $R^{\prime}$ where $a$, $t$, $x$, and $y$ are distinct colors; or
   3. (iii)
-    for some color $e\notin P$, $R^{\prime}$ contains $eP$ and $S^{\prime}$ contains one of $eat$, $aet$, or $ate$
+    $eP\subseteq R^{\prime}$ for some color $e\notin P$ and $eat$, $aet$, or $ate$ is a subordering of $S^{\prime}$.
 Proof.
@@
 Case 1: A sandwich of the form $ded$ blocks $a$ in $X$.
-Note that since $d\notin R^{\prime}$, we have $d\in S^{\prime}$.
-So, if $S^{\prime}$ does not contain $dat$ or $adt$, then $S^{\prime}$ must contain $atde$ and $e\neq t$.
+Note that since $d\notin R^{\prime}$, we have $de\subseteq S^{\prime}$.
+So, if neither $dat\subseteq S^{\prime}$ nor $adt\subseteq S^{\prime}$, then $atde\subseteq S^{\prime}$ and $e\neq t$.
 So to show that $X$ satisfies (i), we only need to show that $e\notin P$.
 First note that $e\neq a$, since $ded$ blocks $a$.
-If $e=r$, then, since we always have $art\subseteq P$, the sequence $S^{\prime}dP\subseteq X$ contains $atdrdart\cong abdtdatb\cong C_{2,5}$, a contradiction.
-We get a similar contradiction if $P=arbrt$ or $P=arcbct$ and $e=b$, since in either case $P$ contains $abt$, so $S^{\prime}dP$ contains $atdbdabt\cong abdtdatb\cong C_{2,5}$.
-If $P=arcbct$ and $e=c$, then $P$ contains $act$, so $S^{\prime}dP$ contains $atdcdact\cong abdtdatb\cong C_{2,5}$.
-This completes the proof of this case.
+If $e=r$, then, since Claim 19 implies that $art\subseteq P$, we have $X\supseteq S^{\prime}dP\supseteq atdrdart\cong abdtdatb\cong C_{2,5}$, a contradiction.
+We get a similar contradiction if $P=arbrt$ or $P=arcbct$ and $e=b$, since in either case $abt\subseteq P$, so $S^{\prime}dP\supseteq atdbdabt\cong abdtdatb\cong C_{2,5}$, a contradiction.
+If $P=arcbct$ and $e=c$, then $act\subseteq P$, so $S^{\prime}dP\supseteq atdcdact\cong abdtdatb\cong C_{2,5}$, a contradiction.
+With Claim 19, this completes the proof of this case.
 Case 2: A sandwich of the form $ede$ blocks $a$ in $X$.
@@
 If $P=arbrt$, then one option is $tarbrt=eP\subseteq R^{\prime}$, and (ii) holds in this case.
-By Claim 37, there is no $rbr$ sandwich that blocks $a$, so the only other possibilities when $P=arbrt$ are for $R^{\prime}$ to contain $atrabrt$ or $atrbart$, and these both contain $trart$, so (ii) holds.
+By Claim 20, there is no $rbr$ sandwich that blocks $a$, so the only other possibilities when $P=arbrt$ are $atrabrt\subseteq R^{\prime}$ or $atrbart\subseteq R^{\prime}$, and, in either case, we have $trart\subseteq R^{\prime}$, so (ii) holds.
 Finally, assume $P=arcbct$.
-If $tarcbct=eP\subseteq R^{\prime}$, then $R^{\prime}$ contains $tacbct$, so (ii) holds.
-Otherwise, because there is no $cbc$ sandwich that blocks $a$, the pattern $R^{\prime}$ must contain one of $atracbct$, $atrcabct$, or $atrcbact$.
-These contain $tacbct$, $tcact$, and $tcact$, respectively, so (ii) holds.
+If $tarcbct=eP\subseteq R^{\prime}$, then $tacbct\subseteq R^{\prime}$, so (ii) holds.
+Otherwise, because there is no $cbc$ sandwich that blocks $a$, at least one of $atracbct\supseteq tacbct$, $atrcabct\supseteq tcact$, or $atrcbact\supseteq tcact$ is a subordering of $R^{\prime}$, so (ii) holds.
 Case 2.2: $e=r$.
 In this case, $r\in S^{\prime}$ and $ra=ea\subseteq R^{\prime}$.
-Note that, by Claim 37, there is no sandwich of the form $rtr$ that blocks $a$ in $X$.
+Recall that, by Claim 20, there is no sandwich of the form $rtr$ that blocks $a$ in $X$.
 Therefore, $r$ must follow $at$ in $S^{\prime}$, so $atr\subseteq S^{\prime}$.
-Since $ar\subseteq P$ and $t\in P$, either $art$, $atr$, or $tar=tI(r)ar$ is in $P\subseteq R$.
-Therefore, Lemma 27 implies that a subpattern of $X$ is in $\Gamma$, a contradiction.
+Because $art\subseteq P$ and $a$ is blocked by a $d$-sandwich, Lemma 10 (with $r$ playing the role of $b$) implies that a subpattern of $X$ is in $\Gamma$, a contradiction.
 Case 2.3: $e\neq r$ and $e\neq t$.
@@
 Recall that Claim 20 implies that every color in $atP$ is unsortable except $a$.
-If $e$ is blocked by a sandwich consisting of colors only in $P$, then the pattern formed by removing all colors except those in $P$ and $e$ from $X-d$ is trivially unsortable.
-This contradicts the assumption that all strict subpatterns of $X$ are sortable.
-Therefore, we can assume that $e$ is not blocked by a sandwich consisting of colors only in $P$ in $X$.
+If $e$ is blocked by a sandwich consisting of colors only in $P$, then the subordering formed by removing all colors except those in $P$ and $e$ from $X-d$ is trivially unsortable.
+This contradicts the assumption that all proper suborderings of $X$ are sortable.
 Now assume that $e\notin R$.
@@
 By the definition of $R^{\prime}$, the fact that $e\in R^{\prime}$ implies that the last $e$ follows the last $d$ in $X$.
-Since $e\notin R$ and the state after $d$ is sorted is $(S,R)$, this means that at some sorting step before the color $d$ was sorted, $de$ appeared on the stack.
-Therefore, $e$ was sorted before $d$.
+Since $e\notin R$ and the state after $d$ is sorted is $(S,R)$, the color $e$ was sorted before the color $d$.
 Since $d$ was sorted before any color in $P$, the color $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
-So, in all cases, $e\notin P$ and $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
-In particular, the pattern $ata$ does not block $e$ in $X$.
+So, in both cases, $e\notin P$ and $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
+In particular, the subordering $ata$ does not block $e$ in $X$.
 This means that $e$ does not follow $a$ in $R^{\prime}$.
@@
 This, with the fact that $e$ does does not follow $a$ in $R^{\prime}$ and $e\in R^{\prime}$, implies that $P=rart$ and $reart\subseteq R^{\prime}$.
-Therefore, $X=S^{\prime}dR^{\prime}$ contains either $eatdreart$, $aetdreart$, or $atedreart$.
-These contain $eatrear\cong dabtdat\cong I_{1,0}$, $aetreart\cong adbtdatb\cong\mathcal{C}_{2,4}$, and $atereart\cong abdtdatb\cong\mathcal{C}_{2,5}$, respectively, which is a contradiction.
+Therefore, at least one of $eatdreart$, $aetdreart$, or $atedreart$ is a subordering of $X=S^{\prime}dR^{\prime}$.
+This is a contradiction, because $eatdreart\supseteq eatrear\cong dabtdat\cong I_{1,0}$, $aetdreart\supseteq aetreart\cong adbtdatb\cong\mathcal{C}_{2,4}$, and $atedreart\supseteq atereart\cong abdtdatb\cong\mathcal{C}_{2,5}$.
 ∎
-When Claim 38(i) holds, then $X$ must contain one of the patterns in the following table.
+When Claim 21(i) holds, one of the orderings in the following table is a subordering of $X$.
 $datdarat\cong\mathcal{C}_{1,1}$ $datdrart\cong\mathcal{C}_{1,7}$ $datdarbrt\cong\mathcal{C}_{1,12}$ $datdarcbct\succeq\mathcal{C}_{1,12}$
@@
 $atdedarat\cong\mathcal{C}_{1,5}$ $atdedrart\cong\mathcal{C}_{1,9}$ $atdedarbrt\cong\mathcal{C}_{1,13}$ $atdedarcbct\succeq\mathcal{C}_{1,13}$
-If Claim 38(iii) holds, then by swapping $d$ and $e$, we can see that $X$ must contain one of these same patterns.
-Finally, if Claim 38(ii) holds, then $X$ must contain one of the patterns in the following table.
+If Claim 21(iii) holds, then by swapping $d$ and $e$, we can see that one of orderings in the preceding table is a subpattern of $X$.
+Finally, if Claim 21(ii) holds, then one of the orderings in the following table is a subpattern of $X$.
 $atdtarat\cong\mathcal{C}_{1,2}$ $atdatrat\cong\mathcal{C}_{1,3}$ $atdtrart\cong\mathcal{C}_{1,6}$ $atdtarbrt\cong\mathcal{C}_{1,10}$
+3. Further Directions
+While the process of foot sorting lends itself naturally to sorting with a stack, one could consider analogs of the sock sorting problem which utilize other stack configurations or other data structures entirely.
+3.1. Sorting with Multiple Stacks
+Defant and Kravitz introduced a notion of foot-sorting with multiple feet in [1], though with more stacks in series, the sorting process becomes increasingly computationally intensive.
+It would be interesting to investigate perhaps the simplest extension of our work, the basis of sock orderings which are foot-sortable with two stacks in series.
+Defant and Kravitz proved in [1] that for any number $n\geq 1$, one may construct a sock ordering which is unsortable with $n$ stacks.
+Using this proof, one can check that the sock ordering of the form $(abcde)^{1440}$ (that is, the block $abcde$, where $a,b,c,d$ and $e$ are distinct sock colors, repeated 1440 times) is unsortable with two stacks in series, but it might be challenging to construct a single sock ordering that is minimally unsortable in this setting.
+3.2. Deque Sorting
+Because a deque, or double-ended queue, allows for more access to the socks for addition and removal, one might think about deque-sorting for socks independently of a foot; a sock could be placed either over an outermost sock or within an innermost sock, and the same is possible for sock removal.111 Using a deque in the context of permutation sorting has been consider previously (e.g.
+see [3], [2], and [4]).
+This modifies the original operations of foot sorting, using the language of the deque, to be
+  1. ($\text{stack}_{\uparrow}$)
+    Take the leftmost sock on the right and place it on top of the deque (potentially over any socks that are already in the deque).
+  2. ($\text{stack}_{\downarrow}$)
+    Take the leftmost sock on the right and place it on the bottom of the deque (potentially under any socks that are already in the deque).
+  3. ($\text{unstack}_{\uparrow}$)
+    Remove the topmost sock from the deque and place it so that it becomes the rightmost sock on the left.
+  4. ($\text{unstack}_{\downarrow}$)
+    Remove the bottommost sock from the deque and place it so that it becomes the rightmost sock on the left.
+While the addition of possible moves in this scenario makes for more difficult computation, some similar ideas hold as in the stack setting: in particular, it is now necessary to avoid the pattern $abab$ from appearing in the deque, just as it is necessary to avoid $aba$ appearing in the stack.
+It would be interesting to enumerate the basis of deque-sortable sock orderings, just as we have enumerated the basis of foot-sortable sock orderings.
+It is straightforward to show that the basis of deque-sortable sock orderings is different from $\Gamma$, as well as to show that this basis is infinite.
+Very closely mirroring the proof of Observation 2 under the relaxed restrictions of the deque, one can generate the class of sock orderings of the form $dabtI_{1,n}$ for any $n\geq 0$ and verify that such sock orderings are minimally unsortable with the deque.
+However, with the increased computational difficulty in the deque setting, we do not know whether other minimal deque-unsortable sock orderings may be generated from $\Gamma$ in a similar fashion.

### v3

@@
 Consider the following problem, which was introduced by Defant and Kravitz [1].
-Starting with an arbitrary line of colored socks to the right of a foot, one aims to line the socks up by color to the left of the foot, but, at any point in time, only one of the following two actions can be performed.
+Starting with an arbitrary line of colored socks to the right of a foot, one aims to move the socks to the left of the foot so that the ordering is sorted by color; that is, the output ordering consists of blocks of socks so that all socks of the same color appear consecutively.
+At any point in time, however, only one of the following two actions can be performed.
   1. (stack)
@@
     Remove the outermost sock from the foot and place it so that it becomes the rightmost sock on the left.
-The initial ordering of socks on the right is a sock ordering and a sock ordering is foot-sortable (or simply sortable) if, using only the two specified operations, one can produce a sorted ordering of the socks on the left.
+The initial ordering of socks on the right is a sock ordering and a sock ordering is foot-sortable (or simply sortable) if using only the two specified operations, one can produce a sorted ordering of the socks on the left.
 Otherwise, it is unsortable.
 Throughout, we represent sock orderings as strings in the obvious way; left-to-right in the string corresponds to left-to-right in the sock ordering.
-For example, if $r$, $g$, and $b$ refer to socks of color red, green, and blue, respectively, the sock ordering $rgbrgb$ is sortable (to $rrbbgg$ on the left) while the sock ordering $rgbgrgb$ is unsortable.
-Let $X$ and $Y$ be sock orderings.
-We say that $Y$ is a subordering of $X$ and write $Y\subseteq X$ (or $X\supseteq Y$) if deleting zero or more socks from $X$ yields $Y$ exactly.
-The subordering $Y\subseteq X$ is proper if $X\neq Y$.
-An unsortable sock ordering is minimally unsortable if every proper subordering is sortable.
-We write $a\in X$ if the color $a$ appears in $X$ and we let $X-a$ be the subordering of $X$ obtained by deleting all socks of color $a$ from $X$.
-We use $\emptyset$ to denote the empty ordering and $a$ to denote the ordering consisting of a single sock of color $a$.
-We let $XY$ be the concatenation of the orderings $X$ and $Y$.
-We say $X$ and $Y$ are equivalent and write $X\cong Y$ if $Y$ can be obtained from $X$ via a bijective mapping of the colors of $X$ to the colors of $Y$.
-We call $Y$ a subpattern of $X$ and write $Y\preceq X$ (or $X\succeq Y$) if there exists a sock ordering $Z$ such that $Z\subseteq X$ and $Z\cong Y$.
-Otherwise, we say that $X$ avoids $Y$.
-Furthermore, $X$ avoids a collection of sock orderings if it avoids every ordering in the collection.
-Modulo the relation $\cong$ and with respect to the partial ordering $\preceq$, Defant and Kravitz called the set of minimal elements among the collection of unsortable sock orderings the basis of foot-sortable sock orderings [1].
-One may equivalently view the basis of foot-sortable sock orderings as a collection of pairwise inequivalent minimally unsortable sock orderings that no unsortable sock ordering avoids.
-In this paper, we give an explicit description of the basis of foot-sortable sock orderings (Section 2.2) and a proof that no unsortable sock ordering avoids this basis (Theorem 3).
+For example, if $r$, $g$, and $b$ refer to socks of color red, green, and blue, respectively, the sock ordering rgbrgbrgbrgb is sortable (to rrbbggrrbbgg on the left) while the sock ordering rgbgrgbrgbgrgb is unsortable.
+We will use the following two definitions to describe our result.
+Two sock orderings are equivalent if one can be obtained from the other by a bijective mapping of the colors.
+So, for example, we think of rgbgrgbrgbgrgb as equivalent to gbrbgbrgbrbgbr.
+We call a sock ordering minimally unsortable if it is unsortable and removing any sock yields a sortable sock ordering.
+For example, rgbgrgbrgbgrgb is minimally unsortable, because it is unsortable and the sock orderings gbgrgbgbgrgb, rbgrgbrbgrgb, rggrgbrggrgb, rgbrgbrgbrgb, rgbggbrgbggb, rgbgrbrgbgrb, and rgbgrgrgbgrg are each sortable (to gggrbbgggrbb, rrggbbrrggbb, rrgggbrrgggb, rrbbggrrbbgg, rgggbbrgggbb, ggbbrrggbbrr, and bgggrrbgggrr, respectively).
+Modulo equivalence, Defant and Kravitz called the unique set of minimally unsortable sock orderings the basis of foot-sortable sock orderings [1].
+In this paper, we explicitly described this set.
+This description appears in Section 2.3, and the rest of the paper is devoted to proving that this set is indeed the basis of foot-sortable sock orderings.
+(For reference, rgbgrbgrgbgrbg is equivalent to the ordering $\mathcal{T}_{6}$ that is defined in Section 2.3.)
 1.1. Prior Work
@@
 Defant and Kravitz also considered several variants of this problem, including employing multiple stacks (in particular two stacks or two feet) and restricting the number of socks of a given color.
-Of particular interest with respect to socks, is the case when there are at most two socks of every color.
+Of particular interest with respect to socks is the case when there are at most two socks of every color.
 Independent of our work, Yu explicitly described the basis for foot-sortable orderings in which each color appears at most twice [6].
@@
 Also independent of our work, Yu devised a fast deterministic algorithm that decides if a sock ordering of length $N$ is sortable in time $O(N\log N)$.
-Provided that the ordering is sortable, this algorithm can also procedure a valid sorting.
+Provided that the ordering is sortable, this algorithm can also produce a valid sorting.
 This problem was also investigated by Xia, who considered a deterministic algorithm that forces the stack to remain sorted [5].
-2. Results
-2.1. Preliminaries
+2. Preliminary discussion and a description of the basis.
+2.1. Additional Definitions and Notation
+Let $X$ and $Y$ be sock orderings.
+We say that $Y$ is a subordering of $X$ and write $Y\subseteq X$ (or $X\supseteq Y$) if deleting zero or more socks from $X$ yields $Y$ exactly.
+The subordering $Y\subseteq X$ is proper if $X\neq Y$.
+So, an unsortable sock ordering is minimally unsortable if every proper subordering is sortable.
+We write $a\in X$ if the color $a$ appears in $X$ and we let XaX-a be the subordering of $X$ obtained by deleting all socks of color $a$ from $X$.
+We use $\emptyset$ to denote the empty ordering and $a$ to denote the ordering consisting of a single sock of color $a$.
+We let $XY$ be the concatenation of the orderings $X$ and $Y$ and we write $X\cong Y$ when $X$ and $Y$ are equivalent.
+We call $Y$ a subpattern of $X$ and write $Y\preceq X$ (or $X\succeq Y$) if there exists a sock ordering $Z$ such that $Z\subseteq X$ and $Z\cong Y$.
+Otherwise, we say that $X$ avoids $Y$.
+Furthermore, $X$ avoids a collection of sock orderings if it avoids every ordering in the collection.
 Throughout, we will refer to the sorting process using primarily stack-centered language.
@@
 As such, we implicitly assume throughout that the stack $S$ is sorted.
+We call a state $(S^{\prime},R^{\prime})$ a substate of $(S,R)$ if $S^{\prime}\subseteq S$ and $R^{\prime}\subseteq R$, and it is proper if at least one of the two containment relations is strict.
 Instead of dealing with the two operations stack and unstack, we work almost exclusively at a slightly higher level.
-At each point, we are mainly concerned with the next sock color that will be moved completely to the left and we say that a color is sortable if it is possible to do this while keeping the stack is sorted.
+At each point, we are mainly concerned with the next sock color that will be moved completely to the left, and we say that a color is sortable if it is possible to do this while keeping the stack sorted.
+Of great importance in the context of moving a color to the left is the concept of a sandwich: a configuration of (possibly non-consecutive) socks taking the form bcbbcb, where $b$ and $c$ are distinct colors.
+We refer to a sandwich of the form bcbbcb as either a $b$-sandwich or a $c$-sandwich.
 More explicitly, given a state $(S,R)$, a color $a\in SR$ is sortable if the following two conditions hold.
@@
     If $a$ appears in $R$, then $bcba\not\subseteq SR$ for any colors $b,c\in SR$ such that $a$, $b$, and $c$ are distinct.
-If a color $a\in R$ does not meet the second condition, we say that the color $a$ is blocked by a sandwich.
-We refer to a sandwich of the form $bcb$ as either a $b$-sandwich or a $c$-sandwich.
+If a color $a\in R$ does not meet the second condition for colors $b,c$ in $SR$, we say that the color $a$ is blocked by the sandwich bcbbcb.
 Given a state $(S,R)$, a color $a\in SR$ is a good sortable color if $a$ is sortable in $(S,R)$ and what remains after sorting $a$ in $(S,R)$ is foot-sortable.
@@
 For a sock ordering $X$, we let $\theta(X):=\theta(\emptyset,X)$.
-2.2. The basis of foot-sortable sock orderings
-We use $\Gamma$ to denote the basis of foot-sortable sock orderings. As mentioned previously, the set $\Gamma$ is infinite and we use the following definition to describe the infinite classes.
-Definition 1.
+2.2. Interlaced sequences
+As mentioned previously, the basis of foot-sortable sock orderings is infinite. We will use the following definition to help describe the infinite classes.
 Let $T:=(t_{0},\dotsc,t_{n})$ be a sequence of $n+1$ colors. Define $I(t_{0}):=I(T)$ to be the empty ordering when $n=0$, and when $n\geq 1$, let
 $I(T):=I(t_{0},\dotsc,t_{n}):=I(t_{0},\dotsc,t_{n-1})t_{n}t_{n-1}=t_{1}t_{0}t_{2}t_{1}\dotsc t_{n}t_{n-1}.$
-Observation 2.
+Observation 1.
 Let $n\geq 0$ and $a,b,t_{0},\dotsc,t_{n}$ be distinct colors.
 For every $0\leq i\leq n$, if $abt_{i}\subseteq S$ and $I(t_{i},\dotsc,t_{n})at_{n}\subseteq R$, then the state $(S,R)$ is unsortable.
-Furthermore, if $S^{\prime}\subseteq abt_{i}\subseteq S$ and $R^{\prime}\subseteq I(t_{i},\dotsc,t_{n})at_{n}\subseteq R$ are such that $S^{\prime}R^{\prime}\neq abt_{i}I(t_{i},\dotsc,t_{n})at_{n}$, then $(S^{\prime},R^{\prime})$ is sortable.
+Furthermore, for any $0\leq i\leq n$, if $(S^{\prime},R^{\prime})$ is a proper substate of $(abt_{i},I(t_{i},\dotsc,t_{n})at_{n})$, then $(S^{\prime},R^{\prime})$ is sortable.
 Proof.
-The proof is by induction on $n-i$. For the base case, we have $i=n$, so
+The proof is by induction on nin-i. For the base case, we have $i=n$, so
 $I(t_{i},\dotsc,t_{n})at_{n}=at_{n}$
@@
 leads to $(abt_{i+1},I(t_{i+1},\dotsc,t_{n})at_{n})$.
-Note that when $n-i=1$ the only sortable color is $t_{i}$, so the first statement follows by induction.
+Note that when ni=1n-i=1 the only sortable color is $t_{i}$, so the first statement follows by induction.
 When $n-i\geq 2$, the only sortable colors are $t_{i}$ and $t_{i+1}$ and sorting $t_{i+1}$ in
@@
 ∎
-The following tables describe the basis $\Gamma$. The first table consists of trivially unsortable orderings. It is not difficult to verify that the orderings are minimally unsortable.
+2.3. The basis of foot-sortable sock orderings
+The tables below describe the basis of foot-sortable sock orderings. We use $\Gamma$ to denote this set, so our main theorem is the following.
+Theorem 2.
+If $X$ is an unsortable sock ordering, then a subpattern of $X$ is in $\Gamma$.
+With Theorem 2, to show that $\Gamma$ is the basis of foot-sortable sock orderings we also need to show that every ordering in $\Gamma$ is minimally unsortable.
+The following table consists of the trivially unsortable orderings in $\Gamma$.
 $\mathcal{T}_{1}\cong abacaba$ $\mathcal{T}_{2}\cong abacbab$ $\mathcal{T}_{3}\cong abcabca$ $\mathcal{T}_{4}\cong abcacba$
@@
 $\mathcal{T}_{8}\cong abcadcab$ $\mathcal{T}_{9}\cong abacdcba$ $\mathcal{T}_{10}\cong abacdcab$
-Note that all of the sock orderings in the next table have either $atdarat$, $atdrart$, or $atdarbrt$ as a subordering.
-Furthermore, sorting the color $d$ (preceded or followed by sorting the color $e$ when it appears) for each of these three possible suborderings yields $(at,arat)$, $(at,rart)$, and $(at,arbrt)$, respectively, and each of these three states are terminal.
-Further note that in the orderings below, $d$ and $e$ are the only sortable colors, which implies that the orderings are unsortable.
-It is then not difficult to verify that the orderings are minimally unsortable.
+It is not hard to verify that for each ordering listed above, every color is blocked by sandwich, so the orderings are indeed trivially unsortable.
+It is also not difficult to verify that whenever a sock is removed, a color becomes sortable.
+This is enough to show that $\mathcal{T}_{1},\dotsc,\mathcal{T}_{7}$ are minimally unsortable, because they each contain only three distinct colors, which means that after a color is sorted in one of the suborderings only at most two colors remain, but three distinct colors are necessary for a color to be blocked by a sandwich.
+Similarly, for $\mathcal{T}_{8}$, $\mathcal{T}_{9}$, or $\mathcal{T}_{10}$, it can be checked that removing one sock produces an ordering in which two colors can be successively sorted.
+After this, only two distinct colors then remain, so the suborderings are sortable.
+All of the sock orderings in the next table have either atdaratatdarat, atdrartatdrart, or atdarbrtatdarbrt as a subordering.
+Each of these three suborderings is sortable if one first sorts the color $a$.
+However, for these three suborderings, sorting the color $d$ yields $(at,arat)$, $(at,rart)$, and $(at,arbrt)$, respectively, and each of these three states are terminal.
+So, to construct the unsortable sock orderings in this table, either a new sock is inserted so that a $d$-sandwich blocks $a$, or $ed$ is placed immediately after $d$, After this addition, the only sortable colors are $d$ and if $ed$ was added, $e$.
+This implies that the orderings are unsortable.
 $\mathcal{C}_{1,1}\cong datdarat$ $\mathcal{C}_{1,6}\cong atdtrart$ $\mathcal{C}_{1,10}\cong atdtarbrt$
@@
 $\mathcal{C}_{1,5}\cong atdedarat$
-The next table consists of the sock orderings that have either $abtdabt$ or $abtdatb$ as a subordering.
-In a similar fashion to the other classes, sorting $d$ (and possibly $e$) first in these orderings yields some $(S,R)$ where $abt\subseteq S$ and $at\subseteq R)$.
-Since $(abt,at)$ is unsortable and $d$ and $e$ are the only sortable colors in the following orderings, we have that orderings in this class are unsortable, and it is not difficult to verify that these are minimally unsortable.
+To see that these orderings are minimally unsortable, first consider the suborderings of $\mathcal{C}_{1,3}\cong atdatrat$ formed by removing a single sock:
+$tdatrat,adatrat,atatrat,atdtrat,atdarat,atdatat,atdatrt,\text{ and }atdatra.$
+For adatratadatrat, atatratatatrat, and atdaratatdarat, we have removed a sock from the $d$-sandwich that blocks $a$, so we can fully sort each of these three subordering by first sorting $a$, and then sorting $r$.
+For the remaining suborderings (tdatrattdatrat, atdtratatdtrat, atdatatatdatat, atdatrtatdatrt, and atdatraatdatra), we can first sort $d$ and then one additional color ($a$, $r$, $t$, $t$, and $t$, respectively), leaving at most two colors, so the suborderings are sortable.
+For every other subordering formed by removing one sock from an ordering in the table, one can verify that one of the following two possibilities hold:
+  • One of the socks in the $d$-sandwich that blocks $a$ has been removed and the subordering can then be fully sorted when one first sorts the color $a$, or
+  • sorting $d$ (and then $e$ if it appears) yields either $(S,R)$ or $(S,tR)$ where $(S,R)$ is a proper substate of one of the states $(at,arat)$, $(at,rart)$, or $(at,arbrt)$, thereby implying that the subordering is sortable.
+The next table consists of the sock orderings that have either abtdabtabtdabt or abtdatbabtdatb as a subordering.
 $\mathcal{C}_{2,1}\cong abtdatab$ $\mathcal{C}_{2,4}\cong adbtdatb$ $\mathcal{C}_{2,6}\cong abtdedabt$
@@
 $\mathcal{C}_{2,3}\cong adbtdabt$
-All that remains are the five infinite classes of the basis.
+Sorting $d$ (and possibly $e$) first in these orderings yields $(abt,R)$ where $at\subseteq R$.
+Since $(abt,at)$ is terminal and $d$ and $e$ are the only sortable colors in these orderings, the orderings in this table are unsortable.
+To address minimality, note that each of the following holds for every orderings in the table:
+  • If the final $t$ is removed, then $d$ is a good sortable color.
+  • If the first $a$ or $b$ is removed, then $d$ is a good sortable color.
+  • If the final $a$ or $b$ is removed, then the color of the removed sock becomes a good sortable color.
+  • If $d$ or $e$ is removed, then $a$ is a good sortable color.
+  • If there is only one sandwich that blocks $a$, then removing a sock from that sandwich makes $a$ a good sortable color.
+This covers all possibilities except removing the middle $a$ when there are three $a$’s or the first $t$ when it is not part of the only sandwich that blocks $a$.
+This leaves the following suborderings to consider:
+$abtdtab\subseteq\mathcal{C}_{2,1},\quad adbdabt\subseteq\mathcal{C}_{2,3},\quad adbdatb\subseteq\mathcal{C}_{2,4},\quad abdedabt\subseteq\mathcal{C}_{2,6},\quad\text{and}\quad abdedatb\subseteq\mathcal{C}_{2,7},$
+and $d$ is a good sortable color in each of these three suborderings.
 For every $n\geq 0$, let $a$, $b$, $d$, $e$, and $t_{0},\dotsc,t_{n}$ be distinct colors.
-Let $T:=(t_{0},\dotsc,t_{n})$ and define $t:=t_{0}$.
-Let $E:=I(T)at_{n}$ (so $E=at_{0}=at$ when $n=0$) and define
-$I_{1,n}:=dabtdE$ $I_{2,n}:=adbtdbE$ $I_{3,n}:=abdtdbE$ $I_{4,n}:=abtdtbE$ $I_{5,n}:=abtdedbE$
-With Observation 2 it is not hard to verify that these orderings are also minimally unsortable.
-As we have now specified a collection $\Gamma$ of minimally unsortable sock orderings, to prove that $\Gamma$ is the basis of foot-sortable sock orderings, we only need the following theorem.
-Theorem 3.
-If $X$ is an unsortable sock ordering, then a subpattern of $X$ is in $\Gamma$.
-2.3. Algorithm
-Our proof of Theorem 3 is inherently algorithmic, and we feel it is instructive to explicitly describe a deterministic algorithm that, given an arbitrary sock order, either sorts the ordering using only the operations stack and unstack, or indicates that the given sock ordering is unsortable.
-Our proof can be used to extend this algorithm so that, when the given sock ordering is unsortable, it also identifies a subpattern which is an element of the basis, thereby justifying the indication that the sock ordering is unsortable.
-For simplicity, we do not explicitly describe this extension.
-We made no attempt to optimize the running time of this algorithm.
+Let $T_{n}:=(t_{0},\dotsc,t_{n})$, $E_{n}:=I(T_{n})at_{n}$ and define $t:=t_{0}$.
+(Note that $E_{0}=at_{0}=at$).
+We define the following five orderings which each contain $abtdE_{n}$.
+$I_{1,n}:=dabtdE_{n}$ $I_{2,n}:=adbtdbE_{n}$ $I_{3,n}:=abdtdbE_{n}$ $I_{4,n}:=abtdtbE_{n}$ $I_{5,n}:=abtdedbE_{n}$
+By Observation 1, the state $(S,R)$ is unsortable if $abt\subseteq S$ and $E_{n}\subseteq R$.
+Therefore, none of the orderings described below can be fully sorted by first sorting $d$ (or $d$ and $e$ if $e$ is present).
+Since this is the only option for $I_{2,n}$, $I_{3,n}$, $I_{4,n}$, and $I_{5,n}$, those orderings are unsortable.
+In $I_{1,n}$, both the colors $b$ and $d$ are sortable, but sorting $b$ leads to $(da,tdE_{n})$ which is terminal, so $I_{1,n}$ is also unsortable.
+Let $(S,R)$ be a proper substate of $(abt,E_{n})$.
+Recall that Observation 1 implies that $(S,R)$ is sortable.
+We also claim that $(S,bR)$ is also sortable.
+To see this, first note that by adding socks we can assume that we either have $S=abt$ or $R=E_{n}$.
+Therefore, $t_{0}=t\in SR$.
+If $n\geq 1$, then $t$ is a good sortable color in $(S,bR)$, because $t=t_{0}$ is sortable in $(abt,bE_{n})$ and sorting $t$ in the state $(S,bR)$ yields a proper substate of $(abbt_{1},I(t_{1},t_{2},\dotsc,t_{n})at_{n})$ which is sortable by Observation 1.
+If $n=0$, then $(S,R)$ is a proper substate of $(abt,at)$, and it is not hard to then verify that $(S,bR)$ is then sortable.
+The preceding argument implies that both $(S,R)$ and $(S,bR)$ are sortable when $(S,R)$ is a proper substate of $(abt,E_{n})$.
+This immediately implies that if a sock in the initial abtabt or the terminal $E_{n}$ is removed from $I_{1,n}$, $I_{2,n}$, $I_{3,n}$, and $I_{5,n}$, then the subordering is sortable.
+For $I_{4,n}$, the same statement also holds.
+To see this, first note that we can sort $d$ and then $b$ in $abdtbE_{n}\subseteq I_{4,n}$ to reach the state $(at,E_{n})$.
+In all other cases when we delete a sock in the initial abtabt or the terminal $E_{n}$ from $I_{4,n}$, we can sort $d$ and push $t$ onto the stack to reach either the state $(att,bE_{n})$, the state $(btt,bE_{n})$ or a state $(abtt,bR)$ where $R$ is a proper subordering of $E_{n}$.
+For $I_{2,n}$, $I_{3,n}$, $I_{4,n}$, and $I_{5,n}$, the only other possible options are to remove a sock of color $d$, $e$, or the second sock of color $b$, or, in $I_{4,n}$, the second sock of color $t$.
+For the suborderings $adbtbE_{n}\subseteq I_{2,n}$ and $abdtbE_{n}\subseteq I_{3,n}$, first sorting $d$ and then $b$ leads to the state $(at,E_{n})$ which is sortable by Observation 1.
+For all other possibilities, if we first sort $b$ and then $d$ and $e$ if they are present, we either enter the state $(at,E_{n})$, the state $(att,E_{n})$, or the state $(at,tE_{n})$ and each of these states are sortable by Observation 1.
+For $I_{1,n}$, the only other options are to remove one of the two socks of color $d$.
+If the first $d$ is removed, then we can sort $b$ followed by $d$ which leaves $(at,E_{n})$, which is sortable by Observation 1.
+If the second $d$ is removed, then we can again arrive at the state $(at,E_{n})$ by first sorting $d$, then sorting $b$, and then pushing $t$ onto the stack.
+3. Overview of the proof of Theorem 2
+Our proof of Theorem 2 is inherently algorithmic, and we explicitly describe a deterministic recursive algorithm in Appendix A which is based on the proof.
+Central to our proof is the following notion of a greedy algorithm.
+3.1. The greedy algorithm
+One’s natural instinct may be to iteratively sort the first sortable color appearing in the sock ordering; that is, the sortable color appearing in $(S,R)$ whose final sock appears in $SR$ before the final sock of any other sortable color in $(S,R)$.
+We refer to this process as the greedy algorithm, and have named the corresponding process in our algorithm as such.
+In some instances, the greedy algorithm may fail.
+In effect, it can trap certain socks in the stack which then leads to issues later on in the sorting process.
+For a simple example, consider the sock ordering abcdabcabcdabc; the greedy algorithm instructs us to sort the color $d$ first since $d$ is sortable in abcdabcabcdabc and the final sock of color $d$ appears before the final socks of any other sortable color (which, in this case, is only $a$).
+Sorting $d$ yields $(abc,abc)$, which is terminal.
+On the other hand, ignoring the greedy algorithm and instead sorting the color $a$ first yields $(bcd,bc)$.
+From this point, one may use the sorting sequence $\theta=d,c,b$ and observe $\theta(bcd,bc)=(\emptyset,\emptyset)$, and as such, the sorting sequence $\theta^{\prime}=a,d,c,b$ fully sorts abcdabcabcdabc.
+3.2. Proof overview
+Let $X$ be unsortable sock ordering. We prove that $X$ does not avoid $\Gamma$ by induction on the number of socks in $\Gamma$.
+We start by applying the greedy algorithm to $X$. As $X$ is unsortable, the algorithm returns a terminal state $(S,R)$ where at least one of $S$ or $R$ is nonempty.
+First suppose that the stack $S$ is empty or contains a color $x$ such that $(x,R)$ is terminal.
+This implies that either $R$ or $xR$ is trivially unsortable.
+Since it is relatively straightforward to show that a trivially unsortable pattern does not avoid $\Gamma$ (c.f.
+Lemma 8), we can easily handle this case.
+That is, we can show that $R$ or $xR$ does not avoid $\Gamma$, so $X$ does not avoid $\Gamma$.
+Therefore, we can assume that there are at least two colors on the stack $S$ and for no color $x\in S$ is $(x,R)$ terminal.
+Let $t$ be top color on the stack and $b$ the color that appears immediately below $t$ on the stack.
+So, $t\neq b$ and $bt\subseteq S$.
+Since the stack is not empty, the greedy algorithm must have sorted at least one color.
+We use $d$ to denote the last color sorted by the greedy algorithm before termination.
+One simplifying observation is that one of the following two possibilities is true (c.f. Lemma 11):
+  • $(bt,R)$ is terminal, or
+  • there exists a color $a$ on the stack, distinct from $b$ or $t$, such that $(abt,R)$ is terminal. (For technical reasons, in this case, we always assume $a$ is the lowest such color on the stack.)
+Suppose $(bt,R)$ is terminal.
+From this and the assumption that for every $x\in S$ the state $(x,R)$ is not terminal, we deduce that $R$ must contain $P$ where $P$ is one of a few possible subpatterns and $b$ is the only good sortable color in the subordering $btP$.
+For example, one possibility is that $R$ contains brbtbrbt where $r$ is some new color, and the only good sortable color in btbrbtbtbrbt is $b$ (in fact, $b$ is the only sortable color in btbrbtbtbrbt).
+If $b$ is blocked by a $d$-sandwich in $X$, then we can argue that a subpattern of $X$ is contained in $\Gamma$.
+Otherwise, we can argue that either XdX-d is unsortable or that $I_{1,n}$ is a subpattern of $X$ for some $n\geq 0$ (c.f.
+Lemma 7).
+Since $I_{1,n}\in\Gamma$, we are done in the second case.
+In the first case, the induction hypothesis implies that XdX-d, and hence $X$, contains a subpattern that appears in $\Gamma$.
+Now assume the other possibility holds.
+That is, we have $abt\subseteq S$ such that $(abt,R)$ is terminal.
+This case is somewhat more complicated, but we can sometimes argue in a very similar manner to the previous case.
+However, sometimes the terminal state $(S,R)$ does not give us enough information about the original ordering $X$.
+For example, it could be that $S=abt$ and $R=at$.
+When we are in this situation, we backtrack to immediately before the greedy algorithm pushed the final sock of color $b$ onto the stack.
+We then have sufficient information to complete the proof by making an argument similar to the previous cases.
+4. Observations and Lemmas
+In the enumeration of our basis $\Gamma$, the five infinite classes have a common structure. In some sense, the following lemma explains the origin of this commonality.
+Lemma 3.
+Let $X$ be a sock ordering that, for some $n\geq 0$, has distinct colors $b$ and $c$ where the final $b$ precedes the final $c$ in $X$.
+Define $A$ and $C$ to be the strings such that $X=AbC$ where $b\notin C$ and $c\in C$.
+Let $\theta$ be a nonempty sorting sequence for $X$ that sorts a color in $C$ and suppose $c$ is the color in $C$ that is sorted first by $\theta$.
+Taking $(S,R):=\theta(X)$, suppose $b\in S$ and that once $c$ is sorted, $b$ never appears as the top color on the stack.
+Also, suppose that $t_{0},\dots,t_{n}$ (in this order) are the $n+1$ colors which appear immediately above the color $b$ in the stack after $c$ is sorted.
+Then
+$I(c,t_{0},\dotsc,t_{n})R\subseteq C.$
+Furthermore, the colors $t_{0},\dotsc,t_{n}$ are distinct and disjoint from $b$ and $c$.
+Proof.
+We will prove the first statement by induction on $n$.
+For the base case, assume $n=0$, so $t_{0}=t_{n}$ is the only color that appears directly above $b$ after $c$ is sorted.
+Since $c$ follows the last $b$ and $b$ can never appear as the top color on the stack after $c$ is sorted, the color $t_{0}=t_{n}$ must appear above $b$ after $c$ is sorted.
+Therefore, we have
+$C\supseteq t_{0}cR=I(c,t_{0},\dotsc,t_{n})R.$
+Now, suppose $n\geq 1$.
+Let $d$ be the color that is sorted that placed the color $t_{n}$ directly above the color $b$ on the stack.
+And let $\theta_{d}$ be the initial sequence of $\theta$ up to but not including $d$.
+Let $(S^{\prime},R^{\prime}):=\theta_{d}(X)$.
+Note, during the execution of $\theta_{d}$, the color $b$ is never the top color on the stack and that $t_{0},\dotsc,t_{n-1}$ is the sequence of colors that appear immediately above the color $b$ after $c$ is sorted.
+Therefore, by induction, we have
+$I(c,t_{0},\dotsc,t_{n-1})R^{\prime}\subseteq C.$
+Recall that $t_{n-1}$ appears immediately above $b$ on the stack $S^{\prime}$, but after $d$ is sorted $t_{n}$ is the color above $b$ on the stack.
+Since $t_{n-1}\neq t_{n}$, this implies that $d=t_{n-1}$.
+Furthermore, we also have $t_{n}t_{n-1}R\subseteq R^{\prime}$.
+This completes the proof of the first statement because
+$C\supseteq I(c,t_{0},\dotsc,t_{n-1})R^{\prime}\supseteq I(c,t_{0},\dotsc,t_{n-1})t_{n}t_{n-1}R=I(c,t_{0},\dotsc,t_{n-1},t_{n})R.$
+For the final statement, note that, for every $i\in\{0,\dotsc,n\}$, the color $t_{i}\neq b$ by definition and $t_{i}\neq c$ since it appears in the stack after $c$ is sorted.
+Furthermore, the colors $t_{0},\dotsc,t_{n}$ must be $n+1$ distinct colors because the only way for a color to removed from the stack is for it to be sorted.
+∎
+We now present a few simple observations and lemmas that will be useful throughout.
+Lemma 4.
+Let $X$ be a sock ordering.
+Suppose that $x\in X$ is unsortable and that $x$ is one of the first two colors to appear in $X$.
+Then any color that appears after the final $x$ in $X$ is blocked by an $x$-sandwich.
+Proof.
+Let yzyyzy be a sandwich that blocks $x$ in $X$.
+Since $x$ is one of the first two colors to appear, $z$ cannot precede the first occurrence of $x$.
+Therefore, both $y$ and $z$ appear between the first and last occurrence of $x$.
+Therefore, any color that follows the final $x$ in $X$ is blocked by one of the sandwiches xyxxyx or xzxxzx.
+∎
+There is a natural ordering of the sandwiches that appear in a given sock ordering.
+If abaaba and xyxxyx are suborderings of a sock ordering $X$ we say that abaaba precedes xyxxyx if the final $a$ in abaaba precedes the final $x$ in xyxxyx; or when $a=x$ and the final socks in abaaba and xyxxyx are the same, the $b$ in abaaba precedes the $y$ in xyxxyx; or when $a=x$ and $b=y$ and the last two sock in abaaba and xyxxyx are the same, the first $a$ in abaaba precedes the first $x$ in xyxxyx.
+A sandwich abaaba is the first sandwich to appear in a sock ordering $X$ if for all sandwiches xyxxyx in $X$ which are distinct from abaaba, abaaba precedes xyxxyx.
+Observation 5.
+If abaaba is the first sandwich to appear in a sock ordering $X$, then any color that is not $a$ or $b$ is either sortable in $X$ or is blocked by the sandwich abaaba.
+Proof.
+Let $x$ be an unsortable color in $X$ that is neither $a$ nor $b$.
+Because $x$ is unsortable, it must follow a sandwich.
+Since abaaba is the first sandwich to appear, $x$ must follow abaaba.
+Therefore, $x$ is blocked by abaaba.
+∎
+Lemma 6.
+Let $(S,R)$ be a sorting state (so $S$ is sorted) and assume that $S$ has at least two distinct colors.
+Let $a\in S$ where $a$ is not the top color in $S$ and let $S^{\prime}$ be formed by removing the color $a$ from $S$.
+If $x$ is unsortable in $(S,R)$ and $x$ is sortable in $(S^{\prime},R)$, then
+  • $x=a$, or
+  • $ax\subseteq R$, or
+  • there exists a color $y$ distinct from the colors $a$ and $x$ such that $yx\subseteq R$ and $x$ is the top color, $a$ is immediately below the color $x$, and $y$ is immediately below the color $a$ on the stack $S$.
+Proof.
+Assume $x\neq a$.
+Because $x$ is sortable in $(S^{\prime},R)$, we either have $x\in R$, or that $x$ is the top color on $S^{\prime}$.
+Suppose $x$ is the top color on $S^{\prime}$.
+Because $a$ is not the top color on $S$, this implies that $x$ is the top color on $S$.
+Since $x$ is not sortable in $(S,R)$, we have $x\in R$ and a sandwich blocks $x$ in $SR$.
+So, in all cases, we have $x\in R$ and a sandwich blocks $x$ in $SR$.
+Let yzyyzy be the first sandwich in $SR$ that blocks $x$.
+Since $S$ is sorted, yzyyzy is not a subordering of $S$, so we have $yx\subseteq R$.
+Because $x$ is sortable in $S^{\prime}R$, the color $a$ is either $y$ or $z$.
+If $a$ is $y$, then $ax\subseteq R$, so assume $a$ is $z$.
+Since $yzy=yay$ does not block $x$ in $S^{\prime}R$, we must have $ya\subseteq S$.
+Recall that, by assumption, $a$ is not the top color in $S$, so let $w$ be the color that appears directly above $a$ in $S$.
+Note that $w\neq y$ because $S$ is sorted.
+If $w\neq x$, then ywyywy blocks $x$ in $S^{\prime}R$, so we have $w=x$.
+Since $x$ is sortable in $(S^{\prime},R)$ and $x$ is on $S^{\prime}$, the color $x$ must be the top color on the stack $S^{\prime}$.
+Therefore, $x$ is the top color on $S$.
+We now have that $x$ is the top color and $a$ is the color immediately below $x$ on $S$ and that $yx\subseteq R$.
+Therefore, we only need to show that the color $y$ is directly below the color $a$ in $S$.
+This follows from the observation that if there exists a color $u$ that is distinct from $y$, $a$, or $x$ that is between $y$ and $a$ in $S$, then the sandwich yuyyuy blocks $x$ in $S^{\prime}R$, a contradiction.
+∎
+The following lemma (Lemma 7) is a crucial piece of our proof.
+Before we present its statement and proof, we provide the following informal discussion of the statement, the proof, and how it fits into the larger problem.
+Let $X$ be a sock ordering (either sortable or unsortable) and suppose that an attempted sorting fails because it trapped a color on the stack.
+That is, there is a color that we need to sort, but we cannot because it is beneath the top element on the stack.
+Let $a$ be a color we need to sort and let $d$ be the last color sorted in our attempted.
+The hypothesis of Lemma 7 is how we characterize this situation precisely.
+When this happens, it might have been difficult to sort $a$ before sorting $d$ because there is a $d$-sandwich that blocks $a$ in the original ordering $X$ (c.f.
+(I)).
+In this case, we can often argue that there was essentially no way to avoid trapping $a$ on the stack, and that $X$ is unsortable because it does not avoid $\Gamma$ (c.f.
+Lemmas 9 and 10).
+If there is no $d$-sandwich blocking $a$ in $X$, then it might be natural to assume that sorting the color $d$ was not the problem with our attempt.
+That is, maybe we sorted the wrong color earlier on in the process.
+With this in mind, it might be reasonable to recursively try to sort the ordering formed by removing $d$ from $X$, XdX-d.
+If that fails, then the recursive assumption implies that XdX-d is itself unsortable, which further implies that the original ordering, $X$, is unsortable (c.f.
+(II)).
+So, we only need to decide how to proceed when we can sort XdX-d.
+Let $\theta$ be any sorting sequence that fully sorts XdX-d.
+One possibility is that if we apply $\theta$ to the original sequence $X$, then at some point the color $d$ appears only at the top of the stack (c.f.
+(III)).
+Note that this implies that $X$ itself is sortable; if we sort $d$ at this point, then the state matches the corresponding state reached while sorting XdX-d.
+(For the proof of Theorem 2, we could replace (III) with the simpler statement “$X$ is sortable”, but that would not imply that when an attempt to sort XdX-d succeeds, we necessarily have a sorting sequence that can be used to sort $X$, so it would not imply that our recursive algorithm is correct.)
+So let us assume otherwise.
+(That is, we assume that neither (I), (II), nor (III) hold.) This is how we begin the proof of the lemma.
+With this assumption, we can quickly argue that $X=YDW$ where $D$ contains all of the socks of color $d$ and no sock color besides $d$ appears in $D$.
+Let $\theta$ be any sorting sequence that fully sorts Xd=YWX-d=YW.
+We then argue that $\theta$ can be used to sort $X$ up to and including the color $a$.
+Furthermore, when the sequence $\theta$ sorts $a$, it must push $d$ onto the stack.
+We then analyze the situation immediately before a sandwich would appear on the stack when sorting $X$ with $\theta$ (this must happen because the fact that (III) does not hold implies that $\theta$ cannot be used to sort every color in $X$ except $d$, since then $d$ would be the only element on the stack).
+Using Lemmas 3 and 6, we can then determine that $I_{1,n}\in\Gamma$ is a subpattern of $X$ where $n+1$ is the number of distinct colors that appear immediately above $d$ on the stack after we sort the color $a$ (c.f.
+(IV)).
+Lemma 7.
+Let $X$ be a sock ordering and let $\overline{\theta}$ be a nonempty sorting sequence for $X$.
+Define $(S,R):=\overline{\theta}(\emptyset,X)$ and let $d$ be the last color in $\overline{\theta}$.
+Suppose that there exists a color $a$ distinct from $d$ and suborderings $S^{\prime}\subseteq S$ and $R^{\prime}\subseteq R$ such that $a$ is not the top color of $S^{\prime}$, $a$ appears in $R^{\prime}$, and $a$ is the only good sortable color in $S^{\prime}R^{\prime}$.
+Then, one of the following is true:
+  1. (I)
+    there is a $d$-sandwich that blocks $a$ in $X$;
+  2. (II)
+    XdX-d is unsortable;
+  3. (III)
+    XdX-d is sortable and if a sequence $\theta$ fully sorts XdX-d, then when we attempt to sort $X$ with the sequence $\theta$, there is a stage where the color $d$ only appears as the top color on the stack; or
+  4. (IV)
+    for some $n\geq 0$, $I_{1,n}$ is a subpattern of $X$.
+Proof.
+Assume that (I), (II), and (III) do not hold.
+We will show that (IV) must hold.
+Let $x$ be the top color on $S^{\prime}$ and $D\subseteq X$ be such that $X=YDW$ where the first element of $D$ is the first $d$ in $X$ and the final element of $D$ is the final $d$ in $X$.
+Note that the subordering $R$ must follow the last $d$ in $X$, so $R\subseteq W$.
+Similarly, everything in $S$ must precede the last $d$ in $X$, so $S\subseteq YD$.
+Since $R\subseteq W$ and $a\in R^{\prime}\subseteq R$, we have $a\in W$, so the fact there are no $d$-sandwiches that block $a$ in $X$ implies that the only colors that could appear in $D$ are $a$ and $d$.
+In particular, the color $x$ does not appear in $D$, so the fact that $a$ is beneath $x$ on $S^{\prime}$ and $S^{\prime}\subseteq S\subseteq YD$ implies that $ax\subseteq Y$.
+Note that because $d$ is sorted by $\overline{\theta}$ but $a$ and $x$ are not, there is no sandwich that consists only of $a$ and $x$ that blocks $d$ in $X$.
+Therefore, the color $a$ does not appear in $D$ as otherwise the fact that $ax\subseteq Y$ would imply the existence of a sandwich of the form axaaxa before the final $d$ in $X$.
+As such, $d$ is the only color that appears in $D$ and we can assume without loss of generality that the color $d$ only appears once in $X$, so $D=d$.
+By assumption, the statements (II) and (III) are both false, so there exists a sorting sequence $\theta$ that fully sorts Xd=YWX-d=YW such that when we apply $\theta$ to $X=YDW=YdW$ the color $d$ never appears only as the top color on the stack.
+Let $\theta_{a}$ be the list of initial colors in $\theta$ up to, but not including, $a$ (so $\theta_{a}$ is empty if and only if $\theta$ begins with $a$).
+Note that no color in $S^{\prime}R^{\prime}$ could appear before $a$ in $\theta$ because $a$ is the only good sortable color in $S^{\prime}R^{\prime}\subseteq YW$ and $\theta$ fully sorts $YW$.
+We claim that this implies that no color that appears after $ax$ in $YW$ is in $\theta_{a}$.
+To see this, note that sorting such a color would put $ax$ onto the stack, but that would make it impossible to sort $a$ before sorting $x$.
+In particular, because $ax\subseteq Y$, this means that no color that appears in $W$ is in $\theta_{a}$.
+Therefore, $a$ is the color in $W$ that is sorted first by $\theta$, and $\theta_{a}$ is a sorting sequence for $X=YdW$.
+Furthermore, if we let $(S_{1},R^{\prime}_{1})=\theta_{a}(YW)$, then $(S_{1},R_{1})=\theta_{a}(YdW)$ where $R^{\prime}_{1}=R_{1}-d$.
+Therefore, because $a$ is sortable in $(S_{1},R^{\prime}_{1})$ and no $d$-sandwich blocks $a$ in $X$, the color $a$ is also sortable in $(S_{1},R_{1})$.
+Recall that no color that appears after $ax$ is in $\theta_{a}$ and $x$ is not in $\theta_{a}$.
+This, with the fact that $axda\subseteq SdR\subseteq X$, implies that $xda\subseteq R_{1}$.
+Therefore, sorting $a$ in the state $(S_{1},R_{1})$ pushes $xd$ onto the stack.
+Let $y$ be the color that immediately precedes $d$ in $R_{1}$ (so we could have $y=x$).
+Note that since $ax\subseteq Y$ and either $x=y$ or $xy\subseteq Y$, in all cases, we have
+$ay\subseteq Y.$ (1)
+Recall that, by the previous arguments, $d$ is pushed on the stack when we attempt to apply $\theta$ to $X$, so $\theta$ cannot be a sorting sequence for $X$ as otherwise we would have $(d,\emptyset)=\theta(X)$, a contradiction to the fact that the color $d$ never appears as the top color on the stack.
+Therefore, there must exist a color in $\theta$ that causes the stack to be unsorted when $\theta$ is applied to $X$.
+Let $z$ be the first such color and let $\theta_{z}$ be the list of initial colors in $\theta$ up to but not including $z$.
+That is, $\theta_{z}$ is the longest initial sequence of $\theta$ that is a sorting sequence for $X$.
+Let $(S_{2},R_{2}):=\theta_{z}(X)$.
+Note that by our previous arguments, $a$ must appear in $\theta_{z}$ and $a$ is the color in $W$ that is sorted first by $\theta_{z}$ and $d\notin W$.
+By Lemma 3, we have
+$W\supseteq I(a,t_{0},\dotsc,t_{n})R_{2}=t_{0}aI(t_{0},\dotsc,t_{n})R_{2},$ (2)
+where $t_{0},\dotsc,t_{n}$ is the sequence of colors that appear above $d$ on the stack after $a$ is sorted.
+Note that since $y$ is immediately below $d$ on the stack after the sequence $\theta_{a}$ is applied to $X$ and $d$ is never removed from the stack while $\theta_{z}$ is applied to $X$, the color $y$ is immediately below $d$ on the stack $S_{2}$.
+Let $S_{2}^{\prime}:=S_{2}-d$ and note that $(S_{2}^{\prime},R_{2})=\theta_{z}(X-d)$.
+By the selection of $z$, the color $z$ is unsortable in $(S_{2},R_{2})$, but $z$ is sortable in $(S^{\prime}_{2},R_{2})$.
+Since $d\notin S^{\prime}_{2}R_{2}\subseteq X-d$, we have $z\neq d$ and $dz\nsubseteq R_{2}$, so Lemma 6 implies that $y$ is the top color, $d$ is immediately below the color $y$, and $z$ is the color immediately below $d$ on the stack $S_{2}$ and that $yz\subseteq R_{2}$.
+Note that this means that $z=t_{n}$, since $t_{n}$ is defined to be the last color that appears immediately above $d$ on the stack when we sort $X$ with $\theta_{z}$.
+So, $yt_{n}\subseteq R_{2}$.
+This, with (1) and (2), implies that
+$X=YdW\supseteq aydt_{0}aI(t_{0},\dotsc,t_{n})R_{2}\supseteq aydIt_{0}a(t_{0},\dotsc,t_{n})yt_{n}\cong I_{1,n}.\qed$
+The following three lemmas are needed in the proof of our main theorem to show that a given unsortable sock ordering has a subpattern that is in $\Gamma$.
+Unsurprisingly, these proofs have a number of cases and are somewhat technical, so we defer their proofs until Section 6.
+Lemma 8.
+If $X$ is trivially unsortable, then a subpattern of $X$ is in $\Gamma$.
+Lemma 9.
+Let $X$ be a sock ordering and $n\geq 0$.
+Suppose that a nonempty sorting sequence produces $(S,R)$ when it is applied to $X$ and let $d$ be the last color sorted.
+Let $a,b,t_{0},\dots,t_{n}$ be distinct colors in $X$ which are different from $d$ and define $t:=t_{0}$.
+If $abt\subseteq S$ and either
+  • $atb\subseteq R$ or $abt\subseteq R$ and $a$ is blocked by a $d$-sandwich in $X$, or
+  • $bI(t_{0},\dotsc,t_{n})at_{n}\subseteq R$ and $b$ is blocked by a $d$-sandwich in $X$,
+then $X$ is unsortable. If, in addition, every proper subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
+Lemma 10.
+Let $X$ be a sock ordering and suppose that a nonempty sorting sequence produces $(S,R)$ when it is applied to $X$ where $d$ is the last color sorted by the sequence.
+Further suppose that the following two conditions hold:
+  • For every $x\in S$, the state $(x,R)$ is not terminal.
+  • There exist distinct colors $a$ and $t$ in $S$ such that $(at,R)$ is terminal.
+Then there exists $P\subseteq R$ such that $a$ appears in $P$ and $a$ is the only good sortable color in $atP$.
+Furthermore, if $a$ is blocked by a $d$-sandwich in $X$, then $X$ is unsortable.
+If, in addition, every proper subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
+The following lemma shows that if we reach the state $(S,R)$ that is terminal and $S$ has at least three colors, then, in a certain sense, we only need to focus on at most three colors in $S$.
+Furthermore, two of the three colors can be the top two colors on $S$.
+Lemma 11.
+Let $S$ and $R$ be sock orderings where $S$ is sorted and at least three distinct colors appear in $S$.
+Suppose $(S,R)$ is terminal.
+If $t$ is the top color and $b$ is the color immediately below $t$ on $S$, then there exists a color $a\in S$ distinct from $b$ and $t$ such that $(abt,S)$ is terminal.
+Proof.
+Suppose $(S,R)$ is a counterexample.
+We can assume without loss of generality that no colors are repeated in $S$.
+So we can write $S=x_{1}x_{2}\dotsm x_{m}bt$ where $x_{1},\dotsc,x_{m}$, $b$, and $t$ are distinct colors.
+Let $p$ be as small as possible such that there exists $y_{1},\dotsc,y_{p}$, a subsequence of $x_{1},\dotsc,x_{m}$, such that $(y_{1}y_{2}\dotsm y_{p}bt,R)$ is terminal.
+Since $(S,R)$ is a counterexample, we have $p\geq 2$.
+If $p=2$, let $W=b$ and otherwise let $W=y_{3}\dotsm y_{p}b$.
+By the minimality of $p$, for $i\in\{1,2\}$, there exists some $z_{i}$ that is sortable in $(y_{i}Wt,R)$.
+For $i\in\{1,2\}$, since $z_{i}$ is sortable in $(y_{i}Wt,R)$, we have $z_{i}\notin y_{i}W$.
+In particular, this implies that $z_{i}\neq y_{i}$ and $z_{i}\neq b$.
+Furthermore, since $z_{i}$ is sortable in $(y_{i}Wt,R)$, there is no $y_{i}by_{i}$ sandwich blocking $z_{i}$ in $y_{i}WtR$, so $y_{i}z_{i}\nsubseteq R$.
+We will now show that $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
+Recall that $(y_{1}y_{2}Wt,R)=(y_{1}y_{2}\dotsm y_{p}bt,R)$ is terminal, so $z_{1}$ is not sortable in $(y_{1}y_{2}Wt,R)$.
+Therefore, we can apply Lemma 6 to deduce that $z_{1}$ is unsortable in $(y_{2}Wt,R)$, because $z_{1}\neq y_{1}$, $y_{1}z_{1}\nsubseteq R$, and $y_{2}$ is not the second color on the stack $y_{1}y_{2}Wt$.
+Since $z_{1}$ is unsortable in $(y_{2}Wt,R)$, $y_{1}z_{1}\nsubseteq R$, and $z_{1}\notin W$, either $z_{1}=y_{2}$ or there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ in which $y_{1}$ does not appear.
+If $z_{1}=y_{2}$, then, because $z_{1}=y_{2}$ is sortable in $(y_{1}Wt,R)$, we have $y_{2}\in R$, so $y_{1}y_{2}=y_{1}z_{1}\nsubseteq R$ implies that either $y_{1}\notin R$ or that $y_{1}$ is blocked by a sandwich of the form $y_{2}by_{2}$ in $y_{2}WtR$.
+So, $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
+If there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ in which $y_{1}$ does not appear, then $y_{1}z_{1}\nsubseteq R$ implies that $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
+So, in all cases, $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
+Since $y_{1}$ is not sortable in $(y_{2}Wt,R)$, we have $y_{1}\neq z_{2}$.
+This, with the fact that $z_{2}$ is not sortable in $(y_{1}y_{2}Wt,R)$ but is sortable $(y_{2}Wt,R)$ together with Lemma 6, yields $y_{1}z_{2}\subseteq R$.
+Since $y_{1}z_{2}\subseteq R$ and $y_{2}z_{2}\nsubseteq R$, $y_{2}$ is not sortable in $(y_{1}Wt,R)$, so $y_{2}\neq z_{1}$.
+Using this fact with Lemma 6, allows us to conclude that $y_{2}z_{1}\subseteq R$ because $z_{1}$ is not sortable in $(y_{1}y_{2}Wt,R)$ but is sortable $(y_{1}Wt,R)$.
+Since $y_{1}z_{2}\subseteq R$, $y_{2}z_{2}\nsubseteq R$, and $y_{2}z_{1}\subseteq R$, we have $y_{1}z_{1}\subseteq R$.
+But we have already established that $y_{1}z_{1}\nsubseteq R$.
+∎
+Our last lemma is simple to prove despite its somewhat technical statement.
+Informally speaking, we use the lemma in the following situation.
+Suppose there are two distinct colors, say $x$ and $y$, and the final $x$ precedes the final $y$.
+We would expect the greedy algorithm to sort $x$ before it sorts $y$, but sometimes this might not occur.
+One reason $y$ might be sorted before $x$ is that, when $y$ is sorted, $x$ is blocked by a sandwich.
+Note that, when this occurs, the sandwich blocking $x$ must be a $y$-sandwich because the greedy algorithm sorts $y$ and the final $y$ appears after the final $x$.
+The other possibility is that $x$ cannot be sorted because it already appears below the top color on the stack.
+This lemma is used when we are in this second case.
+Lemma 12.
+Let $S$ and $R$ be sock orderings where $S$ is sorted and let $x$ and $y$ be distinct colors in $R$ such that the final $x$ in $R$ appears before the final $y$ in $R$.
+Suppose that, with respect to $(S,R)$, the color $y$ is sortable and the color $x$ is unsortable.
+If no $y$-sandwich blocks $x$ in $SR$, then $y$ is the top color and $x$ is the color immediately below $y$ on $S$.
+Furthermore, no color except possibly $y$ precedes $x$ in $R$.
+Proof.
+Because $y$ is sortable in $SR$, every sandwich in $SR$ that appears before the final $y$ in $R$ must be a $y$-sandwich.
+Therefore, because the final $x$ in $R$ precedes the final $y$ in $R$ and $x$ is not blocked by a $y$-sandwich in $SR$, the color $x$ is not blocked by a sandwich in $SR$.
+The fact that $x$ is unsortable then implies that $x$ is below the top color on the stack $S$.
+Since $y$ is sortable in $SR$, there is no $x$-sandwich that blocks $y$ in $SR$.
+Therefore, $y$ is the top color and $x$ is the color immediately below $y$ on $S$.
+Furthermore, no color except possibly $y$ can precede $x$ in $R$.
+∎
+5. Proof of Theorem 2
+The proof is by induction on the number of socks in $X$.
+The base case is vacuously true.
+So assume that $X$ is an unsortable sock ordering.
+With the induction hypothesis, we can assume that there does not exist a proper subordering of $X$ that is unsortable.
+Assume for a contradiction that $X$ avoids $\Gamma$.
+Claim 13.
+Let $\theta$ be a nonempty sorting sequence for $X$, let $(S,R):=\theta(X)$, and let $d$ be the last color in $\theta$.
+If there are distinct colors $a$ and $d$ and substrings $S^{\prime}\subseteq S\subseteq X$ and $R^{\prime}\subseteq R\subseteq X$ such that
+  • $a$ is not the top color of $S^{\prime}$,
+  • $a\in R^{\prime}$, and
+  • $a$ is the only good sortable color in $S^{\prime}R^{\prime}$,
+then $X$ contains a $d$-sandwich that blocks $a$.
+Proof.
+This claim is simply a restatement of Lemma 7.
+To see this, first note that the hypothesis of the claim is essentially the same as the hypothesis of Lemma 7.
+Furthermore, Lemma 7(II) cannot hold because we have assumed that every proper subordering is sortable; Lemma 7(III) cannot hold because $X$ is unsortable; and Lemma 7(IV) cannot hold because we are assuming that $X$ avoids $\Gamma$.
+Therefore, Lemma 7(I) must hold.
+That is, there is a $d$-sandwich that blocks $a$ in $X$.
+∎
+Let $(S,R)$ be the output of the greedy algorithm (as described in Section 3.1) applied to $X$. This implies that $(S,R)$ is terminal. Since $X$ is unsortable, we have $R\neq\emptyset$.
+Case 1: $S=\emptyset$ or there exists $x\in S$ such that $(x,R)$ is terminal.
+If $S=\emptyset$, then $R$ is trivially unsortable.
+Similarly, if there exists $x\in S$ such that $(x,R)$ is terminal then $xR$ is trivially unsortable.
+In either case, Lemma 8 implies that a subpattern of $X$ is in $\Gamma$.
+From now on, we will assume that we are not in Case 1.
+That is, for every $x\in S$, some color in $xR$ is sortable.
+Since $(S,R)$ is terminal, this implies that there are at least two colors on $S$.
+Define $t$ to be the top color in $S$, and let $b$ be the color in $S$ appearing immediately under $t$.
+We can also now let $d$ be the last color sorted by the greedy algorithm.
+Case 2: $(bt,R)$ is terminal.
+By Lemma 10, there exists $P\subseteq R$ such that $b$ appears in $P$ and such that $b$ is the only good sortable color in $btP$.
+This means that the hypothesis of Claim 13 is satisfied with $b$, $d$, btbt, and $P$ playing the roles of $a$, $d$, $S^{\prime}$, and $R^{\prime}$, respectively.
+Therefore, $b$ is blocked by a $d$-sandwich in $X$.
+Note that this is the hypothesis of the second statement in Lemma 10.
+The final statement of Lemma 10 then implies that a subpattern of $X$ is in $\Gamma$, a contradiction.
+Case 3: $(bt,R)$ is not terminal. With Lemma 11, we can define $a$ to be the lowest color on the stack $S$ such that $(abt,R)$ is terminal.
+Claim 14.
+$at\subseteq R$.
+Proof.
+Suppose, by way of contradiction, that $at\nsubseteq R$.
+Then, because $t$ is unsortable in $(abt,R)$ and $t\neq a$ and $a$ is not the second color on the stack abtabt, Lemma 6 implies that $t$ is unsortable in $(bt,R)$.
+By the case, $(bt,R)$ is not terminal, so since both $b$ and $t$ are not sortable in $(bt,R)$, there must exist a color $y\in R$ that is not $b$ or $t$ which is sortable in $(bt,R)$.
+If $y=a$, then, because $at\nsubseteq R$, the color $y=a$ must follow the last occurrence of $t$ in $btR$.
+If $y\neq a$, then, since $y$ is sortable in $(bt,R)$ and unsortable in $(abt,R)$, Lemma 6 implies that $ay\subseteq R$.
+Because $at\nsubseteq R$, this implies that $y$ follows the last occurrence of $t$ in $btR$.
+So, in both cases, $y$ follows the last occurrence of $t$ in $btR$.
+This with Lemma 4 and the fact that $t$ is the second color in $btR$ and $t$ is unsortable in $btR$ give us that $y$ is unsortable in $btR$, a contradiction.
+∎
+Case 3.1: There exists a color $z\in S$ distinct from $a$ and $t$ such that $azt\subseteq S$ and $z\in R$.
+By Claim 14, we can assume that either atzatz, aztazt, or $zat=zI(t)at$ is in $R$.
+First suppose that either atzatz or aztazt is $R$.
+Because $a$ is the only good sortable color in both aztatzaztatz and aztaztaztazt, Claim 13 implies that $a$ is blocked by a $d$-sandwich.
+Since every proper subordering of $X$ is unsortable, the last statement of Lemma 9 implies that $X$ does not avoid $\Gamma$, a contradiction.
+The other case is when $zat=zI(t)at$ is in $R$ and the argument is similar.
+Since $z$ is the only good sortable color in aztzataztzat, Claim 13 implies that $z$ is blocked by a $d$-sandwich.
+So, like the previous case, Lemma 9 implies that $X$ does not avoid $\Gamma$, a contradiction.
+Case 3.2: For every $z\in S$ distinct from $a$ and $t$ such that $azt\subseteq S$, we have $z\notin R$.
+Recall that $abt\subseteq S$.
+This and the case imply that $b\notin R$.
+Let $A$ and $C$ be such that $X=AbC$ and $b\notin C$.
+Because $b\in S$ and $b\notin R$, there exists a color in $C$ that is sorted by the greedy algorithm.
+Let $c\in C$ be the color in $C$ that is sorted first by the greedy algorithm and let $(S^{\prime},A^{\prime}bC)$ be the state immediately before $c$ is sorted.
+Note that we have
+$S^{\prime}A^{\prime}\subseteq A.$ (3)
+Recall that the color $b$ is never sorted by the greedy algorithm and note that, after $c$ is sorted, the color $b$ only appears on the stack. Therefore, because $ab\subseteq S$, we have
+$a\in S^{\prime}A^{\prime}$ (4)
+Furthermore, since $b$ is not sorted by the greedy algorithm, the color $b$ can never appear as the top color on the stack after $c$ is sorted by the greedy algorithm.
+For $n\geq 0$, let $t_{0},\dotsc,t_{n}$ (in that order from bottom-to-top) be the $n+1$ distinct colors that appear immediately above $b$ after $c$ is sorted by the greedy algorithm.
+Note that then $t=t_{n}$, so Claim 14 implies that $at_{n}=at\subseteq R$.
+With Lemma 3, we have
+$C\supseteq I(c,t_{0},\dotsc,t_{n})R\supseteq I(c,t_{0},\dotsc,t_{n})at_{n}.$ (5)
+Note that if $ca\subseteq A$, then, with (5), we have
+$X=AbC\supseteq cabI(c,t_{0},\dotsc,t_{n})at_{n}=cabt_{0}cI(t_{0},\dotsc,t_{n})at_{n}\cong I_{1,n}.$
+So, for the remainder of the proof, we will assume
+$ca\nsubseteq A.$ (6)
+Claim 15.
+Suppose that there is a color $z\in A^{\prime}bC$ distinct from $a$ and $c$ such that $z$ is not blocked by a $c$-sandwich and the final $z$ appears before the final $c$ in $A^{\prime}bC$.
+Then $z$ is the color immediately beneath the top color on the stack $S^{\prime}$ and $azc\subseteq S^{\prime}$ and $zI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC$.
+Proof.
+Recall that, in the state $(S^{\prime},A^{\prime}bC)$, the greedy algorithm sorted the color $c$.
+Since the final $z$ precedes the final $c$ in $S^{\prime}A^{\prime}bC$, this implies that, with respect to the state $(S^{\prime},A^{\prime}bC)$, the color $c$ is sortable and the color $z$ is not sortable.
+By Lemma 12 (with $S^{\prime}$, $A^{\prime}bC$, $c$ and $z$ playing the roles of $S$, $R$, $y$, and $x$, respectively), $c$ is the top color and $z$ is the color immediately below $c$ on the stack $S^{\prime}$ and no color except possibly $c$ precedes $z$ in $A^{\prime}bC$.
+So, with (5), we have
+$A^{\prime}bC\supseteq zC\supseteq zI(c,t_{0},\dotsc,t_{n})at_{n}.$
+Recall that $a\in S^{\prime}A^{\prime}\subseteq A$ by (4) and (3) and note that (6) implies that the final $a$ in $A$ precedes the first $c$ in $A$.
+Therefore, the assumption that $c$ is on the stack $S^{\prime}$ implies that $a\subseteq S^{\prime}$ as well.
+Furthermore, $c$ is the top color on $S^{\prime}$ and $z$ is the color immediately below $c$ on $S^{\prime}$, we have $azc\subseteq S^{\prime}$.
+∎
+Claim 16.
+There are at least three colors in $S^{\prime}$ and when we define $y$ to be the color immediately beneath the top color in the stack $S^{\prime}$, we have $y\neq a$, $y\neq c$, $ayc\subseteq S^{\prime}$, and $yI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC$.
+Proof.
+If $b$ is not blocked by a $c$-sandwich, then because $b$ is distinct from $a$ and $c$ and the final $b$ occurs before the final $c$ in $S^{\prime}A^{\prime}bC$, the hypothesis of Claim 15 is satisfied with $b$ playing the role of $z$.
+This implies that $y=b$, and that the desired conclusion holds.
+So we can assume that $b$ is blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$.
+Let $\hat{y}$ be the other color in the first $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$.
+By definition,
+$\hat{y}\neq c.$ (7)
+We will show that $\hat{y}=y$ and that the desired conclusion holds.
+To see this, first note that we must have $\hat{y}c\subseteq S^{\prime}A^{\prime}\subseteq A$ and $c\hat{y}\subseteq S^{\prime}A^{\prime}\subseteq A$.
+Furthermore, (6) and $c\hat{y}\subseteq S^{\prime}A^{\prime}\subseteq A$ imply that
+$\hat{y}\neq a$ (8)
+and with (4)
+$a\hat{y}\subseteq S^{\prime}A^{\prime}.$ (9)
+If the final $\hat{y}$ follows the final $c$ in $S^{\prime}A^{\prime}bC$, then, after $c$ is sorted, the color $\hat{y}$ must appear off the stack and, with (9), we will also have $a\hat{y}b$ on the stack.
+Then, the fact that $b$ is never sorted by the greedy algorithm implies that $\hat{y}$ is never sorted by the greedy algorithm.
+Furthermore, since we cannot have a $\hat{y}b\hat{y}-$sandwich on the stack, we must have $a\hat{y}bt_{n}\subseteq S$ and $\hat{y}\in R$.
+But this contradicts the fact that we are in Case 3.2.
+Therefore, the final $\hat{y}$ precedes the final $c$ in $S^{\prime}A^{\prime}bC$.
+Recall that $\hat{y}$ is the color other than $c$ that is in the first $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$.
+Since $b\notin C$, we have $\hat{y}b\subseteq S^{\prime}A^{\prime}b$.
+Because $c$ is sortable in $S^{\prime}A^{\prime}bC$, there is no sandwich consisting of $b$ and $\hat{y}$ which blocks $c$ in $S^{\prime}A^{\prime}bC$ and both the final $b$ and the final $\hat{y}$ in $S^{\prime}A^{\prime}bC$ precede the final $c$ in $S^{\prime}A^{\prime}bC$, this implies that $b\hat{y}\nsubseteq S^{\prime}A^{\prime}bC$.
+In particular, we have $\hat{y}\notin C$.
+We claim that the color $\hat{y}$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$.
+For a contradiction, assume that $\bar{y}$ is the color other than $c$ in such a sandwich.
+Since $\hat{y}\notin C$, this sandwich is in $S^{\prime}A^{\prime}b$, and $\bar{y}\hat{y}\subseteq S^{\prime}A^{\prime}b$.
+Since $b\hat{y}\nsubseteq S^{\prime}A^{\prime}bC$, this implies that $\bar{y}\neq b$.
+Therefore, there is a sandwich consisting of $\bar{y}\neq b$ and $c$ in $S^{\prime}A^{\prime}b$, so this sandwich blocks $b$ as well as $\hat{y}$.
+Furthermore, since there is no sandwich consisting of $\bar{y}$ and $\hat{y}$ that blocks $c$ in $S^{\prime}AbC$, we have $\hat{y}\bar{y}\nsubseteq S^{\prime}Ab$.
+But this contradicts the fact that $\hat{y}$ is in the first $c$-sandwich in $S^{\prime}A^{\prime}bC$ that blocks $b$.
+So we have established that the color $\hat{y}$ is not blocked by a $c$-sandwich and the final $\hat{y}$ precedes the final $c$ in $S^{\prime}A^{\prime}bC$.
+Also, (7) and (8) imply $\hat{y}$ is distinct from $a$ and $c$.
+So, the hypothesis of Claim 15 is satisfied with $\hat{y}$ playing the role of $z$.
+Therefore, $y=\hat{y}$ and the desired conclusion holds.
+∎
+Define $d^{\prime}$ to be the color sorted by the greedy algorithm immediately before reaching the state $(S^{\prime},A^{\prime}bC)$.
+By Claim 16, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC$.
+Note that $y$ and $c$ are the only sortable colors in $aycyI(c,t_{0},\dotsc,t_{n})at_{n}$, because every other color is blocked by a ycyycy sandwich.
+Furthermore, after sorting $c$ we have $(ayyt_{0},I(t_{0},\dotsc,t_{n})at_{n})$ as a substate which Observation 1 implies is unsortable.
+Therefore, with Observation 1, $y$ is the only good sortable color in $aycyI(c,t_{0},\dotsc,t_{n})at_{n}$.
+This means that the hypothesis of Claim 13 is satisfied with $y$, $d^{\prime}$, aycayc, and $yI(c,t_{0},\dotsc,t_{n})at_{n}$ playing the roles of $a$, $d$, $S^{\prime}$ and $R^{\prime}$, respectively.
+Therefore, we can assume that $y$ is blocked by a $d^{\prime}$-sandwich in $X$.
+Then, the last statement of Lemma 9 implies that a subpattern of $X$ is in $\Gamma$.
+6. Proofs of Lemmas 8, 9, and 10
+See 8
+Proof.
+Assume, for a contradiction, that $X$ is a minimal counterexample.
+Let abaaba be the first sandwich in $X$.
+Recall that since $X$ is trivially unsortable, every color is blocked by a sandwich.
+That is, for all $z\in X$, there exist distinct colors $x$ and $y$ in $X$ which are distinct from $z$ such that $xyxz\subseteq X$.
+This implies that every color must follow the first sandwich abaaba.
+Let $z$ be the first of the colors $a$ and $b$ that terminates in $X$.
+Let xyxxyx be the first sandwich in $X$ that blocks $z$.
+Case 1: $x=a$.
+Then $xyx=aya$ blocks $z$, so $z=b$.
+Suppose the $y$ in the sandwich ayaaya follows the first sandwich abaaba, then $X\supseteq abayaba\cong abacaba\cong\mathcal{T}_{1}$.
+Otherwise, because abaaba (and not ayaaya) is the first sandwich, we have $X\supseteq abya$.
+This abyaabya subordering is followed by $ba$ and the color $y$ in $X$.
+Therefore, either $X\supseteq abyabya\cong abcabca\cong\mathcal{T}_{3}$, $X\supseteq abyayba\cong abcacba\cong\mathcal{T}_{4}$, or $X\supseteq abyabay\cong abcabac\cong\mathcal{T}_{5}$.
+Case 2: $x=b$.
+Then $xyx=byb$ blocks $z$, so $z=a$.
+Suppose the $y$ in the sandwich bybbyb follows the first sandwich abaaba.
+Then $X\supseteq abaybab\cong abacbab\cong\mathcal{T}_{2}$.
+Otherwise, because abaaba is the first sandwich (and not ayaaya), we have $X\supseteq abya$.
+Since bybbyb blocks $z=a$ and the final $b$ follows the final $z=a$, the subordering babbab must follow the subordering abyaabya.
+Furthermore, since $y$ must follow the first sandwich, the color $y$ must follow this abyaabya.
+So, as in the previous case, abyaabya is followed by $ba$ and the color $y$.
+Therefore, as in the previous case, either $\mathcal{T}_{3}$, $\mathcal{T}_{4}$, or $\mathcal{T}_{5}$ is a subpattern of $X$.
+Case 3: $x\neq a$ and $x\neq b$.
+Case 3.1: The first $x$ appears before the first sandwich abaaba.
+Then, because abaaba is the first sandwich to appear, the color $x$ must also follow abaaba.
+Therefore, $X\subseteq xabaxab\cong abcbacb\cong\mathcal{T}_{6}$ or $X\subseteq xabaxba\cong abcbabc\cong\mathcal{T}_{7}$.
+Case 3.2: The first $x$ appears between the first and last $a$ in the first sandwich abaaba.
+By the case, an axaaxa sandwich precedes the first xyxxyx sandwich.
+Since the first sandwich that blocks $z$ is xyxxyx, we must have $z=a$.
+Note that, because xyxxyx blocks $z=a$, we also have $y\neq a$.
+The color $x$ must follow the $ab$ in the first sandwich abaaba, as otherwise an axaaxa sandwich would proceed the first abaaba sandwich.
+Since xyxxyx is not the first sandwich in $X$ and $y\neq a$, either $X\supseteq abxayx$ or $X\supseteq abxyax$.
+We can assume that both abaaba and xyxxyx in this subordering are the first such subordering to appear in $X$.
+If $y=b$, then the subordering is abxabxabxabx or abxbaxabxbax.
+But then bxbbxb precedes $xbx=xyx$, so $xbx=xyx$ is not the first sandwich which blocks $z=a$.
+Therefore, we can assume $y\neq b$.
+Therefore, the colors $a$ and $b$ are distinct from the colors $x$ and $y$, so either $X\supseteq abxayxab\cong abcadcab\cong\mathcal{T}_{8}$ or the subordering is abxyaxabxyax.
+So, we can assume the subordering is abxyaxabxyax.
+Recall that every color must follow the first abaaba sandwich, so every color must follow abxyaabxya.
+This with the fact that $a$ is neither $x$ nor $y$ implies that every color that is not $a$ is blocked by either axaaxa or ayaaya.
+Furthermore, $a$ is blocked by xyxxyx.
+Therefore, every color is blocked by one of the three sandwiches axaaxa, ayaaya, or xyxxyx.
+Since none of these three sandwiches includes the color $b$, if we remove the color $b$ from $X$ we have a trivially unsortable sock ordering which contradicts the minimality of $X$.
+Case 3.3: The first $x$ appears after the first sandwich abaaba.
+So, we have abaxyxabaxyx in $X$ where abaaba and xyxxyx in this subordering are the first such subordering to appear in $X$.
+If $y=a$, then we must have $z=b$, but then $axy=axa$ appears before xyxxyx, a contradiction to the fact that xyxxyx is the first sandwich that blocks $z=b$.
+Similarly, if $y=b$, then we must have $z=a$, but then $bxy=bxb$ blocks $z=a$ and appears before xyxxyx.
+Therefore, the colors $x$ and $y$ are distinct from the colors $a$ and $b$.
+Hence, $X\supseteq abaxyxba\cong abacdcba\cong\mathcal{T}_{9}$ or $X\supseteq abaxyxab\cong abacdcab\cong\mathcal{T}_{10}$.
+∎
+See 9
+Proof.
+First note that if a proper subordering of $X$ is unsortable, then $X$ is unsortable.
+Therefore, to prove the lemma, we can assume that every proper subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$, since this will imply that $X$ is unsortable.
+Assume for a contradiction that no subpattern of $X$ is in $\Gamma$.
+Let $T:=(t_{0},\dotsc,t_{n})$ and $P\in\{atb,abt,bI(T)at_{n}\}$ such that $P\subseteq R$.
+We may write $X={\bar{S}}d{\bar{R}}$ where $S\subseteq{\bar{S}}$ and $d\notin{\bar{R}}$ and $R\subseteq{\bar{R}}$.
+Note that $a$ is the only sortable color in abtatbabtatb and abtabtabtabt.
+We claim that $b$ is the only good sortable color in $abtbI(T)at_{n}$.
+To see this note that when $n=0$ the color $b$ is the only sortable color in $abtbI(T)at_{n}=abtbat$ and when $n\geq 1$, $b$ and $t$ are the only sortable colors in $abtbI(T)at_{n}$, but sorting $t=t_{0}$ leaves $(abbt_{1},I(t_{1},\dotsc,t_{n})at_{n})$ which is unsortable by Observation 1.
+Define
+$(z,y):=\begin{cases}(a,b)&\text{if $P\in\{atb,abt\}$,}\\ (b,a)&\text{if $P=bI(T)at_{n}$}.\end{cases}$
+We have the following.
+Claim 17.
+The color $z$ is blocked by a $d$-sandwich in $X$ and the color $z$ is not blocked by a sandwich consisting only of colors in $P$.
+Proof.
+The first conjunct is given by the hypothesis of the lemma.
+For the second conjunct, recall that, by the above discussion, $z$ is the only good sortable color in $abtP\subseteq X-d$.
+Therefore, because XdX-d is a proper subordering of $X$ and is therefore sortable, $z$ is not blocked by a sandwich consisting only of colors in $P$.
+∎
+Claim 18.
+One of the following holds.
+  1. (i)
+    one of dabtdabt, adbtadbt, abdtabdt, is a subordering of $\bar{S}$} or $abtde\subseteq{\bar{S}}$ for some color $e\notin P$; or
+  2. (ii)
+    $atab\subseteq{\bar{R}}$ or $tP\subseteq{\bar{R}}$; or
+  3. (iii)
+    for some color $e\notin P$, $e\in{\bar{S}}$ and $eP\subseteq{\bar{R}}$.
+Proof.
+Recall that by Claim 17, $z$ is blocked by a $d$-sandwich in $X$.
+That is, for some color $e$, we either have a sandwich of the form dedded which blocks $z$ or a sandwich of the form edeede which blocks $z$.
+Case 1: A sandwich of the form dedded blocks $z$.
+In this case, we will show that (i) holds, so suppose that neither dabtdabt, adbtadbt, nor abdtabdt is a subordering of $\bar{S}$}.
+Then, because $d\notin{\bar{R}}$ implies $de\subseteq{\bar{S}}$, we have $abtde\subseteq{\bar{S}}$.
+To show that (i) holds, we then only need to show that $e\notin P$.
+Since dedded blocks $z$, we have $e\neq z$.
+We also have $e\neq y$, since otherwise we would have a ytyyty sandwich in $\bar{S}$} which would block $z$ and this is forbidden by Claim 17.
+Furthermore, $e\neq t$, since we are assuming $abdt\subseteq abtdt$ is not in $\bar{S}$}.
+So, we are left to consider the case when $n\geq 1$, $bI(T)at_{n}$ is in $R$, and $e=t_{i}$ for some $1\leq i\leq n$.
+Then
+$X\supseteq abtdt_{i}dbI(T)at_{n}\supseteq btdt_{i}btt_{i}\cong dabtdat\cong I_{1,0},$
+a contradiction.
+Case 2: A sandwich of the form edeede blocks $z$.
+Since $d\notin{\bar{R}}$, we must have $e\in{\bar{S}}$.
+If $ede\subseteq{\bar{S}}$, then a sandwich of the form dedded blocks $z$, so we are in the first case.
+Therefore, we can assume that $ez\subseteq{\bar{R}}$.
+By Claim 17, the color $z$ is not blocked by a ytyyty sandwich.
+This, with the fact that $yt\subseteq abt\subseteq S\subseteq{\bar{S}}$ and $ez\subseteq{\bar{R}}$, yields $e\neq y$.
+Since edeede blocks $z$, we have $e\neq z$.
+Therefore, $e\notin\{y,z\}=\{a,b\}$.
+Case 2.1: $P\in\{atb,abt,bat=bI(t_{0})at_{0}\}$.
+Recall that Claim 17 implies that a ytyyty sandwich does not block $z$ in $X$.
+Since $yt\subseteq abt\subseteq S$, $z$ cannot follow the $y$ in $P$.
+This, with the fact that $ez\subseteq{\bar{R}}$, implies that we have $eP\subseteq{\bar{R}}$ when $P$ begins with $zy$.
+So, we have $eP\subseteq{\bar{R}}$ when $P\in\{abt,bI(t)at=bat\}$.
+Now suppose $P=atb$ and $eP\nsubseteq{\bar{R}}$.
+Then, because $ea=ez\subseteq{\bar{R}}$ and $z=a$ cannot follow $y=b$ in $P$, we either have $e=t$ and $atab\subseteq{\bar{R}}$ or $e\neq t$ and $aetab\subseteq{\bar{R}}$ or $ateab\subseteq{\bar{R}}$.
+Note that, in all cases, we have $atab\subseteq{\bar{R}}$, so (ii) is satisfied.
+Therefore, we can assume that $eP\subseteq{\bar{R}}$.
+If $e=t$, then $tP\subseteq{\bar{R}}$ and condition (ii) is satisfied.
+If $e\neq t$, then because we have already shown that $e\notin\{a,b\}$ and we are in Case 2.1, we have $e\notin P$, so condition (iii) is satisfied.
+Case 2.2: $n\geq 1$ and $P=bI(T)at_{n}=bt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$.
+In this case, we have $z=b$ and $y=a$.
+The color $b$ cannot follow $t_{1}t$ in $\bar{R}$} because Claim 17 implies that $b$ is not blocked by a $tt_{1}t$ sandwich.
+Therefore, because $P=bt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$ and $eb\subseteq{\bar{R}}$, if we do not have $eP\subseteq{\bar{R}}$, then we must have $bt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\subseteq{\bar{R}}$.
+But then,
+$X\supseteq abtdR\supseteq abtdbt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\supseteq btdt_{1}btt_{1}\cong dabtdat\cong I_{1,0},$
+a contradiction.
+Therefore, we can assume that ${\bar{R}}\supseteq eP=ebt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$.
+Note that if $e=t$, then (ii) holds and if $e\notin P$, then (iii) holds.
+Since we have already shown that $e\notin\{a,b\}$, we can assume $e=t_{i}$ where $1\leq i\leq n$.
+Recall that $t_{i}=e\in{\bar{S}}$, so since Claim 17 implies that no $t_{i}tt_{i}$ sandwich blocks $b$ in $X$, we must have $abtt_{i}\subseteq{\bar{S}}$.
+Therefore,
+$X\supseteq abtt_{i}dt_{i}bt_{1}tI(t_{1},\dotsc t_{n})at_{n}\supseteq btdt_{i}btt_{i}\cong dabtdat\cong I_{1,0},$
+a contradiction. ∎
+Suppose Claim 18(i) holds.
+Note that $at\subseteq P$ or $n\geq 1$ and $I(T)at_{n}\subseteq P$.
+Therefore, if $dabt\subseteq{\bar{S}}$, then either $X\supseteq dabtdat\cong I_{1,0}$ or $X\supseteq dabtdI(T)at_{n}\cong I_{1,n}$, a contradiction.
+Otherwise, one of the orderings in the following table is a subordering of $X$.
+$datb\cong\mathcal{C}_{2,4} adbtdabt$$dabt\cong\mathcal{C}_{2,3}$ $adbtdbI(T)at_{n}\cong I_{2,n}$
+$datb\cong\mathcal{C}_{2,5} abdtdabt$$dabt\supseteq abdtabt\cong I_{1,0}$ $abdtdbI(T)at_{n}\cong I_{3,n}$
+$abtdedatb\cong\mathcal{C}_{2,7}$ $abtdedabt\cong\mathcal{C}_{2,6}$ $abtdedbI(T)at_{n}\cong I_{5,n}$
+If Claim 18(iii) holds, then by swapping the labels of $d$ and $e$, we can see that either $I_{1,n}\preceq X$ or one of the orderings in the table above is a subpattern of $X$.
+Finally, if Claim 18(ii) holds, then one of the four orderings in the following table is a subordering of $X$.
+$abtdatab\cong\mathcal{C}_{2,1}$ $abtdtatb\cong\mathcal{C}_{2,2}$ $abtdtabt\supseteq abdtabt\cong I_{1,0}$ $abtdtbI(T)at_{n}\cong I_{4,n}$
+See 10
+Proof.
+As in the proof of Lemma 9, to prove the lemma, we can assume that every proper subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$.
+For a contradiction, assume that no subpattern of $X$ is in $\Gamma$.
+We may write $X={\bar{S}}d{\bar{R}}$ where $S\subseteq{\bar{S}}$ and $d\notin{\bar{R}}$ and $R\subseteq{\bar{R}}$.
+Since $(at,R)$ is terminal, there exists a sandwich that blocks $t$ in $atR$.
+Therefore, there exists a color that is neither $a$ nor $t$ in $R$.
+Let $r$ be the first color that is neither $a$ nor $t$ in $R$.
+Furthermore, since a color that is neither $a$ nor $t$ must precede $t$ in $R$, we have $rt\subseteq R$.
+Claim 19.
+$ar\subseteq R$ and $at\subseteq R$.
+Proof.
+Let $x\in\{r,t\}$ and suppose, for a contradiction, that $ax\nsubseteq R$.
+Note that then $x$ is one of the first two colors to appear in $tR$.
+Lemma 6 implies that $x$ is unsortable in $(t,R)$, because $x$ is unsortable in $(at,R)$ and $x\neq a$ and $ax\nsubseteq R$.
+Recall that by the hypothesis of the lemma, there exists $y$ that is sortable in $(t,R)$.
+Since $y$ is unsortable in $(at,R)$, Lemma 6 implies that either $y=a$ or $ay\subseteq R$.
+In either case, $y$ must follow the last occurrence of $x$ in $R$ because $ax\nsubseteq R$.
+However, with Lemma 4, the fact that $y$ is sortable in $(t,R)$ while $x$ is unsortable in $(t,R)$, and the fact that $x$ is one of the first two colors to appear in $tR$, imply that $y$ cannot follow the last occurrence of $x$ in $R$, a contradiction.
+∎
+Claim 20.
+There are distinct colors $c$ and $b$ that are distinct from $a$, $r$, and $t$, such that there exists $P\in\{arat,rart,arbrt,arcbct\}$ such that $P\subseteq R$.
+In particular, we have $art\subseteq R$.
+Proof.
+First, suppose that $rat\subseteq R$.
+Since, by Claim 19, we have $ar\subseteq R$, one of aratarat, rartrart, or ratrratr appears in $R$.
+Since $atdratr=atdraI(r)tr\cong I_{1,0}$, we can assume the subordering ratrratr does not appear $R$, so we are done in this case.
+Now, suppose $rat\nsubseteq R$.
+Since $r$ is the first color in $R$ that is not $a$ or $t$, $rat\nsubseteq R$ implies that the first sandwich that blocks $t$ in $atR$ does not include the color $a$.
+So there exist distinct colors $c$ and $b$ that are both neither $a$ nor $t$, such that a sandwich of the form cbccbc blocks $t$ in $atR$.
+Since $r$ is the first color that is not $a$ or $t$ to appear in $R$, we have either $r=c$ and $rbrt\subseteq R$ or $r\notin\{b,c\}$ and $rcbct\subseteq R$.
+By Claim 19, we have $at\subseteq R$.
+This, with the fact that $rat\nsubseteq R$, implies that either $arbrt\subseteq R$ or $arcbct\subseteq R$, where the distinct colors $b$ and $c$ are distinct from $a$, $r$, and $t$.
+∎
+Claim 21.
+For distinct colors $a$, $r$, $t$, $b$, and $c$ and every $P\in\{arat,rart,arbrt,arcbct\}$ the color $a$ is the only sortable color in $atP$ and $a$ is a good sortable color in $atP$.
+Therefore, no sandwich consisting only of the colors that appear in $atP$ blocks $a$ in $X$.
+Proof.
+It is not hard to see that $a$ is the only sortable color in $atP\subseteq{\bar{S}}R$.
+Furthermore, it is not hard to see that $a$ is a good sortable color in $atP\subseteq{\bar{S}}R$.
+Recall that, by assumption, XdX-d is sortable.
+This means that for any sorting sequence that sorts XdX-d, the color $a$ must be sorted first among the colors in $atP$.
+Therefore, there is no sandwich consisting only of the colors that appear in $atP$ that blocks $a$ in $X$.
+∎
+Note that Claims 20 and 21 establish the first statement of the conclusion.
+Claim 22.
+Let $P\in\{arat,rart,arbrt,arcbct\}$ such that $P\subseteq R$ where $a$, $r$, $t$, $b$, and $c$ are distinct colors. One of the following holds.
+  1. (i)
+    datdat or adtadt is a subordering of $\bar{S}$} or $atde\subseteq{\bar{S}}$ for some color $e\notin P$; or
+  2. (ii)
+    taxattaxat, atxatatxat, txaxttxaxt, or taxyxttaxyxt is a subordering of $\bar{R}$} where $a$, $t$, $x$, and $y$ are distinct colors; or
+  3. (iii)
+    $eP\subseteq{\bar{R}}$ for some color $e\notin P$ and eateat, aetaet, or ateate is a subordering of $\bar{S}$}.
+Proof.
+By assumption, there exists a $d$-sandwich that blocks $a$ in $X={\bar{S}}d{\bar{R}}$.
+That is, for some color $e$, we either have a sandwich of the form dedded which blocks $a$ or a sandwich of the form edeede which blocks $a$.
+Therefore, we have the following two cases.
+Case 1: A sandwich of the form dedded blocks $a$ in $X$.
+Note that since $d\notin{\bar{R}}$, we have $de\subseteq{\bar{S}}$.
+So, if neither $dat\subseteq{\bar{S}}$ nor $adt\subseteq{\bar{S}}$, then $atde\subseteq{\bar{S}}$ and $e\neq t$.
+So to show that $X$ satisfies (i), we only need to show that $e\notin P$.
+First note that $e\neq a$, since dedded blocks $a$.
+If $e=r$, then, since Claim 20 implies that $art\subseteq P$, we have $X\supseteq{\bar{S}}dP\supseteq atdrdart\cong abdtdatb\cong C_{2,5}$, a contradiction.
+We get a similar contradiction if $P=arbrt$ or $P=arcbct$ and $e=b$, since in either case $abt\subseteq P$, so ${\bar{S}}dP\supseteq atdbdabt\cong abdtdatb\cong C_{2,5}$, a contradiction.
+If $P=arcbct$ and $e=c$, then $act\subseteq P$, so ${\bar{S}}dP\supseteq atdcdact\cong abdtdatb\cong C_{2,5}$, a contradiction.
+With Claim 20, this completes the proof of this case.
+Case 2: A sandwich of the form edeede blocks $a$ in $X$.
+Since $d\notin R$, we have $e\in{\bar{S}}$.
+Furthermore, if $de\subseteq{\bar{S}}$, then a sandwich of the form dedded blocks $a$ in $X$, and this was handled in Case 1.
+Therefore, we can assume that $ea\subseteq{\bar{R}}$.
+Case 2.1: $e=t$.
+In this case, $ta=ea\subseteq{\bar{R}}$ and, by Claim 21 a sandwich of the form trttrt does not block $a$ in $X={\bar{S}}d{\bar{R}}$.
+The fact that $at\subseteq{\bar{S}}$ then implies that $rta\nsubseteq{\bar{R}}$.
+Therefore, there exists a $t$ in $\bar{R}$} that is before both an $a$ in $\bar{R}$} and the first $r$ in $\bar{R}$}.
+So, if $P=arat$, then we have either $tarat=eP$ or atratatrat in $\bar{R}$}, so (ii) holds.
+Furthermore, if $P=rart$, we have $trart=eP$ in $\bar{R}$}, so (ii) holds.
+If $P=arbrt$, then one option is $tarbrt=eP\subseteq{\bar{R}}$, and (ii) holds in this case.
+By Claim 21, there is no rbrrbr sandwich that blocks $a$, so the only other possibilities when $P=arbrt$ are $atrabrt\subseteq{\bar{R}}$ or $atrbart\subseteq{\bar{R}}$, and, in either case, we have $trart\subseteq{\bar{R}}$, so (ii) holds.
+Finally, assume $P=arcbct$.
+If $tarcbct=eP\subseteq{\bar{R}}$, then $tacbct\subseteq{\bar{R}}$, so (ii) holds.
+Otherwise, because there is no cbccbc sandwich that blocks $a$, at least one of $atracbct\supseteq tacbct$, $atrcabct\supseteq tcact$, or $atrcbact\supseteq tcact$ is a subordering of $\bar{R}$}, so (ii) holds.
+Case 2.2: $e=r$.
+In this case, $r\in{\bar{S}}$ and $ra=ea\subseteq{\bar{R}}$.
+Recall that, by Claim 21, there is no sandwich of the form rtrrtr that blocks $a$ in $X$.
+Therefore, $r$ must follow $at$ in $\bar{S}$}, so $atr\subseteq{\bar{S}}$.
+Because $art\subseteq P$ and $a$ is blocked by a $d$-sandwich, Lemma 9 (with $r$ playing the role of $b$) implies that a subpattern of $X$ is in $\Gamma$, a contradiction.
+Case 2.3: $e\neq r$ and $e\neq t$.
+Recall that we have $e\in{\bar{S}}$ and $ea\subseteq{\bar{R}}$.
+Since $e\in{\bar{S}}$ and $at\subseteq{\bar{S}}$, one of eateat, aetaet, or ateate is in $\bar{S}$}.
+We will first prove that $e\notin P$ and $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
+Assume first that $e\in R$ and recall that $r$ was defined to be the first color in $R$ that is neither $t$ nor $a$.
+Therefore, $rea\subseteq{\bar{R}}$, which implies that a sandwich of the form ereere blocks $a$ in $X\supseteq{\bar{S}}d{\bar{R}}$.
+Note that this, with Claim 21, implies that $e\notin P$.
+Recall that Claim 21 implies that every color in $atP$ is unsortable except $a$.
+If $e$ is blocked by a sandwich consisting of colors only in $P$, then the subordering formed by removing all colors except those in $P$ and $e$ from XdX-d is trivially unsortable.
+This contradicts the assumption that all proper suborderings of $X$ are sortable.
+Now assume that $e\notin R$.
+In this case, $P\subseteq R$ implies that $e\notin P$.
+By the definition of $\bar{R}$}, the fact that $e\in{\bar{R}}$ implies that the last $e$ follows the last $d$ in $X$.
+Since $e\notin R$ and the state after $d$ is sorted is $(S,R)$, the color $e$ was sorted before the color $d$.
+Since $d$ was sorted before any color in $P$, the color $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
+So, in both cases, $e\notin P$ and $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
+In particular, the subordering ataata does not block $e$ in $X$.
+This means that $e$ does not follow $a$ in $\bar{R}$}.
+Note first that if $eP\subseteq{\bar{R}}$, then (iii) holds.
+So we can assume $eP\nsubseteq{\bar{R}}$.
+This, with the fact that $e$ does does not follow $a$ in $\bar{R}$} and $e\in{\bar{R}}$, implies that $P=rart$ and $reart\subseteq{\bar{R}}$.
+Therefore, at least one of eatdrearteatdreart, aetdreartaetdreart, or atedreartatedreart is a subordering of $X={\bar{S}}d{\bar{R}}$.
+This is a contradiction, because $eatdreart\supseteq eatrear\cong dabtdat\cong I_{1,0}$, $aetdreart\supseteq aetreart\cong adbtdatb\cong\mathcal{C}_{2,4}$, and $atedreart\supseteq atereart\cong abdtdatb\cong\mathcal{C}_{2,5}$.
+∎
+When Claim 22(i) holds, one of the orderings in the following table is a subordering of $X$.
+$datdarat\cong\mathcal{C}_{1,1}$ $datdrart\cong\mathcal{C}_{1,7}$ $datdarbrt\cong\mathcal{C}_{1,12}$ $datdarcbct\succeq\mathcal{C}_{1,12}$
+$adtdarat\cong\mathcal{C}_{1,4}$ $adtdrart\cong\mathcal{C}_{1,8}$ $adtdarbrt\cong\mathcal{C}_{1,11}$ $adtdarcbct\succeq\mathcal{C}_{1,11}$
+$atdedarat\cong\mathcal{C}_{1,5}$ $atdedrart\cong\mathcal{C}_{1,9}$ $atdedarbrt\cong\mathcal{C}_{1,13}$ $atdedarcbct\succeq\mathcal{C}_{1,13}$
+If Claim 22(iii) holds, then by swapping $d$ and $e$, we can see that one of orderings in the preceding table is a subpattern of $X$.
+Finally, if Claim 22(ii) holds, then one of the orderings in the following table is a subpattern of $X$.
+$atdtarat\cong\mathcal{C}_{1,2}$ $atdatrat\cong\mathcal{C}_{1,3}$ $atdtrart\cong\mathcal{C}_{1,6}$ $atdtarbrt\cong\mathcal{C}_{1,10}$
+7. Further Directions
+While the process of foot-sorting lends itself naturally to sorting with a stack, one could consider analogs of the sock sorting problem which utilize other stack configurations or other data structures entirely.
+7.1. Sorting with Multiple Stacks
+Defant and Kravitz introduced a notion of foot-sorting with multiple feet in [1], though with more stacks in series, the sorting process becomes increasingly computationally intensive.
+It would be interesting to investigate perhaps the simplest extension of our work, the basis of sock orderings which are foot-sortable with two stacks in series.
+Defant and Kravitz proved in [1] that for any number $n\geq 1$, one may construct a sock ordering which is unsortable with $n$ stacks.
+Using this proof, one can check that the sock ordering of the form $(abcde)^{1440}$ (that is, the block abcdeabcde, where $a,b,c,d$ and $e$ are distinct sock colors, repeated 1440 times) is unsortable with two stacks in series, but it might be challenging to construct a single sock ordering that is minimally unsortable in this setting.
+7.2. Deque Sorting
+Because a deque, or double-ended queue, allows for more access to the socks for addition and removal, one might think about deque-sorting for socks independently of a foot; a sock could be placed either over an outermost sock or within an innermost sock, and the same is possible for sock removal.
+[^1] Using a deque in the context of permutation sorting has been considered previously (e.g.
+see [3], [2], and [4]).
+This modifies the original operations of foot-sorting, using the language of the deque, to be
+  1. ($\text{stack}_{\uparrow}$)
+    Take the leftmost sock on the right and place it on top of the deque (potentially over any socks that are already in the deque).
+  2. ($\text{stack}_{\downarrow}$)
+    Take the leftmost sock on the right and place it on the bottom of the deque (potentially under any socks that are already in the deque).
+  3. ($\text{unstack}_{\uparrow}$)
+    Remove the topmost sock from the deque and place it so that it becomes the rightmost sock on the left.
+  4. ($\text{unstack}_{\downarrow}$)
+    Remove the bottommost sock from the deque and place it so that it becomes the rightmost sock on the left.
+While the addition of possible moves in this scenario makes for more difficult computation, some similar ideas hold as in the stack setting: in particular, it is now necessary to avoid the pattern abababab from appearing in the deque, just as it is necessary to avoid abaaba appearing in the stack.
+It would be interesting to enumerate the basis of deque-sortable sock orderings, just as we have enumerated the basis of foot-sortable sock orderings.
+It is straightforward to show that the basis of deque-sortable sock orderings is different from $\Gamma$, as well as to show that this basis is infinite.
+Very closely mirroring the proof of Observation 1 under the relaxed restrictions of the deque, one can generate the class of sock orderings of the form $dabtI_{1,n}$ for any $n\geq 0$ and verify that such sock orderings are minimally unsortable with the deque.
+However, with the increased computational difficulty in the deque setting, we do not know whether other minimal deque-unsortable sock orderings may be generated from $\Gamma$ in a similar fashion.
+8. Acknowledgement
+We would like to thank the anonymous referees for their careful reading of the manuscript and for their valuable comments.
+Appendix A Algorithm
+In this appendix, we present a recursive algorithm for foot-sorting that corresponds to our inductive proof of Theorem 2.
+We present this algorithm only for expository purposes, and we made no attempt to optimize the running time of this algorithm.
 (Note that, as referenced in the introduction, independent of our work, Yu has already described a fast algorithm [6].)
-The algorithm first attempts to sort using the following simple greedy algorithm.
-This greedy algorithm is the only part of this section that is referenced in the other sections of this paper.
+The algorithm first attempts to sort using the greedy algorithm.
 Greedy algorithm - Input: A sock ordering $X$.
@@
   2. (H.2)
-    Otherwise, we recursively run the main algorithm on the smaller ordering $X-d$.
+    Otherwise, we recursively run the main algorithm on the smaller ordering XdX-d.
   3. (H.3)
-    If it fails, then by our recursive assumption $X-d$ is unsortable, so $X$ is unsortable.
+    If it fails, then by our recursive assumption XdX-d is unsortable, so $X$ is unsortable.
   4. (H.4)
-    Otherwise, $X-d$ is sortable and we have the following two possibilities.
+    Otherwise, XdX-d is sortable and we have the following two possibilities.
     1. (H.4.1)
-      When we attempt to sort $X$ using the same sequence of colors that fully sorted $X-d$, there is a stage in which the color $d$ only appears as the top color of the stack, which means that we can fully sort $X$ by sorting $d$ at this stage and then continuing with the sequence that was used to sort $X-d$.
+      When we attempt to sort $X$ using the same sequence of colors that fully sorted XdX-d, there is a stage in which the color $d$ only appears as the top color of the stack, which means that we can fully sort $X$ by sorting $d$ at this stage and then continuing with the sequence that was used to sort XdX-d.
     2. (H.4.2)
       $X$ is unsortable because $I_{1,n}\in\Gamma$ is a subpattern of $X$.
-We now present the main algorithm. It closely follows the proof Theorem 3 in Section 2.5, so the proof of Theorem 3 also serves as a proof that this algorithm is correct.
+We now present the main algorithm. It closely follows the proof Theorem 2.
 Main Algorithm - Input: A sock ordering $X$.
@@
 Lemma 9 and Claim 16 in the proof of Theorem 2).
-2.4. Observations and Lemmas
-In the enumeration of our basis $\Gamma$, the five infinite classes have a common structure. In some sense, the following lemma explains the origin of this commonality.
-Lemma 4.
-Let $X$ be a sock ordering that, for some $n\geq 0$, has distinct colors $b$, and $c$ where the final $b$ precedes the final $c$ in $X$.
-Define $A$ and $C$ to be the strings such that $X=AbC$ where $b\notin C$ and $c\in C$.
-Let $\theta$ be a nonempty sorting sequence for $X$ that sorts a color in $C$ and suppose $c$ is the color in $C$ that is sorted first by $\theta$.
-Taking $(S,R):=\theta(X)$, suppose $b\in S$ and that once $c$ is sorted, $b$ never appears as the top color on the stack.
-Suppose $t_{0},\dots,t_{n}$ (in this order) are the $n+1$ colors which appear immediately above the color $b$ in the stack after $c$ is sorted.
-Then
-$I(c,t_{0},\dotsc,t_{n})R\subseteq C.$
-Furthermore, the colors $t_{0},\dotsc,t_{n}$ are distinct and disjoint from $b$ and $c$.
-Proof.
-We will prove the first statement by induction on $n$.
-For the base case, assume $n=0$, so $t_{0}=t_{n}$ is the only color that appears directly above $b$ after $c$ is sorted.
-Since $c$ follows the last $b$ and $b$ can never appear as the top color on the stack after $c$ is sorted, the color $t_{0}=t_{n}$ must appear above $b$ after $c$ is sorted.
-Therefore, we have
-$C\supseteq t_{0}cR=I(c,t_{0},\dotsc,t_{n})R.$
-Now, suppose $n\geq 1$.
-Let $d$ be the color that is sorted that placed the color $t_{n}$ directly above the color $b$ on the stack.
-And let $\theta_{d}$ be the initial sequence of $\theta$ up to but not including $d$.
-Let $(S^{\prime},R^{\prime}):=\theta_{d}(X)$.
-Note, during the execution of $\theta_{d}$, the color $b$ is never the top color on the stack and that $t_{0},\dotsc,t_{n-1}$ is the sequence of colors that appear immediately above the color $b$ after $c$ is sorted.
-Therefore, by induction, we have
-$I(c,t_{0},\dotsc,t_{n-1})R^{\prime}\subseteq C.$
-Recall that $t_{n-1}$ appears immediately above $b$ on the stack $S^{\prime}$, but after $d$ is sorted $t_{n}$ is the color above $b$ on the stack.
-Since $t_{n-1}\neq t_{n}$, this implies that $d=t_{n-1}$.
-Furthermore, we also have $t_{n}t_{n-1}R\subseteq R^{\prime}$.
-This completes the proof because
-$C\supseteq I(c,t_{0},\dotsc,t_{n-1})R^{\prime}\supseteq I(c,t_{0},\dotsc,t_{n-1})t_{n}t_{n-1}R=I(c,t_{0},\dotsc,t_{n-1},t_{n})R.\qed$
-For the final statement, note that, for every $i\in\{0,\dotsc,n\}$, the color $t_{i}\neq b$ by definition and $t_{i}\neq c$ since it appears in the stack after $c$ is sorted.
-Furthermore, the colors $t_{0},\dotsc,t_{n}$ must be $n+1$ distinct colors because the only way for a color to removed from the stack is for it to be sorted.
-We now present a few simple observations and lemmas that will be useful throughout.
-Lemma 5.
-Let $X$ be a sock ordering.
-Suppose that $x\in X$ is unsortable and that $x$ is one of the first two colors to appears in $X$.
-Then any color that appears after the final $x$ in $X$ is blocked by an $x$-sandwich.
-Proof.
-Let $yzy$ be a sandwich that blocks $x$ in $X$.
-Since $x$ is one of the first two colors to appear, $z$ cannot precede the first occurrence of $x$.
-Therefore, both $y$ and $z$ appear between the first and last occurrence of $x$.
-Therefore, any color that follows the final $x$ in $X$ is blocked by one of the sandwiches $xyx$ or $xzx$.
-∎
-There is a natural ordering of the sandwiches that appear in a given sock ordering.
-If $aba$ and $xyx$ are suborderings of a sock ordering $X$ we say that $aba$ precedes $xyx$ if the final $a$ in $aba$ precedes the final $x$ in $xyx$; or when $a=x$ and the final socks in $aba$ and $xyx$ are the same, the $b$ in $aba$ precedes the $y$ in $xyx$; or when $a=x$ and $b=y$ and the last two sock in $aba$ and $xyx$ are the same, the first $a$ in $aba$ precedes the first in $xyx$.
-Observation 6.
-If $aba$ is the first sandwich to appear in a sock ordering $X$, then any color that is not $a$ or $b$ is either sortable in $X$ or is blocked by the sandwich $aba$.
-Proof.
-Let $x$ be an unsortable color in $X$ that is neither $a$ nor $b$.
-Because $x$ is unsortable, it must follow a sandwich.
-Since $aba$ is the first sandwich to appear, $x$ must follow $aba$.
-Therefore, $x$ is blocked by $aba$.
-∎
-Lemma 7.
-Let $S$ and $R$ be strings where $S$ is sorted and has at least two distinct colors.
-Let $a\in S$ where $a$ is not the top color in $S$ and let $S^{\prime}$ be formed by removing the color $a$ from $S$.
-If $x$ is unsortable in $(S,R)$ and $x$ is sortable in $(S^{\prime},R)$, then
-  • $x=a$, or
-  • $ax\subseteq R$, or
-  • there exists a color $y$ distinct from the colors $a$ and $x$ such that $yx\subseteq R$ and $y$ is the top color, $a$ is immediately below the color $y$, and $x$ is immediately below the color $a$ on the stack $S$.
-Proof.
-Assume $x\neq a$.
-Because $x$ is sortable in $(S^{\prime},R)$, we either have $x\in R$, or that $x$ is the top color on $S^{\prime}$.
-Suppose $x$ is the top color on $S^{\prime}$.
-Because $a$ is not the top color on $S$, this implies that $x$ is the top color on $S$.
-Since $x$ is not sortable in $(S,R)$, we have $x\in R$.
-So, in all cases, we have $x\in R$.
-Let $yzy$ be the first sandwich in $SR$ that blocks $x$.
-Since $S$ is sorted, $yzy$ is not a subordering of $S$, so we have $yx\subseteq R$.
-Because $x$ is sortable in $S^{\prime}R$, the color $a$ is either $y$ or $z$.
-If $a$ is $y$, then $ax\subseteq R$, so assume $a$ is $z$.
-Since $yzy=yay$ does not block $x$ in $S^{\prime}R$, we must have $ya\subseteq S$.
-Recall that, by assumption, $a$ is not the top color in $S$, so let $w$ be the color that appears directly above $a$ in $S$.
-If $w\neq x$, then $ywy$ blocks $x$ in $S^{\prime}R$, so we have $w=x$.
-Since $x$ is sortable in $(S^{\prime},R)$ and $x$ is on $S^{\prime}$, the color $x$ must be the top color on the stack $S^{\prime}$.
-Therefore, $x$ is the top color on $S$.
-We now have that the colors $a$ and $x$ are the top two colors on $S$, respectively, and $yx\subseteq R$.
-Therefore, we only need to show that the color $y$ is directly below the color $a$ in $S$.
-This follows from the observation that if there exists a color $u$ that is distinct from $y$, $a$, or $x$ that is between $y$ and $a$ in $S$, then the sandwich $yuy$ blocks $x$ in $S^{\prime}R$, a contradiction.
-∎
-The following lemma is a crucial piece of our inductive proof, because if (I) and (IV) do not hold for a sock ordering $X$ which satisfies the hypothesis, then, because (III) implies that $X$ is sortable, we can determine if a sock ordering is sortable and exactly how to sort $X$ by considering a proper subordering.
-Lemma 8.
-Let $X$ be a sock ordering and let $a$ and $d$ be distinct colors in $X$.
-Let $\overline{\theta}$ be a nonempty sorting sequence for $X$ and let $(S,R):=\overline{\theta}(\emptyset,X)$.
-Let $d$ be the last color in $\overline{\theta}$.
-Furthermore, assume there exists a substack $S^{\prime}\subseteq S$ and subordering $R^{\prime}\subseteq R$ such that the color $a$ is beneath the top color on $S^{\prime}$, $a$ appears in $R^{\prime}$, and $a$ is the only good sortable color in $S^{\prime}R^{\prime}$.
-Then, one of the following is true:
-  1. (I)
-    there is a $d$-sandwich that blocks $a$ in $X$;
-  2. (II)
-    $X-d$ is unsortable;
-  3. (III)
-    $X-d$ is sortable and if a sequence $\theta$ fully sorts $X-d$, then when we attempt to sort $X$ with the sequence $\theta$, there is a stage where the color $d$ only appears as the top color on the stack (this condition implies that $X$ is sortable); or
-  4. (IV)
-    for some $n\geq 0$, $I_{1,n}$ is a subpattern of $X$.
-Proof.
-Assume that (I), (II), and (III) do not hold.
-We will show that (IV) must hold.
-Let $x$ be the top color on $S^{\prime}$ and $D\subseteq X$ be such that $X=YDW$ where the first element of $D$ is the first $d$ in $X$ and the final element of $D$ is the final $d$ in $X$.
-Note that the subordering $R$ must follow the last $d$ in $X$, so $R\subseteq W$.
-Similarly, everything in $S$ must precede the last $d$ in $X$, so $S\subseteq YD$.
-Since $R\subseteq W$ and $a\in R^{\prime}\subseteq R$, we have $a\in W$, so the fact there are no $d$-sandwiches that block $a$ in $X$ implies that the only colors that could appear in $D$ are $a$ and $d$.
-In particular, the color $x$ does not appear in $D$, so the fact that $a$ is beneath $x$ on $S^{\prime}$ and $S^{\prime}\subseteq S\subseteq YD$ implies that $ax\subseteq Y$.
-Note that because $d$ is sorted by $\overline{\theta}$ but $a$ and $x$ are not, there is no sandwich that consists only of $a$ and $x$ that blocks $d$ in $X$.
-Therefore, the color $a$ does not appear in $D$ as otherwise the fact that $ax\subseteq Y$ would imply the existence of a sandwich of the form $axa$ before the final $d$ in $X$.
-As such, $d$ is the only color that appears in $D$ and we can assume without loss of generality that the color $d$ only appears once in $X$, so $D=d$.
-By assumption, the statements (II) and (III) are both false, so there exists a sorting sequence $\theta$ that fully sorts $X-d=YW$ such that when we apply $\theta$ to $X=YDW=YdW$ the color $d$ never appears only as the top color on the stack.
-Let $\theta_{a}$ be the list of initial colors in $\theta$ up to, but not including, $a$ (so $\theta_{a}$ is empty if and only if $\theta$ begins with $a$).
-Note that no color in $S^{\prime}R^{\prime}$ could appear before $a$ in $\theta$ because $a$ is the only good sortable color in $S^{\prime}R^{\prime}\subseteq YW$ and $\theta$ fully sorts $YW$.
-We claim that this implies that no color that appears after $ax$ in $YW$ is in $\theta_{a}$.
-To see this, note that sorting such a color would put $ax$ onto the stack, but that would make it impossible to sort $a$ before sorting $x$.
-In particular, because $ax\subseteq Y$, this means that no color that appears in $W$ is in $\theta_{a}$.
-Therefore, $a$ is the color in $W$ that is sorted first by $\theta$, and $\theta_{a}$ is a sorting sequence for $X=YdW$.
-Furthermore, if we let $(S_{1},R^{\prime}_{1})=\theta_{a}(YW)$, then $(S_{1},R_{1})=\theta_{a}(YdW)$ where $R^{\prime}_{1}=R_{1}-d$.
-Therefore, because $a$ is sortable in $(S_{1},R^{\prime}_{1})$ and no $d$-sandwich blocks $a$ in $X$, the color $a$ is also sortable in $(S_{1},R_{1})$.
-Recall that no color that appears after $ax$ is in $\theta_{a}$ and $x$ is not in $\theta_{a}$.
-This, with the fact that $axda\subseteq SdR\subseteq X$, implies that $xda\subseteq R_{1}$.
-Therefore, sorting $a$ in the state $(S_{1},R_{1})$ pushes $xd$ onto the stack.
-Let $y$ be the color that immediately precedes $d$ in $R_{1}$ (so we could have $y=x$).
-Note that since $ax\subseteq Y$ and either $x=y$ or $xy\subseteq Y$, in all cases, we have
-$ay\subseteq Y.$ (1)
-Recall that, by the previous arguments, $d$ is pushed on the stack when we attempt to apply $\theta$ to $X$, so $\theta$ cannot be a sorting sequence for $X$ as otherwise we would have $(d,\emptyset)=\theta(X)$, a contradiction to the fact that the color $d$ never appears as the top color on the stack.
-Therefore, there must exist a color in $\theta$ that causes the stack to be unsorted when $\theta$ is applied to $X$.
-Let $z$ be the first such color and let $\theta_{z}$ be the list of initial colors in $\theta$ up to but not including $z$.
-That is, $\theta_{z}$ is the longest initial sequence of $\theta$ that is a sorting sequence for $X$.
-Let $(S_{2},R_{2}):=\theta_{z}(X)$.
-Note that by our previous arguments, $a$ must appear in $\theta_{z}$ and $a$ is the color in $W$ that is sorted first by $\theta_{z}$ and $d\notin W$.
-By Lemma 4, we have
-$W\supseteq I(a,t_{0},\dotsc,t_{n})R_{2}=t_{0}aI(t_{0},\dotsc,t_{n})R_{2},$ (2)
-where $t_{0},\dotsc,t_{n}$ is the sequence of colors that appear above $d$ on the stack after $a$ is sorted.
-Note that since $y$ is immediately below $d$ on the stack after the sequence $\theta_{a}$ is applied to $X$ and $d$ is never removed from the stack while $\theta_{z}$ is applied to $X$, the color $y$ is immediately below $d$ on the stack $S_{2}$.
-Let $S_{2}^{\prime}:=S_{2}-d$ and note that $(S_{2}^{\prime},R_{2})=\theta_{z}(X-d)$.
-By the selection of $z$, the color $z$ is unsortable in $(S_{2},R_{2})$, but $z$ is sortable in $(S^{\prime}_{2},R_{2})$.
-Since $d\notin S^{\prime}_{2}R_{2}\subseteq X-d$, we have $z\neq d$ and $dz\nsubseteq R_{2}$, so Lemma 7 implies that $y$, $d$, and $z$ are the top three colors in the stack $S_{2}$ and that $yz\subseteq R_{2}$.
-Note that this means that $z=t_{n}$, since $t_{n}$ is defined to be the last color that appears immediately above $d$ on the stack when we sort $X$ with $\theta_{z}$.
-So, $yt_{n}\subseteq R_{2}$.
-This, with (1) and (2), implies that
-$X=YdW\supseteq aydt_{0}aI(t_{0},\dotsc,t_{n})R_{2}\supseteq aydIt_{0}a(t_{0},\dotsc,t_{n})yt_{n}\cong I_{1,n}.\qed$
-The following three lemmas are needed in the proof of our main theorem to show that a given unsortable sock ordering has a subpattern that is in $\Gamma$.
-Unsurprisingly, these proofs have a number of cases and are somewhat technical, so we defer their proofs until Section 2.6.
-Lemma 9.
-If $X$ is trivially unsortable, then a subpattern of $X$ is in $\Gamma$.
-Lemma 10.
-Let $X$ be a sock ordering where, for some $n\geq 0$, the distinct colors $a,b,d,t_{0},\dotsc,t_{n}$ appear in $X$ and define $t:=t_{0}$.
-Suppose that a nonempty sorting sequence produces $(S,R)$ when it is applied to $X$ and let $d$ be the last color sorted.
-If $abt\subseteq S$ and either
-  • $atb\subseteq R$ or $abt\subseteq R$ and $a$ is blocked by a $d$-sandwich in $X$; or
-  • $bI(t_{0},\dotsc,t_{n})at_{n}\subseteq R$ and $b$ is blocked by a $d$-sandwich in $X$,
-then $X$ is unsortable. If, in addition, every proper subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
-Lemma 11.
-Let $X$ be a sock ordering and suppose that a nonempty sorting sequence produces $(S,R)$ when it is applied to $X$ where $d$ is the last color sorted by the sequence.
-Further suppose that the following two conditions hold:
-  • For every $x\in S$, the state $(x,R)$ is not terminal.
-  • There exist distinct colors $a$ and $t$ in $S$ such that $(at,R)$ is terminal.
-Then there exists $P\subseteq R$ such that $a$ appears in $P$ and $a$ is the only good sortable color in $atP$.
-Furthermore, if $a$ is blocked by a $d$-sandwich in $X$, then $X$ is unsortable.
-If, in addition, every proper subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
-The following lemma shows that if we reach the state $(S,R)$ that is terminal and $S$ has at least three colors, then, in a certain sense, we only need to focus on at most three colors in $S$.
-Furthermore, two of the three colors can be the top two colors on $S$.
-Lemma 12.
-Let $S$ and $R$ be sock orderings where $S$ is sorted and at least three distinct colors appear in $S$.
-Suppose $(S,R)$ is terminal.
-If $b$ and $t$ are the top two colors on $S$, then there exists a color $a\in S$ distinct from $b$ and $t$ such that $(abt,S)$ is terminal.
-Proof.
-Suppose $(S,R)$ is a counterexample.
-We can assume without loss of generality that no colors are repeated in $S$.
-So we can write $S=x_{1}x_{2}\dotsm x_{m}bt$ where $x_{1},\dotsc,x_{m}$, $b$, and $t$ are distinct colors.
-Let $p$ be as small as possible such that there exists $y_{1},\dotsc,y_{p}$, a subsequence of $x_{1},\dotsc,x_{m}$, such that $(y_{1}y_{2}\dotsm y_{p}bt,R)$ is terminal.
-Since $(S,R)$ is a counterexample, we have $p\geq 2$.
-If $p=2$, let $W=b$ and otherwise let $W=y_{3}\dotsm y_{p}b$.
-By the minimality of $p$, for $i\in\{1,2\}$, there exists some $z_{i}$ that is sortable in $(y_{i}Wt,R)$.
-For $i\in\{1,2\}$, since $z_{i}$ is sortable in $(y_{i}Wt,R)$, we have $z_{i}\notin y_{i}W$.
-In particular, this implies that $z_{i}\neq y_{i}$ and $z_{i}\neq b$.
-Furthermore, since $z_{i}$ is sortable in $(y_{i}Wt,R)$, there is no $y_{i}by_{i}$ sandwich blocking $z_{i}$ in $y_{i}WtR$, so $y_{i}z_{i}\nsubseteq R$.
-We will now show that $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
-Recall that $(y_{1}y_{2}Wt,R)=(y_{1}y_{2}\dotsm y_{p}bt,R)$ is terminal, so $z_{1}$ is not sortable in $(y_{1}y_{2}Wt,R)$.
-Therefore, we can apply Lemma 7 to deduce that $z_{1}$ is unsortable in $(y_{2}Wt,R)$, because $z_{1}\neq y_{1}$, $y_{1}z_{1}\nsubseteq R$, and $y_{2}$ is not the second color on the stack $y_{1}y_{2}Wt$.
-Since $z_{1}$ is unsortable in $(y_{2}Wt,R)$, $y_{1}z_{1}\nsubseteq R$, and $z_{1}\notin W$, either $z_{1}=y_{2}$ or there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ in which $y_{1}$ does not appear.
-If $z_{1}=y_{2}$, then, because $z_{1}=y_{2}$ is sortable in $(y_{1}Wt,R)$, we have $y_{2}\in R$, so $y_{1}y_{2}=y_{1}z_{1}\nsubseteq R$ implies that either $y_{1}\notin R$ or that $y_{1}$ is blocked by a sandwich of the form $y_{2}by_{2}$ in $y_{2}WtR$.
-So, $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
-If there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ in which $y_{1}$ does not appear, then $y_{1}z_{1}\nsubseteq R$ implies that $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
-So, in all cases, $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
-Since $y_{1}$ is not sortable in $(y_{2}Wt,R)$, we have $y_{1}\neq z_{2}$.
-This, with the fact that $z_{2}$ is not sortable in $(y_{1}y_{2}Wt,R)$ but is sortable $(y_{2}Wt,R)$ together with Lemma 7, yields $y_{1}z_{2}\subseteq R$.
-Since $y_{1}z_{2}\subseteq R$ and $y_{2}z_{2}\nsubseteq R$, $y_{2}$ is not sortable in $(y_{1}Wt,R)$, so $y_{2}\neq z_{1}$.
-Using this fact with Lemma 7, allows us to conclude that $y_{2}z_{1}\subseteq R$ because $z_{1}$ is not sortable in $(y_{1}y_{2}Wt,R)$ but is sortable $(y_{1}Wt,R)$.
-Since $y_{1}z_{2}\subseteq R$, $y_{2}z_{2}\nsubseteq R$, and $y_{2}z_{1}\subseteq R$, we have $y_{1}z_{1}\subseteq R$.
-But we have already established that $y_{1}z_{1}\nsubseteq R$.
-∎
-Our last lemma is simple despite the somewhat technical statement.
-Recall the greedy algorithm described in Section 2.3.
-This lemma is used in the case when a color, say $x$, is not sorted by the greedy algorithm at a particular stage despite terminating before another color, say $y$, which is sorted by the greedy algorithm at that stage.
-Note that, in this case, if a sandwich blocks $x$, then it must be a $y$-sandwich because $y$ is sortable.
-The lemma deals with the other possibility.
-Namely, $x$ already being beneath the top color on the stack.
-Lemma 13.
-Let $S$ and $R$ be sock orderings where $S$ is sorted and let $x$ and $y$ be distinct colors in $R$ such that the final $x$ in $R$ appears before the final $y$ in $R$.
-Suppose that, with respect to $(S,R)$, the color $y$ is sortable and the color $x$ is unsortable.
-If no $y$-sandwich blocks $x$ in $SR$, then $x$ and $y$ are the top two colors on $S$.
-Furthermore, no color except possibly $y$ precedes $x$ in $R$.
-Proof.
-Because $y$ is sortable in $SR$, every sandwich in $SR$ that appears before the final $y$ in $R$ must be a $y$-sandwich.
-Therefore, because the final $x$ in $R$ precedes the final $y$ in $R$ and $x$ is not blocked by a $y$-sandwich in $SR$, the color $x$ is not blocked by a sandwich in $SR$.
-The fact that $x$ is unsortable then implies that $x$ is below the top color on the stack $S$.
-Since $y$ is sortable in $SR$, there is no $x$-sandwich that blocks $y$ in $SR$.
-Therefore, the colors $x$ and $y$ are the top two colors on $S$.
-Furthermore, no color except possibly $y$ can precede $x$ in $R$.
-∎
-2.5. Proof of Theorem 3
-The proof is by induction on the number of socks in $X$.
-The base case is vacuously true.
-So assume that $X$ is an unsortable sock ordering.
-With the induction hypothesis, we can assume that there does not exist a proper subordering of $X$ that is unsortable.
-Note that if we have colors $a$ and $d$ and substrings $S^{\prime}\subseteq S\subseteq X$ and $R^{\prime}\subseteq R\subseteq X$ that satisfy the hypothesis of Lemma 8, then we can assume Lemma 8(I) holds.
-That is, we can assume that $a$ is blocked by a $d$-sandwich.
-To see this, note first that Lemma 8(II) cannot hold because every proper subordering of $X$ is sortable.
-Additionally, Lemma 8(III) cannot hold because $X$ is unsortable.
-Furthermore, if Lemma 8(IV) holds, then we satisfy the conclusion.
-This will make it possible to apply either Lemma 10 or 11 to complete the proof.
-Let $(S,R)$ be output of the greedy algorithm (as described in Section 2.3) applied to $X$. This implies that $(S,R)$ is terminal. Since $X$ is unsortable, we have $R\neq\emptyset$.
-Case 1: $S=\emptyset$ or there exists $x\in S$ such that $(x,R)$ is terminal.
-If $S=\emptyset$, then $R$ is trivially unsortable, so Lemma 9 implies that a subpattern of $R\subseteq X$ is in $\Gamma$.
-Similarly, if there exists $x\in S$ such that $(x,R)$ is terminal, then $xR$ is trivially unsortable.
-Lemma 9 then implies that a subpattern of $xR\subseteq X$ is in $\Gamma$.
-From now on, we will assume that we are not in Case 1.
-That is, for every $x\in S$, some color in $xR$ is sortable.
-Since $(S,R)$ is terminal, this implies that there are at least two colors on $S$.
-Define $t$ to be the top color in $S$, and let $b$ be the color in $S$ appearing immediately under $t$.
-We can also now let $d$ be the last color sorted by the greedy algorithm.
-Case 2: $(bt,R)$ is terminal.
-By Lemma 11, there exists $P\subseteq R$ such that $b$ appears in $P$ and such that $b$ is the only good sortable color in $btP$.
-This means that the hypothesis of Lemma 8 is satisfied with $b$, $d$, $bt$, and $P$ playing the roles of $a$, $d$, $S^{\prime}$, and $R^{\prime}$, respectively.
-Therefore, we can assume that $b$ is blocked by a $d$-sandwich in $X$.
-But then Lemma 11 implies that a subpattern of $X$ is in $\Gamma$.
-Case 3: $(bt,R)$ is not terminal. With Lemma 12, we can define $a$ to be the lowest color on the stack $S$ such that $(abt,R)$ is terminal.
-Claim 14.
-$at\subseteq R$.
-Proof.
-Suppose, by way of contradiction, that $at\nsubseteq R$.
-Then, because $t$ is unsortable in $(abt,R)$ and $t\neq a$ and $a$ is not the second color on the stack $abt$, Lemma 7 implies that $t$ is unsortable in $(bt,R)$.
-By the case, $(bt,R)$ is not terminal, so since both $b$ and $t$ are not sortable in $(bt,R)$, there must exist a color $y\in R$ that is not $b$ or $t$ which is sortable in $(bt,R)$.
-If $y=a$, then, because $at\nsubseteq R$, the color $y=a$ must follow the last occurrence of $t$ in $btR$.
-If $y\neq a$, then, since $y$ is sortable in $(bt,R)$ and unsortable in $(abt,R)$, Lemma 7 implies that $ay\subseteq R$.
-Because $at\nsubseteq R$, this implies that $y$ follows the last occurrence of $t$ in $btR$.
-So, in both cases, $y$ follows the last occurrence of $t$ in $btR$.
-This with Lemma 5 and the fact that $t$ is the second color in $btR$ and $t$ is unsortable in $btR$ give us that $y$ is unsortable in $btR$, a contradiction.
-∎
-Case 3.1: There exists a color $z\in S$ distinct from $a$ and $t$ such that $azt\subseteq S$ and $z\in R$.
-By Claim 14, we can assume that either $atz$, $azt$, or $zat=zaI(t)t$ is in $R$.
-But then Lemma 8 and Lemma 10 give us the desired conclusion, because $a$ is the only good sortable color in both $aztatz$ and $aztazt$, and $z$ is the only good sortable color in $aztzat$.
-Case 3.2: For every $z\in S$ distinct from $a$ and $t$ such that $azt\subseteq S$, we have $z\notin R$.
-Recall that $abt\subseteq S$.
-This and the case, implies that $b\notin R$.
-Let $A$ and $C$ be such that $X=AbC$ and $b\notin C$.
-Because $b\in S$ and $b\notin R$, there exists a color in $C$ that is sorted by the greedy algorithm.
-Let $c\in C$ be the color in $C$ that is sorted first by the greedy algorithm and let $(S^{\prime},A^{\prime}bC)$ be the state immediately before $c$ is sorted.
-Note that we have
-$S^{\prime}A^{\prime}\subseteq A.$ (3)
-Recall that the color $b$ is never sorted by the greedy algorithm and note that, after $c$ is sorted, the color $b$ only appears on the stack. Therefore, because $ab\subseteq S$, we have
-$a\in S^{\prime}A^{\prime}$ (4)
-Furthermore, since $b$ is not sorted by the greedy algorithm, the color $b$ can never appear as the top color on the stack after $c$ is sorted by the greedy algorithm.
-For $n\geq 0$, let $t_{0},\dotsc,t_{n}$ be the $n+1$ distinct colors that appear immediately above $b$ after $c$ is sorted by the greedy algorithm.
-Note that then $t=t_{n}$, so Claim 14 implies that $at_{n}=at\subseteq R$.
-With Lemma 4, we have
-$C\supseteq I(c,t_{0},\dotsc,t_{n})R\supseteq I(c,t_{0},\dotsc,t_{n})at_{n}.$ (5)
-Note that if $ca\subseteq A$, then, with (5), we have
-$X=AbC\supseteq cabI(c,t_{0},\dotsc,t_{n})at_{n}=cabt_{0}cI(t_{0},\dotsc,t_{n})at_{n}\cong I_{1,n}.$
-So, for the remainder of the proof, we will assume
-$ca\nsubseteq A$ (6)
-Claim 15.
-There are at least three colors in $S^{\prime}$.
-Furthermore, if $y$ is the color immediately beneath the top color in $S^{\prime}$, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC$.
-Proof.
-First assume that $b$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$.
-Then, because the greedy algorithm sorted $c$ and not $b$ at the stage $(S^{\prime},A^{\prime}bC)$ and the final $b$ occurs before the final $c$ in $S^{\prime}A^{\prime}bC$, with respect to $(S^{\prime},A^{\prime}bC)$, the color $b$ is unsortable and the color $c$ is sortable.
-Then, Lemma 13 implies that $b$ and $c$ are the top two colors on the stack $S^{\prime}$, so $y=b$.
-Since (3), (4), and (6), all hold and $b$ and $c$ are the top two colors on the stack $S^{\prime}$, we must have $ayc=abc\subseteq S^{\prime}$.
-By (5), we also have
-$yI(c,t_{0},\dotsc,t_{n})at_{n}=bI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC.$
-Now assume that $b$ is blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$.
-Let $\hat{y}$ be the other color in the first $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$.
-We will show that $\hat{y}=y$ and our conclusion holds.
-To see this, first note that we must have $\hat{y}c\subseteq S^{\prime}A^{\prime}\subseteq A$ and $c\hat{y}\subseteq S^{\prime}A^{\prime}\subseteq A$.
-Furthermore, (4) and (6) and $c\hat{y}\subseteq S^{\prime}A^{\prime}$ imply that $\hat{y}\neq a$ and
-$a\hat{y}\subseteq S^{\prime}A^{\prime}.$ (7)
-If the final $\hat{y}$ follows the final $c$ in $S^{\prime}A^{\prime}bC$, then, after $c$ is sorted, the color $\hat{y}$ must appear off the stack and, with (7), we will also have $a\hat{y}b$ on the stack.
-Then, the fact that $b$ is never sorted by the greedy algorithm implies that $\hat{y}$ is never sorted by the greedy algorithm.
-Furthermore, since we cannot have a $\hat{y}b\hat{y}$-sandwich on the stack, we must have $a\hat{y}bt_{n}\subseteq S$ and $\hat{y}\in R$.
-But this contradicts the fact that we are in Case 3.2.
-Therefore, the final $\hat{y}$ precedes the final $c$ in $S^{\prime}A^{\prime}bC$.
-Recall that $\hat{y}$ is the color other than $c$ that is in the first $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$.
-Since $b\notin C$, we have $\hat{y}b\subseteq S^{\prime}A^{\prime}b$.
-Because there is no sandwich consisting of $b$ and $\hat{y}$ which blocks $c$ in $S^{\prime}A^{\prime}bC$ and both the final $b$ and the final $\hat{y}$ in $S^{\prime}A^{\prime}bC$ preceed the final $c$ in $S^{\prime}A^{\prime}bC$, this implies that $b\hat{y}\nsubseteq S^{\prime}A^{\prime}bC$.
-In particular, we have $\hat{y}\notin C$.
-We claim that the color $\hat{y}$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$.
-For a contradiction, assume that $\bar{y}$ is the other color in such a sandwich.
-Since $\hat{y}\notin C$, this sandwich is in $S^{\prime}A^{\prime}b$, and $\bar{y}\hat{y}\subseteq S^{\prime}A^{\prime}b$.
-Since $b\hat{y}\nsubseteq S^{\prime}A^{\prime}bC$, this implies that $\bar{y}\neq b$.
-Therefore, there is a sandwich consisting of $\bar{y}\neq b$ and $c$ in $S^{\prime}A^{\prime}b$, so this sandwich blocks $b$ as well as $\hat{y}$.
-Furthermore, since there is no sandwich consisting of $\bar{y}$ and $\hat{y}$ that blocks $c$ in $S^{\prime}AbC$, we have $\hat{y}\bar{y}\nsubseteq S^{\prime}Ab$.
-But this contradicts the fact that $\hat{y}$ is in the first $c$-sandwich in $S^{\prime}A^{\prime}bC$ that blocks $b$.
-So we have established that the color $\hat{y}$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$.
-Recall that the final $\hat{y}$ precedes the final $c$ in $S^{\prime}A^{\prime}bC$, so Lemma 13 implies that $c$ is on top of the stack $S^{\prime}$ and the color $\hat{y}$ is immediately below the color $c$ on $S^{\prime}$.
-Note that we now have $\hat{y}=y$ since $y$ was defined to be the color below the top color on $S^{\prime}$.
-Since $y=\hat{y}$ is in a $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$ this also implies that $y=\hat{y}\in A^{\prime}$.
-Furthermore, with (5),
-$A^{\prime}bC\supseteq ybI(c,t_{0},\dotsc,t_{n})at_{n}\supseteq yI(c,t_{0},\dotsc,t_{n})at_{n}.$
-Since (3), (4), and (6), all hold and $y$ and $c$ are the top two colors on $S^{\prime}$, we have $ayc\subseteq S^{\prime}$. ∎
-Define $d^{\prime}$ to be the color sorted by the greedy algorithm immediately before reaching the state $(S^{\prime},A^{\prime}bC)$.
-By Claim 15, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC$.
-Note that $y$ and $c$ are the only sortable colors in $aycyI(c,t_{0},\dotsc,t_{n})at_{n}$, because every other color is blocked by a $ycy$ sandwich.
-Furthermore, after sorting $c$ we have $(ayyt_{0},I(t_{0},\dotsc,t_{n})at_{n})$ which Observation 2 implies is unsortable.
-Therefore, with Observation 2, $y$ is the only good sortable color in $aycyI(c,t_{0},\dotsc,t_{n})at_{n}$.
-This means that the hypothesis of Lemma 8 is satisfied with $y$, $d^{\prime}$, $ayc$, and $yI(c,t_{0},\dotsc,t_{n})at_{n}$ playing the roles of $a$, $d$, $S^{\prime}$ and $R^{\prime}$, respectively.
-Therefore, we can assume that $y$ is blocked by a $d^{\prime}$-sandwich in $X$.
-Then Lemma 10 implies that a subpattern of $X$ is in $\Gamma$.
-2.6. Proofs of Lemmas 9, 10, and 11
-See 9
-Proof.
-Assume, for a contradiction, that $X$ is a minimal counterexample.
-Let $aba$ be the first sandwich in $X$.
-Recall that since $X$ is trivially unsortable, every color is blocked by a sandwich.
-That is, for all $z\in X$, there exist distinct colors $x$ and $y$ in $X$ which are distinct from $z$ such that $xyxz\subseteq X$.
-This implies that every color must follow the first sandwich $aba$.
-Let $z$ be the first of the colors $a$ and $b$ that terminates in $X$.
-Let $xyx$ be the first sandwich in $X$ that blocks $z$.
-Case 1: $x=a$.
-Then $xyx=aya$ blocks $z$, so $z=b$.
-Suppose the $y$ in the sandwich $aya$ follows the first sandwich $aba$, then $X\supseteq abayaba\cong abacaba\cong\mathcal{T}_{1}$.
-Otherwise, because $aba$ (and not $aya$) is the first sandwich, we have $X\supseteq abya$.
-This $abya$ subordering is followed by $ba$ and the color $y$ in $X$.
-Therefore, either $X\supseteq abyabya\cong abcabca\cong\mathcal{T}_{3}$, $abyayba\cong abcacba\cong\mathcal{T}_{4}$, or $X\supseteq abyabay\cong abcabac\cong\mathcal{T}_{5}$.
-Case 2: $x=b$.
-Then $xyx=byb$ blocks $z$, so $z=a$.
-Suppose the $y$ in the sandwich $byb$ follows the first sandwich $aba$.
-Then $X\supseteq abaybab\cong abacbab\cong\mathcal{T}_{2}$.
-Otherwise, because $aba$ is the first sandwich (and not $aya$), we have $X\supseteq abya$.
-Since $byb$ blocks $z=a$ and the final $b$ follows the final $z=a$, the subordering $bab$ must follow the subordering $abya$.
-Furthermore, since $y$ must follow the first sandwich, the color $y$ must follow this $abya$.
-So, as in the previous case, $abya$ is followed by $ba$ and the color $y$.
-Therefore, as in the previous case, either $\mathcal{T}_{3}$, $\mathcal{T}_{4}$, or $\mathcal{T}_{5}$ is a subpattern of $X$.
-Case 3: $x\neq a$ and $x\neq b$.
-Case 3.1: The first $x$ appears before the first sandwich $aba$.
-Then, because $aba$ is the first sandwich to appear, the color $x$ must also follow $aba$.
-Therefore, $X\subseteq xabaxab\cong abcbacb\cong\mathcal{T}_{6}$ or $X\subseteq xabaxba\cong abcbabc\cong\mathcal{T}_{7}$.
-Case 3.2: The first $x$ appears between the first and last $a$ in the first sandwich $aba$.
-By the case, an $axa$ sandwich precedes the first $xyx$ sandwich. Since the first sandwich that blocks $z$ is $xyx$, we must have $z=a$. Note that, because $xyx$ blocks $z=a$, we also have $y\neq a$.
-The color $x$ must follow the $ab$ in the first sandwich $aba$, as otherwise an $axa$ sandwich would proceed the first $aba$ sandwich.
-Since $xyx$ is not the first sandwich in $X$ and $y\neq a$, either $X\supseteq abxayx$ or $X\supseteq abxyax$.
-We can assume that both $aba$ and $xyx$ in this subordering are the first such subordering to appear in $X$.
-If $y=b$, then the subordering is $abxabx$ or $abxbax$.
-But then $bxb$ precedes $xbx=xyx$, so $xbx=xyx$ is not the first sandwich which blocks $z=a$.
-Therefore, we can assume $y\neq b$.
-Therefore, the colors $a$ and $b$ are distinct from the colors $x$ and $y$, so either $X\supseteq abxayxab\cong abcadcab\cong\mathcal{T}_{8}$ or the subordering is $abxyax$.
-So, we can assume the subordering is $abxyax$.
-Recall that every color must follow the first $aba$ sandwich, so every color must follow $abxya$.
-This with the fact that $a$ is neither $x$ nor $y$ implies that every color that is not $a$ is blocked by either $axa$ or $aya$.
-Furthermore, $a$ is blocked by $xyx$.
-Therefore, every color is blocked by one of the three sandwiches $axa$, $aya$, or $xyx$.
-Since none of these three sandwiches includes the color $b$, if we remove the color $b$ from $X$ we have a trivially unsortable sock ordering which contradicts the minimality of $X$.
-Case 3.3: The first $x$ appears after the first sandwich $aba$.
-So, we have $abaxyx$ in $X$ were $aba$ and $xyx$ in this subordering are the first such subordering to appear in $X$.
-If $y=a$, then we must have $z=b$, but then $axy=axa$ appears before $xyx$, a contradiction to the fact that $xyx$ is the first sandwich that blocks $z=b$.
-Similarly, if $y=b$, then we must have $z=a$, but then $bxy=bxb$ blocks $z=a$ and appears before $xyx$.
-Therefore, the colors $x$ and $y$ are distinct from the colors $a$ and $b$.
-Hence, $X\supseteq abaxyxba\cong abacdcba\cong\mathcal{T}_{9}$ or $X\supseteq abaxyxab\cong abacdcab\cong\mathcal{T}_{10}$.
-∎
-See 10
-Proof.
-First note that if a proper subordering of $X$ is unsortable, then $X$ is unsortable.
-Therefore, to prove the lemma, we can assume that every proper subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$, since this will imply that $X$ is unsortable.
-Assume for a contradiction that no subpattern of $X$ is in $\Gamma$.
-Let $T:=(t_{0},\dotsc,t_{n})$ and $P\in\{atb,abt,bI(T)at_{n}\}$ such that $P\subseteq R$.
-We may write $X=S^{\prime}dR^{\prime}$ where $S\subseteq S^{\prime}$ and $d\notin R^{\prime}$ and $R\subseteq R^{\prime}$.
-Note that $a$ is the only sortable color in $abtatb$ and $abtabt$.
-We claim that $b$ is the only good sortable color in $abtbI(T)at_{n}$.
-To see this note that when $n=0$ the color $b$ is the only sortable color in $abtbI(T)at_{n}=abtbat$ and when $n\geq 1$, $b$ and $t$ are the only sortable colors in $abtbI(T)at_{n}$, but sorting $t=t_{0}$ leaves $(abbt_{1},I(t_{1},\dotsc,t_{n})at_{n})$ which is unsortable by Observation 2.
-  • If $P\in\{atb,abt\}$, then let $z=a$ and $y=b$.
-  • If $P=bI(T)at_{n}$, then let $z=b$ and $y=a$.
-We have the following.
-Claim 16.
-The color $z$ is blocked by a $d$-sandwich in $X$ and the color $z$ is not blocked by a sandwich consisting only of colors in $P$.
-Proof.
-The first conjunct is given by the hypothesis of the lemma.
-For the second conjunct, recall that, by the above discussion, $z$ is the only good sortable color in $abtP\subseteq X-d$.
-Therefore, because $X-d$ is a proper subordering of $X$ and is therefore sortable, $z$ is not blocked by a sandwich consisting only of colors in $P$.
-∎
-Claim 17.
-One of the following holds.
-  1. (i)
-    one of $dabt$, $adbt$, $abdt$, is a subordering of $S^{\prime}$ or $abtde\subseteq S^{\prime}$ for some color $e\notin P$; or
-  2. (ii)
-    $atab\subseteq R^{\prime}$ or $tP\subseteq R^{\prime}$; or
-  3. (iii)
-    for some color $e\notin P$, $e\in S^{\prime}$ and $eP\subseteq R^{\prime}$.
-Proof.
-Recall that by Claim 16, $z$ is blocked by a $d$-sandwich in $X$.
-That is, for some color $e$, we either have a sandwich of the form $ded$ which blocks $z$ or a sandwich of the form $ede$ which blocks $z$.
-Case 1: A sandwich of the form $ded$ blocks $z$.
-In this case, we will show that (i) holds, so suppose that neither $dabt$, $adbt$, nor $abdt$ is a subordering of $S^{\prime}$.
-Then, because $d\notin R^{\prime}$ implies $de\subseteq S^{\prime}$, we have $abtde\subseteq S^{\prime}$.
-To show that (i) holds, we then only need to show that $e\notin P$.
-Since $ded$ blocks $z$, we have $e\neq z$.
-We also have $e\neq y$, since otherwise we would have a $yty$ sandwich in $S^{\prime}$ which would block $z$ and this is forbidden by Claim 16.
-Furthermore, $e\neq t$, since we are assuming $abdt\subseteq abtdt$ is not in $S^{\prime}$.
-So, we are left to consider the case when $n\geq 1$, $bI(T)at_{n}$ is in $R$, and $e=t_{i}$ for some $1\leq i\leq n$.
-Then
-$X\supseteq abtdt_{i}dbI(T)at_{n}\supseteq btdt_{i}btt_{i}\cong dabtdat\cong I_{1,0},$
-a contradiction.
-Case 2: A sandwich of the form $ede$ blocks $z$.
-Since $d\notin R^{\prime}$, we must have $e\in S^{\prime}$.
-If $ede\subseteq S^{\prime}$, then a sandwich of the form $ded$ blocks $z$, so we are in the first case.
-Therefore, we can assume that $ez\subseteq R^{\prime}$.
-By Claim 16, the color $z$ is not blocked by a $yty$ sandwich.
-This, with the fact that $abt\subseteq S\subseteq S^{\prime}$ and $ez\subseteq R^{\prime}$, yields $e\neq y$.
-Since $ede$ blocks $z$, we have $e\neq z$.
-Therefore, $e\notin\{y,z\}=\{a,b\}$.
-Case 2.1: $P\in\{atb,abt,bat=bI(t_{0})at_{0}\}$.
-Recall that Claim 16 implies that a $yty$ sandwich does not block $z$ in $X$.
-Since $yt\subseteq abt\subseteq S$, $z$ cannot follow the $y$ in $P$.
-This, with the fact that $ez\subseteq R^{\prime}$, implies that we have $eP\subseteq R^{\prime}$ when $P$ begins with $zy$.
-So, we have $eP\subseteq R^{\prime}$ when $P\in\{abt,bI(t)at=bat\}$.
-If $P=atb$ and $eP\nsubseteq R^{\prime}$, then, because $ea=ez\subseteq R^{\prime}$ and $z=a$ cannot follow $y=b$ in $P$, we either have $aetab\subseteq R^{\prime}$ or $ateab\subseteq R^{\prime}$.
-Note that, in both cases, we have $atab\subseteq R^{\prime}$.
-First note that if $e=t$, then the preceding argument implies that condition (ii) is satisfied, so assume $e\neq t$.
-Recall that we have already shown that $e\notin\{a,b\}$, so, since we are in Case 2.1, we have $e\notin P$.
-Recall that $eP\subseteq R^{\prime}$ when $P\in\{abt,bI(t)at=bat\}$.
-Therefore, to show that (iii) is satisfied, we only need to show that $eP\subseteq R^{\prime}$ when $P=atb$.
-For a contradiction, assume $P=atb$ and $eP\nsubseteq R^{\prime}$.
-Recall that we have previously shown that $e\in S^{\prime}$ and that, in this case, we either have $aetab\subseteq R^{\prime}$ or $ateab\subseteq R^{\prime}$.
-Since $e$ and $t$ are both in $S^{\prime}$, a sandwich consisting of the colors $e$ and $t$ blocks both $a$ and $b$ in $X=S^{\prime}dR^{\prime}$.
-Since $abt\subseteq S^{\prime}$, a sandwich of the form $ata$ blocks $e$ in $X=S^{\prime}dR^{\prime}$ and a sandwich of the form $aba$ blocks $t$ in $X=S^{\prime}dR^{\prime}$.
-Therefore, if we consider the subordering $X^{\prime}$ formed by removing all colors that are not $a$, $b$, $t$, or $e$ from $X$, we have that $X^{\prime}$ is a subordering of $X-d$ that is trivially unsortable.
-This contradicts the fact that every proper subordering of $X$ is sortable.
-Case 2.2: $n\geq 1$ and $P=bI(T)at_{n}=bt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$.
-In this case, we have $z=b$ and $y=a$.
-The color $b$ cannot follow $t_{1}t$ in $R^{\prime}$ because Claim 16 implies that $b$ is not blocked by a $tt_{1}t$ sandwich.
-Therefore, because $P=bt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$ and $eb\subseteq R^{\prime}$, if we do not have $eP\subseteq R^{\prime}$, then we must have $bt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\subseteq R^{\prime}$.
-But then,
-$X\supseteq abtdR\supseteq abtdbt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\supseteq btdt_{1}btt_{1}\cong dabtdat\cong I_{1,0},$
-a contradiction.
-Therefore, we can assume that $R^{\prime}\supseteq eP=ebt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$.
-Note that if $e=t$, then (ii) holds and if $e\notin P$, then (iii) holds.
-Since we have already shown that $e\notin\{a,b\}$, we can assume $e=t_{i}$ where $1\leq i\leq n$.
-Recall that $t_{i}=e\in S^{\prime}$, so since Claim 16 implies that no $t_{i}tt_{i}$ sandwich blocks $b$ in $X$, we must have $abtt_{i}\subseteq S^{\prime}$.
-Therefore,
-$X\supseteq abtt_{i}dt_{i}bt_{1}tI(t_{1},\dotsc t_{n})at_{n}\supseteq btdt_{i}btt_{i}\cong dabtdat\cong I_{1,0},$
-a contradiction. ∎
-Suppose Claim 17(i) holds.
-Note that $at\subseteq P$ or $n\geq 1$ and $I(T)at_{n}\subseteq P$.
-Therefore, if $dabt\subseteq S^{\prime}$, then either $X\supseteq dabtdat\cong I_{1,0}$ or $X\supseteq dabtdI(T)at_{n}\cong I_{1,n}$.
-Otherwise, one of the orderings in the following table is a subordering of $X$.
-$adbtdatb\cong\mathcal{C}_{2,4}$ $adbtdabt\cong\mathcal{C}_{2,3}$ $adbtdbI(T)at_{n}\cong I_{2,n}$
-$abdtdatb\cong\mathcal{C}_{2,5}$ $abdtdabt\supseteq abdtabt\cong I_{1,0}$ $abdtdbI(T)at_{n}\cong I_{3,n}$
-$abtdedatb\cong\mathcal{C}_{2,7}$ $abtdedabt\cong\mathcal{C}_{2,6}$ $abtdedbI(T)at_{n}\cong I_{5,n}$
-If Claim 17(iii) holds, then by swapping the labels of $d$ and $e$, we can see that $X$ either $I_{1,n}\preceq X$ or one of the orderings in the table above is a subpattern of $X$.
-Finally, if Claim 17(ii) holds, then one of the four orderings in the following table is a subordering of $X$.
-$abtdatab\cong\mathcal{C}_{2,1}$ $abtdtatb\cong\mathcal{C}_{2,2}$ $abtdtabt\supseteq abdtabt\cong I_{1,0}$ $abtdtbI(T)at_{n}\cong I_{4,n}$
-See 11
-Proof.
-As in the proof of Lemma 10, to prove the lemma, we can assume that every proper subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$.
-For a contradiction, assume that no subpattern of $X$ is in $\Gamma$.
-We may write $X=S^{\prime}dR^{\prime}$ where $S\subseteq S^{\prime}$ and $d\notin R^{\prime}$ and $R\subseteq R^{\prime}$.
-Since $(at,R)$ is terminal, there exists a sandwich that blocks $t$ in $atR$.
-Therefore, there exists a color that is neither $a$ nor $t$ in $R$.
-Let $r$ be the first color that is neither $a$ nor $t$ in $R$.
-Furthermore, since a color that is neither $a$ nor $t$ must precede $t$ in $R$, we have $rt\subseteq R$.
-Claim 18.
-$ar\subseteq R$ and $at\subseteq R$
-Proof.
-Let $x\in\{r,t\}$ and suppose, for a contradiction, that $ax\nsubseteq R$.
-Note that then $x$ is one of the first two colors to appear in $tR$.
-Lemma 7 implies that $x$ is unsortable in $(t,R)$, because $x$ is unsortable in $(at,R)$ and $x\neq a$ and $ax\nsubseteq R$.
-Recall that by the hypothesis of the lemma, there exists $y$ that is sortable in $(t,R)$.
-Since $y$ is unsortable in $(at,R)$, Lemma 7 implies that either $y=a$ or $ay\subseteq R$.
-In either case, $y$ must follow the last occurrence of $x$ in $R$, because $ax\nsubseteq R$.
-However, with Lemma 5, the fact that $y$ is sortable in $(t,R)$ while $x$ is unsortable in $(t,R)$, and the fact that $x$ is one of the first two colors to appear in $tR$, imply that $y$ cannot follow the last occurrence of $x$ in $R$, a contradiction.
-∎
-Claim 19.
-There are distinct colors $c$ and $b$ that are distinct from $a$, $r$, and $t$, such that there exists $P\in\{arat,rart,arbrt,arcbct\}$ such that $P\subseteq R$.
-In particular, we have $art\subseteq R$.
-Proof.
-First, suppose that $rat\subseteq R$.
-Since, by Claim 18, we have $ar\subseteq R$, one of $arat$, $rart$, or $ratr$ appears in $R$.
-Since $atdratr=atdraI(r)tr\cong I_{1,0}$, we can assume the subordering $ratr$ does not appear $R$, so we are done in this case.
-Now, suppose $rat\nsubseteq R$.
-Since $r$ is the first color in $R$ that is not $a$ or $t$, $rat\nsubseteq R$ implies that the first sandwich that blocks $t$ in $atR$ does not include the color $a$.
-So there exist distinct colors $c$ and $b$ that are both neither $a$ nor $t$, such that a sandwich of the form $cbc$ blocks $t$ in $atR$.
-Since $r$ is the first color that is not $a$ or $t$ to appear in $R$, we have either $r=c$ and $rbrt\subseteq R$ or $r\notin\{b,c\}$ and $rcbct\subseteq R$.
-By Claim 18, we have $at\subseteq R$.
-This, with the fact that $rat\nsubseteq R$, implies that either $arbrt\subseteq R$ or $arcbct\subseteq R$, where the distinct colors $b$ and $c$ are distinct from $a$, $r$, and $t$.
-∎
-Claim 20.
-For distinct colors $a$, $r$, $t$, $b$, and $c$ and every $P\in\{arat,rart,arbrt,arcbct\}$ the color $a$ is the only sortable color in $atP$ and $a$ is a good sortable color in $atP$.
-Therefore, no sandwich consisting only of the colors that appear in $atP$ blocks $a$ in $X$.
-Proof.
-It is not hard to see that $a$ is the only sortable color in $atP\subseteq S^{\prime}R$.
-Furthermore, it is not hard to see that $a$ is a good sortable color in $atP\subseteq S^{\prime}R$.
-Recall that, by assumption, $X-d$ is sortable.
-This means that for any sorting sequence that sorts $X-d$, the color $a$ must be sorted first among the colors in $atP$.
-Therefore, there is no sandwich consisting only of the colors that appear in $atP$ that blocks $a$ in $X$.
-∎
-Note that Claims 19 and 20 establish the first statement of the conclusion.
-Claim 21.
-Let $P\in\{arat,rart,arbrt,arcbct\}$ such that $P\subseteq R$ where $a$, $r$, $t$, $b$, and $c$ are distinct colors. One of the following holds.
-  1. (i)
-    $dat$ or $adt$ is a subordering of $S^{\prime}$ or $atde\subseteq S^{\prime}$ for some color $e\notin P$; or
-  2. (ii)
-    $taxat$, $atxat$, $txaxt$, or $taxyxt$ is a subordering of $R^{\prime}$ where $a$, $t$, $x$, and $y$ are distinct colors; or
-  3. (iii)
-    $eP\subseteq R^{\prime}$ for some color $e\notin P$ and $eat$, $aet$, or $ate$ is a subordering of $S^{\prime}$.
-Proof.
-By assumption, there exists a $d$-sandwich that blocks $a$ in $X=S^{\prime}dR^{\prime}$.
-That is, for some color $e$, we either have a sandwich of the form $ded$ which blocks $a$ or a sandwich of the form $ede$ which blocks $a$.
-Therefore, we have the following two cases.
-Case 1: A sandwich of the form $ded$ blocks $a$ in $X$.
-Note that since $d\notin R^{\prime}$, we have $de\subseteq S^{\prime}$.
-So, if neither $dat\subseteq S^{\prime}$ nor $adt\subseteq S^{\prime}$, then $atde\subseteq S^{\prime}$ and $e\neq t$.
-So to show that $X$ satisfies (i), we only need to show that $e\notin P$.
-First note that $e\neq a$, since $ded$ blocks $a$.
-If $e=r$, then, since Claim 19 implies that $art\subseteq P$, we have $X\supseteq S^{\prime}dP\supseteq atdrdart\cong abdtdatb\cong C_{2,5}$, a contradiction.
-We get a similar contradiction if $P=arbrt$ or $P=arcbct$ and $e=b$, since in either case $abt\subseteq P$, so $S^{\prime}dP\supseteq atdbdabt\cong abdtdatb\cong C_{2,5}$, a contradiction.
-If $P=arcbct$ and $e=c$, then $act\subseteq P$, so $S^{\prime}dP\supseteq atdcdact\cong abdtdatb\cong C_{2,5}$, a contradiction.
-With Claim 19, this completes the proof of this case.
-Case 2: A sandwich of the form $ede$ blocks $a$ in $X$.
-Since $d\notin R$, we have $e\in S^{\prime}$.
-Furthermore, if $de\subseteq S^{\prime}$, then a sandwich of the form $ded$ blocks $a$ in $X$, and this was handled in Case 1.
-Therefore, we can assume that $ea\subseteq R^{\prime}$.
-Case 2.1: $e=t$.
-In this case, $ta=ea\subseteq R^{\prime}$ and, by Claim 20 a sandwich of the form $trt$ does not block $a$ in $X=S^{\prime}dR^{\prime}$.
-The fact that $at\subseteq S^{\prime}$ then implies that $rta\nsubseteq R^{\prime}$.
-Therefore, there exists a $t$ in $R^{\prime}$ that is before both an $a$ in $R^{\prime}$ and the first $r$ in $R^{\prime}$.
-So, if $P=arat$, then we have either $tarat=eP$ or $atrat$ in $R^{\prime}$, so (ii) holds.
-Furthermore, if $P=rart$, we have $trart=eP$ in $R^{\prime}$, so (ii) holds.
-If $P=arbrt$, then one option is $tarbrt=eP\subseteq R^{\prime}$, and (ii) holds in this case.
-By Claim 20, there is no $rbr$ sandwich that blocks $a$, so the only other possibilities when $P=arbrt$ are $atrabrt\subseteq R^{\prime}$ or $atrbart\subseteq R^{\prime}$, and, in either case, we have $trart\subseteq R^{\prime}$, so (ii) holds.
-Finally, assume $P=arcbct$.
-If $tarcbct=eP\subseteq R^{\prime}$, then $tacbct\subseteq R^{\prime}$, so (ii) holds.
-Otherwise, because there is no $cbc$ sandwich that blocks $a$, at least one of $atracbct\supseteq tacbct$, $atrcabct\supseteq tcact$, or $atrcbact\supseteq tcact$ is a subordering of $R^{\prime}$, so (ii) holds.
-Case 2.2: $e=r$.
-In this case, $r\in S^{\prime}$ and $ra=ea\subseteq R^{\prime}$.
-Recall that, by Claim 20, there is no sandwich of the form $rtr$ that blocks $a$ in $X$.
-Therefore, $r$ must follow $at$ in $S^{\prime}$, so $atr\subseteq S^{\prime}$.
-Because $art\subseteq P$ and $a$ is blocked by a $d$-sandwich, Lemma 10 (with $r$ playing the role of $b$) implies that a subpattern of $X$ is in $\Gamma$, a contradiction.
-Case 2.3: $e\neq r$ and $e\neq t$.
-Recall that we have $e\in S^{\prime}$ and $ea\subseteq R^{\prime}$.
-Since $e\in S^{\prime}$ and $at\subseteq S^{\prime}$, one of $eat$, $aet$, or $ate$ is in $S^{\prime}$.
-We will first prove that $e\notin P$ and $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
-Assume first that $e\in R$ and recall that $r$ was defined to be the first color in $R$ that is neither $t$ nor $a$.
-Therefore, $rea\subseteq R^{\prime}$, which implies that a sandwich of the form $ere$ blocks $a$ in $X\supseteq S^{\prime}dR^{\prime}$.
-Note that this, with Claim 20, implies that $e\notin P$.
-Recall that Claim 20 implies that every color in $atP$ is unsortable except $a$.
-If $e$ is blocked by a sandwich consisting of colors only in $P$, then the subordering formed by removing all colors except those in $P$ and $e$ from $X-d$ is trivially unsortable.
-This contradicts the assumption that all proper suborderings of $X$ are sortable.
-Now assume that $e\notin R$.
-In this case, $P\subseteq R$ implies that $e\notin P$.
-By the definition of $R^{\prime}$, the fact that $e\in R^{\prime}$ implies that the last $e$ follows the last $d$ in $X$.
-Since $e\notin R$ and the state after $d$ is sorted is $(S,R)$, the color $e$ was sorted before the color $d$.
-Since $d$ was sorted before any color in $P$, the color $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
-So, in both cases, $e\notin P$ and $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
-In particular, the subordering $ata$ does not block $e$ in $X$.
-This means that $e$ does not follow $a$ in $R^{\prime}$.
-Note first that if $eP\subseteq R^{\prime}$, then (iii) holds.
-So we can assume $eP\nsubseteq R^{\prime}$.
-This, with the fact that $e$ does does not follow $a$ in $R^{\prime}$ and $e\in R^{\prime}$, implies that $P=rart$ and $reart\subseteq R^{\prime}$.
-Therefore, at least one of $eatdreart$, $aetdreart$, or $atedreart$ is a subordering of $X=S^{\prime}dR^{\prime}$.
-This is a contradiction, because $eatdreart\supseteq eatrear\cong dabtdat\cong I_{1,0}$, $aetdreart\supseteq aetreart\cong adbtdatb\cong\mathcal{C}_{2,4}$, and $atedreart\supseteq atereart\cong abdtdatb\cong\mathcal{C}_{2,5}$.
-∎
-When Claim 21(i) holds, one of the orderings in the following table is a subordering of $X$.
-$datdarat\cong\mathcal{C}_{1,1}$ $datdrart\cong\mathcal{C}_{1,7}$ $datdarbrt\cong\mathcal{C}_{1,12}$ $datdarcbct\succeq\mathcal{C}_{1,12}$
-$adtdarat\cong\mathcal{C}_{1,4}$ $adtdrart\cong\mathcal{C}_{1,8}$ $adtdarbrt\cong\mathcal{C}_{1,11}$ $adtdarcbct\succeq\mathcal{C}_{1,11}$
-$atdedarat\cong\mathcal{C}_{1,5}$ $atdedrart\cong\mathcal{C}_{1,9}$ $atdedarbrt\cong\mathcal{C}_{1,13}$ $atdedarcbct\succeq\mathcal{C}_{1,13}$
-If Claim 21(iii) holds, then by swapping $d$ and $e$, we can see that one of orderings in the preceding table is a subpattern of $X$.
-Finally, if Claim 21(ii) holds, then one of the orderings in the following table is a subpattern of $X$.
-$atdtarat\cong\mathcal{C}_{1,2}$ $atdatrat\cong\mathcal{C}_{1,3}$ $atdtrart\cong\mathcal{C}_{1,6}$ $atdtarbrt\cong\mathcal{C}_{1,10}$
-3. Further Directions
-While the process of foot sorting lends itself naturally to sorting with a stack, one could consider analogs of the sock sorting problem which utilize other stack configurations or other data structures entirely.
-3.1. Sorting with Multiple Stacks
-Defant and Kravitz introduced a notion of foot-sorting with multiple feet in [1], though with more stacks in series, the sorting process becomes increasingly computationally intensive.
-It would be interesting to investigate perhaps the simplest extension of our work, the basis of sock orderings which are foot-sortable with two stacks in series.
-Defant and Kravitz proved in [1] that for any number $n\geq 1$, one may construct a sock ordering which is unsortable with $n$ stacks.
-Using this proof, one can check that the sock ordering of the form $(abcde)^{1440}$ (that is, the block $abcde$, where $a,b,c,d$ and $e$ are distinct sock colors, repeated 1440 times) is unsortable with two stacks in series, but it might be challenging to construct a single sock ordering that is minimally unsortable in this setting.
-3.2. Deque Sorting
-Because a deque, or double-ended queue, allows for more access to the socks for addition and removal, one might think about deque-sorting for socks independently of a foot; a sock could be placed either over an outermost sock or within an innermost sock, and the same is possible for sock removal.111 Using a deque in the context of permutation sorting has been consider previously (e.g.
-see [3], [2], and [4]).
-This modifies the original operations of foot sorting, using the language of the deque, to be
-  1. ($\text{stack}_{\uparrow}$)
-    Take the leftmost sock on the right and place it on top of the deque (potentially over any socks that are already in the deque).
-  2. ($\text{stack}_{\downarrow}$)
-    Take the leftmost sock on the right and place it on the bottom of the deque (potentially under any socks that are already in the deque).
-  3. ($\text{unstack}_{\uparrow}$)
-    Remove the topmost sock from the deque and place it so that it becomes the rightmost sock on the left.
-  4. ($\text{unstack}_{\downarrow}$)
-    Remove the bottommost sock from the deque and place it so that it becomes the rightmost sock on the left.
-While the addition of possible moves in this scenario makes for more difficult computation, some similar ideas hold as in the stack setting: in particular, it is now necessary to avoid the pattern $abab$ from appearing in the deque, just as it is necessary to avoid $aba$ appearing in the stack.
-It would be interesting to enumerate the basis of deque-sortable sock orderings, just as we have enumerated the basis of foot-sortable sock orderings.
-It is straightforward to show that the basis of deque-sortable sock orderings is different from $\Gamma$, as well as to show that this basis is infinite.
-Very closely mirroring the proof of Observation 2 under the relaxed restrictions of the deque, one can generate the class of sock orderings of the form $dabtI_{1,n}$ for any $n\geq 0$ and verify that such sock orderings are minimally unsortable with the deque.
-However, with the increased computational difficulty in the deque setting, we do not know whether other minimal deque-unsortable sock orderings may be generated from $\Gamma$ in a similar fashion.